### v1

# Optimality of quasi-Monte Carlo methods and suboptimality of the sparse-grid Gauss–Hermite rule

Abstract
Optimality of several quasi-Monte Carlo methods and suboptimality of the sparse-grid quadrature based on the univariate Gauss–Hermite rule is proved in the Sobolev spaces of mixed dominating smoothness of order $\alpha$, where the optimality is in the sense of worst-case convergence rate. For sparse-grid Gauss–Hermite quadrature, lower and upper bounds are established, with rates coinciding up to a logarithmic factor. The dominant rate is found to be only $N^{-\alpha/2}$ with $N$ function evaluations, although the optimal rate is known to be $N^{-\alpha}(\ln N)^{(d-1)/2}$. The lower bound is obtained by exploiting the structure of the Gauss–Hermite nodes and is independent of the quadrature weights; consequently, no modification of the weights can improve the rate $N^{-\alpha/2}$. In contrast, several quasi-Monte Carlo methods with a change of variables are shown to achieve the optimal rate, some up to, and one including, the logarithmic factor.
1 Introduction
This paper is concerned with numerical integration of multivariate functions over the whole space in high dimensions. Our focus is on integration with respect to the standard Gaussian measure:
$I(f):=\int_{\mathbb{R}^{d}}f(\boldsymbol{x})\frac{\mathrm{e}^{-|\boldsymbol{x}|^{2}/2}}{(2\pi)^{d/2}}\mathrm{d}\boldsymbol{x}\approx Q_{N}(f),$
where $|\boldsymbol{x}|=\sqrt{x_{1}^{2}+\dotsb+x_{d}^{2}}$ for $\boldsymbol{x}\in\mathbb{R}^{d}$ is the Euclidean norm, and $Q_{N}$ is a suitable numerical integration rule using $N$ evaluations of $f$. Such problems arise across a broad range of applications, including data assimilation [36], robotics [3], life insurance [23], physics [45]. They also serve as simpler problems to which more complex integration tasks can be reduced [20].
In high dimensions $d\gg 1$, two popular classes of numerical integration rules $Q_{N}$ are sparse-grid quadrature [6] and quasi-Monte Carlo (QMC) methods [13]. Sparse-grid quadrature is a numerical integration technique based on univariate quadrature rules. Unlike fully tensorized quadrature, it significantly reduces the number of evaluation points, making it attractive for high-dimensional problems. A commonly used univariate quadrature for approximating $I(f)$ is the Gauss–Hermite rule [38, 56, 19, 4, 66], which is the focus of this paper. Quasi-Monte Carlo, on the other hand, is an equal-weight numerical integration. Its core idea is to carefully choose the quadrature points to improve accuracy. Examples of point sets include low-discrepancy point sets and sequences, lattice rules, and digital nets; these classes are not mutually exclusive. These methods are also widely used for integration with respect to the Gaussian measure and have found applications across a diverse range of scientific fields, such as uncertainty quantification [9, 42], computational physics [37], finance [24, 43], and machine learning [1, 47].
In this paper, we show that sparse-grid Gauss–Hermite quadrature is suboptimal, whereas several quasi-Monte Carlo methods are optimal in terms of convergence rate. More precisely, for $L^{2}$-Sobolev spaces of degree $\alpha$, the sparse-grid Gauss–Hermite rule achieves the convergence rate $O(N^{-\alpha/2})$; several quasi-Monte Carlo (QMC) methods together with change of variables achieve $O(N^{-\alpha})$ up to a logarithmic factor. The rate $O(N^{-\alpha})$ is in fact optimal, as there exists a matching lower bound established by Dick et al. [11]. In this sense, QMC methods considered herein are rate-optimal. In contrast, we also show that the rate $O(N^{-\alpha/2})$ for the sparse-grid Gauss–Hermite rule likewise unimprovable, by proving a corresponding lower bound.
Such a gap in convergence rates between the sparse-grid Gauss–Hermite rule and QMC methods has also been observed numerically by Dick et al. [11] and in subsequent work by Nuyens and one of the present authors [55]. Our theoretical results provide a rigorous explanation of these empirical findings.
Analysis of the sparse-grid Gauss–Hermite rule and QMC methods on the whole space has been carried out before, with many papers set in the context of partial differential equations with random coefficients [7, 15, 17, 21, 30, 31, 32, 40, 49, 58]. This paper provides theoretical results for a standard function space, without specific structure of the underlying model problem such as the form of the random field or partial differential equations whose differential operator is given by such a field. In the same vein, aforementioned work by Dick et al. [11] studied quasi-Monte Carlo methods in a Sobolev space, the space we also consider herein. They showed that the QMC method considered there attains the optimal convergence rate by proving matching upper and lower bounds, although the logarithmic factor was not optimal. Nuyens and one of the present authors considered a rank-$1$ lattice QMC over $\mathbb{R}^{d}$ [55] for functions with suitable decay. We revisit their results and show that the resulting QMC method achieves the optimal rate in the Sobolev space considered herein, although the logarithmic factor is again not optimal. In [64], Hyvönen, Karvonen, and one of the present authors considered a change of variables that maps functions on the real line to the unit interval. Although such an approach typically incurs undesirable unboundedness near the boundary, they showed that, with a carefully chosen Möbius transformation, the resulting mapping sends univariate functions from the Gaussian Sobolev space on the real line to the periodic Sobolev space of the same smoothness on the unit interval. In this paper, by applying such a transformation component-wise to $d$-variate functions, we show that the resulting function belongs to two classes of Sobolev spaces that are important in QMC: Korobov spaces and unanchored Sobolev spaces. Using this, we establish that QMC rules with the Möbius transformation inherit the convergence rates of the respective spaces. These rates turn out to be optimal in Gaussian Sobolev spaces: rank-$1$ lattice rules achieve the optimal rate up to a logarithmic factor, whereas higher-order digital nets achieve the optimal rate including the logarithmic factor.
Dũng and Nguyen [16] proposed two numerical integration methods that are optimal – including the logarithmic factor – in $L^{p}$-Sobolev spaces for $1<p<\infty$. One method requires implementing a partition of unity with suitable properties. The other method appears to be implementable, but the quadrature weights depend on various indices. The results in this paper show that equal-weight quadrature rules suffice to obtain an optimal numerical integration rule.
The remainder of the paper is organized as follows. In Section 2 we introduce necessary concepts and state preliminary results. Sections 3 and 4 present our main results on the sparse-grid Gauss–Hermite quadrature and QMC methods, respectively. These results build on the authors’ previous works [29, 41, 64]. Finally, Section 5 concludes the paper.
2 Preliminary
2.1 The Sobolev space
We consider the integration problem in the Gaussian Sobolev space, also known as the Hermite space of finite smoothness. This space is defined by
$\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}:=\|f\|_{H_{\rho}^{\alpha}}:=\left(\sum_{|\mathbf{r}|_{\infty}\leq\alpha}\|D^{\mathbf{r}}f\|_{L_{\rho}^{2}(\mathbb{R}^{d})}^{2}\right)^{\frac{1}{2}},$
where $\rho(\boldsymbol{x})=\frac{\mathrm{e}^{-|\boldsymbol{x}|^{2}/2}}{(2\pi)^{d/2}}$ and $D^{\mathbf{r}}f$ is the weak derivative of $f$. Here, $\|\cdot\|_{L_{\rho}^{2}(\mathbb{R}^{d})}$ is the norm induced by the weighted $L^{2}$-inner product $\langle f,g\rangle_{L_{\rho}^{2}}:=\int_{\mathbb{R}^{d}}f(\boldsymbol{x})g(\boldsymbol{x})\rho(\boldsymbol{x})\mathrm{d}\boldsymbol{x}$ for $f,g\in L_{\rho}^{2}(\mathbb{R}^{d})$.
It is known that this space and the so-called Hermite space coincide as vector spaces, with equivalent norms. More specifically, Dick et al. [11] studied quasi-Monte Carlo methods in the Hermite space and showed that every element of this space also belongs to $H_{\rho}^{\alpha}(\mathbb{R}^{d})$, with the embedding being continuous; the reverse inclusion is discussed in [41, 16]. This space is also known to coincide with the so-called modulation space, as discussed in [18]. Further properties of Hermite spaces are discussed in [25, 26].
In the space $H_{\rho}^{\alpha}(\mathbb{R}^{d})$, a lower bound on the worst-case integration error is known for general algorithms.
Theorem 2.1 ([11]).
For $N\geq 2$, let $A_{N}\colon H_{\rho}^{\alpha}(\mathbb{R}^{d})\to\mathbb{R}$ be a mapping (linear or nonlinear) that uses only $N$ values as information about the argument, i.e., it is of the form $A_{N}(f)=\mathcal{I}_{N}(f(\boldsymbol{x}_{1}),\dots,f(\boldsymbol{x}_{N}))$ for a mapping $\mathcal{I}_{N}\colon\mathbb{R}^{N}\to\mathbb{R}$. Then we have
$\sup_{\begin{subarray}{c}f\in H_{\rho}^{\alpha}(\mathbb{R}^{d})\\ \|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}\leq 1\end{subarray}}|I(f)-A_{N}(f)|\geq c_{\alpha,d}\frac{(\ln N)^{\frac{d-1}{2}}}{N^{\alpha}}.$
Hence, any algorithm achieving this convergence rate $\frac{(\ln N)^{\frac{d-1}{2}}}{N^{\alpha}}$ may be regarded as optimal, in terms of the worst-case error convergence rate. We will show that several QMC methods attain a rate of the form $\frac{(\ln N)^{\gamma}}{N^{\alpha}}$, where the exponent $\gamma$ depends on the method. The rate $\gamma=\frac{d-1}{2}$ turns out to be attainable, hence optimal in the sense above. In contrast, the sparse-grid quadrature based on the Gauss–Hermite rule is of the rate $N^{-\alpha/2}$, up to a logarithmic factor, which is subotimal in the sense of worst-case error convergence rate.
Note that [11, Theorem 1] is formulated for linear algorithms. As noted in [11, Section 3], however, in view of for example [2] (see also [53, Section 4.2]), their lower bound also holds for nonlinear algorithms. Moreover, the implied constant in the result above may differ from that in [11, Theorem 1], since the present bound is formulated for the Sobolev space, whereas [11, Theorem 1] considers a norm-equivalent space, namely the Hermite space.
Let $H_{k}$ for $k\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$ denote the normalized probabilistic Hermite polynomial of degree $k$, defined by
$H_{k}(x)=\frac{(-1)^{k}}{\sqrt{k!}}\mathrm{e}^{x^{2}/2}\frac{\mathrm{d}^{k}}{\mathrm{d}x^{k}}\mathrm{e}^{-x^{2}/2}.$
The multivariate counterpart is defined by $H_{\mathbf{k}}(\boldsymbol{x})=\prod_{j=1}^{d}H_{k_{j}}(x_{j})$ for $\mathbf{k}\in\mathbb{N}_{0}^{d}$. The collection $(H_{\mathbf{k}})_{\mathbf{k}\in\mathbb{N}_{0}^{d}}$ forms a complete orthonormal system for $L_{\rho}^{2}(\mathbb{R}^{d})$, with the normalization above ensuring $\|H_{\mathbf{k}}\|_{L_{\rho}^{2}}=1$ for all $\mathbf{k}\in\mathbb{N}_{0}^{d}$.
Throughout the paper, for mappings $f,g\colon S\to[0,\infty]$ on any set $S$, we write $f\lesssim g$ if there exists a constant $c\in(0,\infty)$ such that $f(x)\leq cg(x)$ holds for all $x\in S$. We write $f\asymp g$ if both $f\lesssim g$ and $g\lesssim f$. If the constant may depend on specific parameters, we indicate this by subscripts. For example, $f\lesssim_{d}g$ means there exists a constant $c_{d}\in(0,\infty)$, which may depend on $d$, such that $f(x)\leq c_{d}\,g(x)$ for all $x\in S$; likewise, $f\asymp_{d}g$ means $f\lesssim_{d}g$ and $g\lesssim_{d}f$.
2.2 Tensorised operators
In this section we define the tensorised operator $T_{1}\otimes\dotsb\otimes T_{d}\colon H_{\rho}^{\alpha}(\mathbb{R}^{d})\to H_{\rho}^{\alpha}(\mathbb{R}^{d})$ based on bounded linear operators $T_{j}\colon H_{\rho}^{\alpha}(\mathbb{R})\to H_{\rho}^{\alpha}(\mathbb{R})$, $j=1,\dots,d$. The main motivation for this subsection is to construct such an expression on $H_{\rho}^{\alpha}(\mathbb{R}^{d})$ without using the norm-equivalence with the Hermite space, which will be used to discuss the Smolyak algorithm.
Proposition 2.2.
The $\mathbb{R}$-vector space spanned by $d$-variate polynomials $\mathcal{P}:=\mathrm{span}\Big\{\prod_{j=1}^{d}H_{k_{j}}\mid k_{j}\in\mathbb{N}_{0},\,j=1,\dots,d\Big\}$ is dense in $H_{\rho}^{\alpha}(\mathbb{R}^{d})$.
Proof.
Fix $f\in H_{\rho}^{\alpha}$. Following the argument in [41, Lemma 2.1] we see
$\langle D^{\mathbf{r}}f,H_{\mathbf{k}}\rangle_{L_{\rho}^{2}}=\biggl(\prod_{j=1}^{d}\prod_{m=1}^{\mathrm{r}_{j}}\bigl(k_{j}+m\bigr)\biggr)^{1/2}\langle f,H_{\mathbf{k}+\mathbf{r}}\rangle_{L_{\rho}^{2}}\quad\text{for }\text{$\mathbf{k}$}\in\mathbb{N}_{0}^{d},\ \mathbf{r}\in\{0,\dots,\alpha\}^{d}.$
With this, using the Parseval identity for derivatives we see that the polynomial $f_{N}=\sum_{|\mathbf{k}|\leq N}\langle f,H_{\mathbf{k}}\rangle_{L_{\rho}^{2}}H_{\mathbf{k}}$ can be made arbitrarily close to $f$ in $H_{\rho}^{\alpha}$:
$\|f-f_{N}\|_{H_{\rho}^{\alpha}}^{2}=\sum_{|\mathbf{r}|_{\infty}\leq\alpha}\sum_{|\mathbf{k}|>N}\biggl(\prod_{j=1}^{d}\prod_{m=1}^{\mathrm{r}_{j}}\bigl(k_{j}+m\bigr)\biggr)|\langle f,H_{\mathbf{k}+\mathbf{r}}\rangle_{L_{\rho}^{2}}|^{2}.$
Let $T_{j}\colon H_{\rho}^{\alpha}(\mathbb{R})\to H_{\rho}^{\alpha}(\mathbb{R})$ be bounded linear operators. Define $T_{1}\otimes\dotsb\otimes T_{d}:\mathcal{P}\to H_{\rho}^{\alpha}(\mathbb{R}^{d})$ by defining
$\left[T_{1}\otimes\dotsb\otimes T_{d}\Bigg(\prod_{j=1}^{d}H_{k_{j}}\Bigg)\right](\boldsymbol{x})=\prod_{j=1}^{d}[T_{j}(H_{k_{1}})](x_{j})\qquad\text{for }\prod_{j=1}^{d}H_{k_{j}}\in\mathcal{P}$
and extending it linearly to $\mathcal{P}$.
Then, using the argument for tensor product Hilbert spaces [57, Section VIII.10] we see $\sup_{p\in\mathcal{P},\|p\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}=1}\|T_{1}\otimes\dotsb\otimes T_{d}(p)\|=\prod_{j=1}^{d}\|T_{j}\|_{H_{\rho}^{\alpha}(\mathbb{R})\to H_{\rho}^{\alpha}(\mathbb{R})}$ on $\mathcal{P}$, and from Proposition 2.2, the operator $T_{1}\otimes\dotsb\otimes T_{d}$ extends continuously to $H_{\rho}^{\alpha}(\mathbb{R}^{d})$ with the same operator norm.
Note that any continuous linear functional on $H_{\rho}^{\alpha}(\mathbb{R}^{d})$ is a bounded linear operator in our setting. Indeed, since the total measure is normalised to $1$, any constant belongs to $H_{\rho}^{\alpha}(\mathbb{R}^{d})$, and if $T_{j}$ is a functional then we have $T_{j}(f)\in H_{\rho}^{\alpha}(\mathbb{R})$ with $\|T_{j}(f)\|_{H_{\rho}^{\alpha}(\mathbb{R})}=|T_{j}(f)|$ for $f\in H_{\rho}^{\alpha}(\mathbb{R})$. Operators with the symbol $\otimes$ in this paper are constructed explicitly in this manner.
2.3 Sobolev embedding for spaces with dominating mixed smoothness
Let $W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})$ denote the Sobolev space of dominating mixed first-order smoothness, equipped with the norm
$\|v\|_{W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})}:=\left(\sum_{|\mathbf{r}|_{\infty}\leq 1}\|D^{\mathbf{r}}v\|_{L^{2}(\mathbb{R}^{d})}^{2}\right)^{\frac{1}{2}},$
where $\|\cdot\|_{L^{2}(\mathbb{R}^{d})}$ is the standard $L^{2}(\mathbb{R}^{d})$-norm. Later, we will use the following Sobolev inequality for $W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})$, which is also of independent interest.
Proposition 2.3.
Every element of $W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})$ admits a bounded continuous representative $v$ satisfying
$\sup_{\boldsymbol{x}\in\mathbb{R}^{d}}|v(\boldsymbol{x})|$ $\leq\|v\|_{W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})},$ (1)
and
$|v(\boldsymbol{x})-v(\boldsymbol{y})|$ $\leq d^{3/4}\|v\|_{W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})}|\boldsymbol{x}-\boldsymbol{y}|_{2}^{1/2}\qquad\text{for all }\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}.$ (2)
Proof.
We will first prove the statement for infinitely differentiable functions with compact support, $C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$, and then extend it by a density argument. For $v\in C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$, following [5, Proof of Theorem 8.8] we get
$|v(x_{1},\dots,x_{d})|^{2}$ $\leq 2\sqrt{\int_{\mathbb{R}}|v(s,x_{2},\dots,x_{d})|^{2}\mathrm{d}s}\sqrt{\int_{\mathbb{R}}|D^{(1,0,\dots,0)}v(s,x_{2},\dots,x_{d})|^{2}\mathrm{d}s}$
$\leq\int_{\mathbb{R}}|v(s,x_{2},\dots,x_{d})|^{2}\mathrm{d}s+\int_{\mathbb{R}}|D^{(1,0,\dots,0)}v(s,x_{2},\dots,x_{d})|^{2}\mathrm{d}s.$
Applying the same argument for $v(s,x_{2},\dots,x_{d})$ and $D^{(1,0,\dots,0)}v(s,x_{2},\dots,x_{d})$ with respect to the second argument $x_{2}$, and repeating the procedure for $x_{3}$ up to $x_{d}$, we obtain
$|v(x_{1},\dots,x_{d})|^{2}\leq\sum_{r_{1},\dots,r_{d}\in\{0,1\}}\int_{\mathbb{R}}\dotsb\int_{\mathbb{R}}|D^{(r_{1},\dots,r_{d})}v(s_{1},\dots,s_{d})|^{2}\mathrm{d}s_{1}\dotsb\mathrm{d}s_{d},$
which gives $\sup_{\boldsymbol{x}\in\mathbb{R}^{d}}|v(\boldsymbol{x})|\leq\|v\|_{W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})}$ for $v\in C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$.
Now fix any representative $v\in W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})$. The density of $C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$ in $W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})$ can be checked using a standard argument from the Sobolev space theory, for example by approximation via smoothly truncated mollified functions. Thus, there exists a sequence $(v_{n})\subset C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$ that approximates $v$ in $W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})$. The sequence $(v_{n})$ also belongs to the Banach space $C_{\mathrm{b}}(\mathbb{R}^{d})$ of bounded continuous functions equipped with the supremum norm. The inequality already proved gives
$\sup_{\boldsymbol{y}\in\mathbb{R}^{d}}|v_{n}(\boldsymbol{y})-v_{m}(\boldsymbol{y})|$ $\leq\|v_{n}-v_{m}\|_{W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})},$
so, $(v_{n})$ is Cauchy in $C_{\mathrm{b}}(\mathbb{R}^{d})$ and thus has a limit $\tilde{v}$ in $C_{\mathrm{b}}(\mathbb{R}^{d})$, which in particular converges to $\tilde{v}$ pointwise. Since a subsequence of $(v_{n})$ converges to $v$ a.e., we conclude $\tilde{v}=v$ a.e. This also implies that both sides of $\sup_{\boldsymbol{x}\in\mathbb{R}^{d}}|v_{n}(\boldsymbol{x})|\leq\|v_{n}\|_{W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})}$ admit limits, which yields the desired inequality for $v\in W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})$.
To show the second inequality, again take $v\in C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$. For any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}$, we have
$|v(\boldsymbol{x})-v(\boldsymbol{y})|$ $\leq|v(\boldsymbol{x})-v(y_{1},x_{2},\dots,x_{d})|$
$\quad+\sum_{j=1}^{d-2}|v(y_{1},\dots,y_{j},x_{j+1},\dots,x_{d})-v(y_{1},\dots,y_{j+1},x_{j+2},\dots,x_{d})|$
$\quad+|v(y_{1},\dots,y_{d-1},x_{d})-v(\boldsymbol{y})|.$
The first term is controlled by
$|v(\boldsymbol{x})-v(y_{1},x_{2},\dots,x_{d})|\leq\left(\int_{\mathbb{R}}|D^{(1,0,\dots,0)}v(s,x_{2},\dots,x_{d})|^{2}\mathrm{d}s\right)^{1/2}|x_{1}-y_{1}|^{1/2},$
which can be further bounded by $\|v\|_{W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})}|x_{1}-y_{1}|^{1/2}$, following the argument above. The other terms are treated similarly. Therefore, we obtain $|v(\boldsymbol{x})-v(\boldsymbol{y})|\leq\|v\|_{W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})}\sum_{j=1}^{d}|x_{j}-y_{j}|^{1/2}$. Finally, using the Hölder inequality with $p=4,q=4/3$ we get
$\sum_{j=1}^{d}|x_{j}-y_{j}|^{1/2}\leq\bigg(\sum_{j=1}^{d}|x_{j}-y_{j}|^{2}\bigg)^{1/4}\big(\sum_{j=1}^{d}\;1\big)^{3/4}=d^{3/4}|\boldsymbol{x}-\boldsymbol{y}|_{2}^{1/2},$
which shows the second inequality for $C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$. Using the density of $C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$ completes the proof. ∎
3 Gauss–Hermite sparse-grid quadrature
A sparse-grid quadrature is constructed based on univariate quadrature rules. For $\ell\in\mathbb{N}_{0}$, let $Q_{\ell}^{\mathrm{uni}}\colon H_{\rho}^{\alpha}(\mathbb{R})\to\mathbb{R}$ a quadrature rule with $n_{\ell}$ quadrature points, with $Q_{0}^{\mathrm{uni}}:=0$, i.e., $Q_{0}^{\mathrm{uni}}(f)=0$ for all $f\in H_{\rho}^{\alpha}(\mathbb{R})$. Define $\Delta_{0}:=Q_{0}^{\mathrm{uni}}=0$, and for $\ell\geq 1$, $\Delta_{\ell}:=Q_{\ell}^{\mathrm{uni}}-Q_{\ell-1}^{\mathrm{uni}}$. Since constant functions belong to $H_{\rho}^{\alpha}(\mathbb{R})$ and $H_{\rho}^{\alpha}(\mathbb{R}^{d})$, the operator $\Delta_{\ell_{1}}\otimes\cdots\otimes\Delta_{\ell_{d}}\colon H_{\rho}^{\alpha}(\mathbb{R}^{d})\to\mathbb{R}$ is well-defined; see Section 2.2. Given an index set $\Lambda\subset\mathbb{N}^{d}$, the sparse-grid quadrature $S_{\Lambda}\colon H_{\rho}^{\alpha}(\mathbb{R}^{d})\to\mathbb{R}$ is defined by
$S_{\Lambda}:=\sum_{(\ell_{1},\dots,\ell_{d})\in\Lambda}\Delta_{\ell_{1}}\otimes\dotsb\otimes\Delta_{\ell_{d}}.$ (3)
The purpose of this section is to analyse the error of $S_{\Lambda}$ when $Q_{\ell}^{\mathrm{uni}}$ is the Gauss–Hermite quadrature. In what follows, unless otherwise stated, we use $S_{\Lambda}$ to denote the operator in (3) with Gauss–Hermite quadrature.
3.1 Lower bounds
In this section, we present lower bounds for two types of Smolyak quadratures: one for a general class of index sets (Theorem 3.1) and the other for the classical isotropic case (Theorem 3.2). In both cases, the dominant decay rate turns out to be half of the optimal rate as in Theorem 2.1.
For the isotropic case, we also provide a matching upper bound in Section 3.3, implying that our lower bound is sharp up to a logarithmic factor. Furthermore, a lower bound with the same polynomial factor holds for general index sets, implying that using an anisotropic index set, e.g., to exploit anisotropy in the target functions, does not improve the convergence rate.
To make the statement of the theorem concise, we say that an index set $\Lambda\subset\mathbb{N}^{d}$ is downward-closed in one direction if there is a coordinate index $j\in\{1,\dots,d\}$ such that whenever $(1,\dots 1,\ell,1\dots,1)\in\Lambda$ holds, where $\ell$ is in the $j$-th position, then for all integers $k\leq\ell$ we also have $(1,\dots 1,k,1\dots,1)\in\Lambda$. Clearly, downward-closed set is downward-closed in one direction.
Theorem 3.1.
Suppose that a finite index set $\Lambda\subset\mathbb{N}^{d}$ is downward-closed in one direction. Let $S_{\Lambda}$ be the corresponding sparse-grid quadrature based on the Gauss–Hermite rule. Let $N$ be the total number of the point evaluations used in $S_{\Lambda}$. Then, we have
$\sup_{f\neq 0}\frac{|I(f)-S_{\Lambda}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}\geq c_{\alpha}N^{-\alpha/2}.$
where the constant $c_{\alpha}>0$ is independent of $N$ and $d$.
Proof.
Our strategy follows [41, Section 3.1]. The heart of the matter is to construct a suitable fooling function, which we denote by $h_{\Lambda}$ below. The function is based on the earlier construction from [41, Section 3.1]. Without loss of generality, we assume that $\Lambda$ is downward-closed in the first coordinate. Then, for some $\overline{\ell}_{1}\in\mathbb{N}$ we have $(\ell_{1},1\dots,1)\in\Lambda$ for all $\ell_{1}=1,\dots,\overline{\ell}_{1}$. For example, if $\Lambda$ is the classical isotropic index set $\Lambda=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid\,|\boldsymbol{\ell}|_{1}\leq L\}$ then $\overline{\ell}_{1}=L-d+1$.
Let $n:=n_{\overline{\ell}_{1}}$ denote the number of quadrature points used in the univariate Gauss–Hermite rule $Q_{\overline{\ell}_{1}}^{\mathrm{uni}}$, and let $\xi_{1},\dots,\xi_{n}$ be the roots of the univariate Hermite polynomial $H_{n}$ of degree $n$. Following [41], define the function $p_{n}(t)\colon\mathbb{R}\to\mathbb{R}$ by
[…]
With this, define
$h_{\Lambda}(\boldsymbol{x}):=p_{n}(x_{1})\quad\text{for }\boldsymbol{x}=(x_{1},\dots,x_{d})\in\mathbb{R}^{d}.$ (4)
First, we will show
$\frac{|I(h_{\Lambda})-S_{\Lambda}(h_{\Lambda})|}{\|h_{\Lambda}\|_{H_{\rho}^{\alpha}}}=\frac{\int_{\mathbb{R}}p_{n}(x)\rho(x)\mathrm{d}x}{\left(\sum_{r=0}^{\alpha}\|D^{r}p_{n}\|_{L_{\rho}^{2}}^{2}\right)},$ (5)
so that the problem reduces to a one-dimensional integration problem. Recall $\Delta_{1}(1)=1$ and $\Delta_{\ell}(1)=0$ for $\ell\neq 1$ because the Gauss–Hermite rule integrates constant functions exactly. Thus, it follows
$S_{\Lambda}(h_{\Lambda})$ $=\sum_{(\ell_{1},\dots,\ell_{d})\in\Lambda}\Delta_{\ell_{1}}(p_{n})\Delta_{\ell_{2}}(1)\cdots\Delta_{\ell_{d}}(1)$
$=\sum_{\begin{subarray}{c}(\ell_{1},\dots,\ell_{d})\in\Lambda\\ \ell_{2}=\dotsb=\ell_{d}=1\end{subarray}}\Delta_{\ell_{1}}(p_{n})\Delta_{\ell_{2}}(1)\cdots\Delta_{\ell_{d}}(1)=Q_{\overline{\ell}_{1}}^{\mathrm{uni}}(p_{n}),$
Hence, $S_{\Lambda}(h_{\Lambda})=0$ and $|I(h_{\Lambda})-S_{\Lambda}(h_{\Lambda})|=\int_{\mathbb{R}}p_{n}(x)\rho(x)\mathrm{d}x$. As for the norm, since $h_{\Lambda}(\boldsymbol{x})=p_{n}(x_{1})$ we obtain
$\|h_{\Lambda}\|_{H_{\rho}^{\alpha}}$ $=\left(\sum_{|\mathbf{r}|_{\infty}\leq\alpha}\|D^{(r_{1},0,\dots,0)}h_{\Lambda}\|_{L_{\rho}^{2}}^{2}\right)^{\frac{1}{2}}=\left(\sum_{r=0}^{\alpha}\|D^{r}p_{n}\|_{L_{\rho}^{2}}^{2}\right)^{\frac{1}{2}}$
Therefore, (5) holds.
The right-hand side of (5) corresponds to the expression in the last displayed equation in [41, Proof of Lemma 3.1], which is further analysed in [41, Theorem 3.2]. There, it is shown
$\frac{\int_{\mathbb{R}}p_{n}(x)\rho(x)\mathrm{d}x}{\left(\sum_{r=0}^{\alpha}\|D^{r}p_{n}\|_{L_{\rho}^{2}}^{2}\right)}\geq c_{\alpha}n^{-\alpha/2},$
for some constant $c_{\alpha}>0$ independent of $N$. From
$n\leq\#\left(\sum_{\begin{subarray}{c}(\ell_{1},\dots,\ell_{d})\in\Lambda\\ \ell_{2}=\dotsb=\ell_{d}=1\end{subarray}}\Delta_{\ell_{1}}\otimes\dotsb\otimes\Delta_{\ell_{d}}\right)\leq\#\left(\sum_{(\ell_{1},\dots,\ell_{d})\in\Lambda}\Delta_{\ell_{1}}\otimes\dotsb\otimes\Delta_{\ell_{d}}\right)=N,$
the statement follows. ∎
For the particular choice of index set
$\Lambda_{L}=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid|\boldsymbol{\ell}|_{1}\leq L\},$
we establish a lower bound of a slower decay, now including a logarithmic factor that grows with the dimension; see Theorem 3.2.
The index set $\Lambda_{L}$ is of particular interest because it corresponds to the classical (isotropic) sparse-grid quadrature, which was Smolyak’s original choice of index set. Note that $S_{\Lambda_{L}}=S_{\Lambda_{L}^{0}}$, where
$\Lambda_{L}^{0}=\{\boldsymbol{\ell}\in\mathbb{N}_{0}^{d}\mid|\boldsymbol{\ell}|_{1}\leq L\}$
because $\Delta_{0}=0$. Moreover, if $L<d$ then $S_{\Lambda_{L}}=0$, since $\ell_{1}+\dotsb+\ell_{d}<L<d$ implies that at least one of the indices is zero, say $\ell_{j}=0$, and thus $\Delta_{\ell_{j}}=0$.
Theorem 3.2.
Let $S_{\Lambda_{L}}$ be the sparse-grid quadrature based on the Gauss–Hermite rule for $\Lambda_{L}=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid|\boldsymbol{\ell}|_{1}\leq L\}$. Suppose that the Gauss–Hermite rule $Q_{\ell}^{\mathrm{uni}}$ for univariate functions employs $n_{\ell}$ quadrature points for every $\ell\geq 1$, and that for some $M_{0}>0$ and $M>1$ we choose $n_{\ell}$ as
$M^{\ell-1}\leq n_{\ell}\leq M_{0}(M^{\ell}-1)\quad\text{and}\qquad n_{L-d+1}\geq\mathrm{e}^{d-1}.$
Let $N$ denote the total number of the function evaluations used in $S_{\Lambda_{L}}$. Then, we have
$\sup_{f\neq 0}\frac{|I(f)-S_{\Lambda_{L}}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}\geq c_{\alpha,d}N^{-\alpha/2}(\ln N)^{\frac{a}{2}(d-1)}.$
where the constant $c_{\alpha,d}>0$ is independent of $N$.
Proof.
Define $h_{\Lambda_{L}}(\boldsymbol{x}):=p_{n_{L-d+1}}(x_{1})$ for $\boldsymbol{x}=(x_{1},\dots,x_{d})\in\mathbb{R}^{d}$. From the proof of Theorem 3.1 we know
$\frac{|I(h_{\Lambda_{L}})-S_{\Lambda_{L}}(h_{\Lambda_{L}})|}{\|h_{\Lambda_{L}}\|_{H_{\rho}^{\alpha}}}\geq c_{\alpha}n_{L-d+1}^{-\alpha/2},$
so it remains to relate NN and $n_{L-d+1}$.
Given the choice $\Lambda_{L}=\bigl\{\boldsymbol{\ell}=(\ell_{1},\dots,\ell_{d})\in\mathbb{N}^{d}:|\boldsymbol{\ell}|_{1}\leq L\bigr\}$, and in view of the well-known representation formula of the Smolyak algorithm (see e.g., [54, Lemma 15.1]) the total number of point evaluation $N$ can be written explicitly as
$N=\sum_{\boldsymbol{\ell}\in\Lambda_{L}}\prod_{k=1}^{d}n_{\ell_{k}}.$ (6)
Hence $M^{\ell-1}\leq n_{\ell}\leq M_{0}(M^{\ell}-1)$ implies
$M^{-d}\sum_{\boldsymbol{\ell}\in\Lambda_{L}}M^{|\boldsymbol{\ell}|_{1}}\leq N\leq M_{0}^{d}\sum_{\boldsymbol{\ell}\in\Lambda_{L}}M^{|\boldsymbol{\ell}|_{1}},$
and thus together with $\sum_{\boldsymbol{\ell}\in\Lambda_{L}}M^{|\boldsymbol{\ell}|_{1}}=\sum_{k=d}^{L}M^{k}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{N}^{d}\\ |\boldsymbol{\ell}|_{1}=k\end{subarray}}1$ and $\{\boldsymbol{\ell}\in\mathbb{N}^{d}:|\boldsymbol{\ell}|_{1}=k\}=\frac{(k-1)!}{(k-d)!(d-1)!}=\Theta_{d}(k^{d-1})$ we obtain
$\underline{C}_{d}M^{L}L^{d-1}\leq\underline{C}^{\prime}_{d}M^{L}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{N}^{d}\\ |\boldsymbol{\ell}|_{1}=L\end{subarray}}1\leq N\leq\overline{C}_{d}L^{d-1}\sum_{k=0}^{L}M^{k}\leq\frac{\overline{C}_{d}M}{M-1}L^{d-1}M^{L}.$ (7)
Here, the implied constants depend on $d$. Hence, $N\asymp_{d}n_{L}(\ln n_{L})^{d-1}\asymp_{d}n_{L-d+1}(\ln n_{L-d+1})^{d-1}$.
Since $n_{L-d+1}\geq\mathrm{e}^{d-1}$ we have $(\ln n_{L-d+1})^{d-1}\leq n_{L-d+1}$, thus $\ln N\lesssim_{d}2\ln n_{L-d+1}$. Hence, we have
$n_{L-d+1}^{-\alpha/2}\asymp_{d}N^{-\alpha/2}(\ln n_{L-d+1})^{\frac{a}{2}(d-1)}\gtrsim_{d}N^{-\alpha/2}(\ln N)^{\frac{a}{2}(d-1)},$
which completes the proof. ∎
These bounds turn out to be sharp in the sense that for the standard isotropic Smolyak’s algorithm we have a matching upper bound, as we will see in Section 3.3.
3.2 Application: lower bounds for sparse-grid interpolation
Using similar arguments, analogous lower bounds follow for sparse-grid polynomial approximation $\mathscr{I}_{\Lambda}$ based on Gauss–Hermite interpolation. The approximation $\mathscr{I}_{\Lambda}$ is defined by
$\mathscr{I}_{\Lambda}$ $:=\sum_{\boldsymbol{\ell}\in\Lambda}(\mathscr{I}_{\ell_{1}}^{\mathrm{uni}}-\mathscr{I}_{\ell_{1}-1}^{\mathrm{uni}})\otimes\dotsb\otimes(\mathscr{I}_{\ell_{d}}^{\mathrm{uni}}-\mathscr{I}_{\ell_{d}-1}^{\mathrm{uni}})\quad\text{with }\mathscr{I}_{0}^{\mathrm{uni}}=0,$
where $\mathscr{I}_{\ell}^{\mathrm{uni}}$ is the polynomial interpolation operator using the zeros of $H_{n_{\ell}+1}$.
Theorem 3.3.
Suppose that a finite index set $\Lambda\subset\mathbb{N}^{d}$ is downward-closed in one direction. Let $\mathscr{I}_{\Lambda}$ be the corresponding sparse-grid approximation, and let $N$ be the total number of the point evaluations used in $\mathscr{I}_{\Lambda}$. Then, we have
$\sup_{f\neq 0}\frac{\|f-\mathscr{I}_{\Lambda}(f)\|_{L_{\rho}^{1}}}{\|f\|_{H_{\rho}^{\alpha}}}\geq c_{\alpha,d}N^{-\alpha/2}.$
For the isotropic index set $\Lambda_{L}$, under the assumptions of Theorem 3.2, we have
$\sup_{f\neq 0}\frac{\|f-\mathscr{I}_{\Lambda_{L}}(f)\|_{L_{\rho}^{1}}}{\|f\|_{H_{\rho}^{\alpha}}}\geq c_{\alpha,d}N^{-\alpha/2}(\ln N)^{\frac{a}{2}(d-1)}.$
Proof.
Without loss of generality, we assume that $\Lambda$ is downward-closed in the first coordinate. Then, for some $\overline{\ell}_{1}\in\mathbb{N}$ we have $(\ell_{1},1\dots,1)\in\Lambda$ for all $\ell_{1}=1,\dots,\overline{\ell}_{1}$. For $h_{\Lambda}$ defined in (4) it follows that
$\mathscr{I}_{\Lambda}(h_{\Lambda})$ $=\mathscr{I}_{\overline{\ell}_{1}}^{\mathrm{uni}}(p_{n}).$
Then, $\|h_{\Lambda}-\mathscr{I}_{\Lambda}(h_{\Lambda})\|_{L_{\rho}^{1}(\mathbb{R}^{d})}=\|p_{n}-\mathscr{I}_{\overline{\ell}_{1}}^{\mathrm{uni}}(p_{n})\|_{L_{\rho}^{1}(\mathbb{R})},$ and since $\mathscr{I}_{\overline{\ell}_{1}}^{\mathrm{uni}}$ is interpolatory,
$\biggl|\int_{\mathbb{R}}p_{n}(x)\rho(x)\mathrm{d}x-Q_{\overline{\ell}_{1}}^{\mathrm{uni}}(p_{n})\biggr|$ $=\biggl|\int_{\mathbb{R}}\bigl[p_{n}(x)-\mathscr{I}_{\overline{\ell}_{1}}^{\mathrm{uni}}(p_{n})(x)\bigr]\rho(x)\mathrm{d}x\biggr|$
$\leq\|p_{n}-\mathscr{I}_{\overline{\ell}_{1}}^{\mathrm{uni}}(p_{n})\|_{L_{\rho}^{1}(\mathbb{R})}.$
Hence, using $\|h_{\Lambda}\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}=\left(\sum_{r=0}^{\alpha}\|D^{r}p_{n}\|_{L_{\rho}^{2}(\mathbb{R})}^{2}\right)^{\frac{1}{2}}$, the argument reduces to bounding the worst-case error for the integration problem of a univariate algorithm:
$\frac{\bigl|\int_{\mathbb{R}}p_{n}(x)\rho(x)\mathrm{d}x-Q_{\overline{\ell}_{1}}^{\mathrm{uni}}(p_{n})\bigr|}{\left(\sum_{r=0}^{\alpha}\|D^{r}p_{n}\|_{L_{\rho}^{2}(\mathbb{R})}^{2}\right)^{\frac{1}{2}}}$ $\leq\frac{\|h_{\Lambda}-\mathscr{I}_{\Lambda}(h_{\Lambda})\|_{L_{\rho}^{1}(\mathbb{R}^{d})}}{\|h_{\Lambda}\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}}.$
Following the arguments in the proofs of Theorems 3.1 and 3.2, the statement follows. ∎
3.3 Upper bound
In this section, we use the index set
$\Lambda_{L}:=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid\,|\boldsymbol{\ell}|_{1}\leq L\}.$
As mentioned earlier, this choice corresponds to the classical sparse-grid quadrature, also known as the isotropic sparse grid or the classical Smolyak sparse grid.
To derive an upper bound, we will invoke the general theory of Smolyak algorithm [65]; see also Smolyak’s foundational work [61]. The theory therein is for general approximation problem of tensor product of continuous linear functionals defined on the tensor product of a reproducing kernel Hilbert space. Although such theory can be applied here in our setting through the norm-equivalence with the Hermite spaces, such a strategy will induce an extra constant possibly dependent on $d$. Instead, we will directly verify that an analogous result holds in $H_{\rho}^{\alpha}$. For notational convenience, define $e(S_{\Lambda}(f)):=\sup_{f\neq 0}\frac{|I(f)-S_{\Lambda}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}.$ The following result is essentially by Wasilkowski and Woźniakowski [65].
Proposition 3.4.
If the worst-case error of $Q_{n_{\ell}}^{\mathrm{uni}}$ for the univariate problem is bounded as
$\sup_{0\neq g\in H_{\rho}^{\alpha}}\frac{\bigl|\int_{\mathbb{R}}g(x)\rho(x)\mathrm{d}x-Q_{n_{\ell}}^{\mathrm{uni}}(g)\bigr|}{\|g\|_{H_{\rho}^{\alpha}}}\leq C_{1}\,M^{-\alpha\ell/2}$
for all $\ell\geq 0$ with $M>1$ independent of $\ell$, then, for the multivariate case, the worst-case error of $S_{\Lambda_{L}}$ is bounded as
$e(S_{\Lambda_{L}}(f))\leq C_{2}\,M^{-\alpha L/2}\binom{L}{d-1}$
with $C_{2}:=C_{1}\Big(\max\Big\{M^{\alpha/2},c_{\alpha}M^{\alpha/2}+c_{\alpha}M^{\alpha}\Big\}\Big)^{d-1}$.
Proof.
We closely follow the argument in [65, Proof of Lemma 2]. The same reasoning applies if the functional norm on $H_{\rho}^{\alpha}(\mathbb{R}^{d})$ is bounded by the product of functional norms on $H_{\rho}^{\alpha}(\mathbb{R})$. Since such a bound is established in Section 2, the result follows. ∎
Then, the final piece for deriving the numerical integration error is an error analysis in the space that gives $H_{\rho}^{\alpha}$ upon taking the tensor product. Such result is obtained in our previous paper [41].
Proposition 3.5.
Suppose that the Gauss–Hermite rule $Q_{\ell}^{\mathrm{uni}}$ for univariate functions uses $n_{\ell}$ quadrature points for $\ell\geq 1$. Suppose further that for some $M_{0}>0$ and $M>1$ we choose $n_{\ell}$ as
$M^{\ell-1}\leq n_{\ell}\leq M_{0}(M^{\ell}-1).$
Let $N$ be the resulting quadrature points used by $S_{\Lambda_{L}}$ with $L\geq d+1$. Then, we have
$|I(f)-S_{\Lambda_{L}}(f)|\leq c_{\alpha,d}\|f\|_{H_{\rho}^{\alpha}}(\ln N)^{(d-1)(1+\frac{\alpha}{2})}N^{-\alpha/2},$
where the constant $c_{\alpha,d}$ is independent of $f$.
Proof.
From [41, Proposition 3.4], we know that the worst-case error in the Sobolev space for univariate functions is given by
$\sup_{g\neq 0}\frac{\big|\int_{\mathbb{R}}g(x)\rho(x)\mathrm{d}x-Q_{n_{\ell}}^{\mathrm{uni}}(g)\big|}{\|g\|_{H_{\rho}^{\alpha}}}\leq c_{\alpha}n_{\ell}^{-\alpha/2}\leq C_{1}\,M^{-\alpha\ell/2}$
with $C_{1}:=c_{\alpha}M^{\alpha/2}$. This bound enables us to invoke Proposition 3.4, yielding
$e(S_{\Lambda_{L}}(f))\leq C_{2}\,M^{-\alpha L/2}\binom{L}{d-1}$
with $C_{2}:=C_{1}\Big(\max\Big\{M^{\alpha/2},c_{\alpha}M^{\alpha/2}+c_{\alpha}M^{\alpha}\Big\}\Big)^{d-1}$.
To express the right-hand side above in terms of $N$, we obtain some bounds for powers of $M$. The lower bound $M^{L-d}\leq n_{L-d+1}\leq N$ can be obtained by considering the contribution for the quadrature corresponding to the index $(\ell_{1},\dots,\ell_{d})$ with $\ell_{1}=L-d+1$ and $\ell_{2}=\dotsb=\ell_{d}=1$. On the other hand, following the calculations in [65, Section 5] we get $N\leq M_{0}^{d}\frac{M}{M-1}M^{L}\binom{L-1}{d-1}$. From $\binom{L-1}{d-1}=\big(1-\frac{d-1}{L}\big)\binom{L}{d-1}\leq\binom{L}{d-1}$
$M^{-L}\leq N^{-1}M_{0}^{d}\frac{M}{M-1}\binom{L}{d-1}$.
Hence, together with $\binom{L}{d-1}\leq\frac{L^{d-1}}{(d-1)!}$ we have
$e(S_{\Lambda}(f))$ $\leq C_{2}\,M^{-\alpha L/2}\binom{L}{d-1}\leq C_{2}\,N^{-\alpha/2}M_{0}^{d\alpha/2}\left(\frac{M}{M-1}\right)^{\alpha/2}\binom{L}{d-1}^{1+\alpha/2}$
$\leq\frac{C_{2}}{((d-1)!)^{1+\alpha/2}}\left(\frac{M_{0}^{d}M}{M-1}\right)^{\alpha/2}N^{-\alpha/2}L^{(d-1)(1+\alpha/2)}.$
[…]
$e(S_{\Lambda_{L}}(f))\leq\frac{C_{2}((1+d)/\ln M)^{(d-1)(1+\alpha/2)}}{((d-1)!)^{1+\alpha/2}}\left(\frac{M_{0}^{d}M}{M-1}\right)^{\alpha/2}\,N^{-\alpha/2}(\ln N)^{(d-1)(1+\alpha/2)}.$
Hence, the statement follows with $c_{\alpha,d}:=\frac{C_{2}((1+d)/\ln M)^{(d-1)(1+\alpha/2)}}{((d-1)!)^{1+\alpha/2}}\left(\frac{M_{0}^{d}M}{M-1}\right)^{\alpha/2}$. ∎
4 Quasi-Monte Carlo methods
A quasi-Monte Carlo method is an equal-weight integration rule. It approximates the integral of a function $g$ by
$\frac{1}{N}\sum_{j=1}^{N}g(\boldsymbol{t}^{(j)})\approx\int_{[0,1]^{d}}g(\boldsymbol{x})\mathrm{d}\boldsymbol{x}$
with nodes $\boldsymbol{t}^{(j)}\in[0,1]^{d}$, $j=1,\dots,N$. Popular classes for the choice of $\boldsymbol{t}_{j}$ include lattice points [60], higher-order digital nets [14], and low discrepancy points/sequences [52, Chapter 3]. These classes are not mutually exclusive: lattice points and digital nets may also have low discrepancy; see, for instance, [12, Chapter 5] and [14, Chapter 5].
An important point to mention is that if one views $\{\boldsymbol{t}_{j}\}$ as a low-discrepancy point set and uses the Koksma–Hlawka inequality as a framework to bound the QMC error, then the best achievable rate is linear convergence. This is because the Koksma–Hlawka inequality is based on first-order derivatives. To achieve higher-order convergence, one needs a framework that exploits higher-order derivatives.
QMC points are typically defined on the unit cube $0,1]^{d}$. To integrate functions over $\mathbb{R}^{d}$, we therefore apply domain truncation and a change of variables:
$\int_{\mathbb{R}^{d}}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}$ $=\int_{\Omega}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}+\int_{\mathbb{R}^{d}\setminus\Omega}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}$ (8)
and
$\int_{\Omega}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}$ $=\int_{[0,1]^{d}}f(\boldsymbol{\Psi}(\boldsymbol{t}))\,\rho(\boldsymbol{\Psi}(\boldsymbol{t}))\,|\det(D\boldsymbol{\Psi}(\boldsymbol{t}))|\,\mathrm{d}\boldsymbol{t}$
$\approx\frac{1}{N}\sum_{j=1}^{N}f\bigl(\boldsymbol{\Psi}(\boldsymbol{t}^{(j)})\bigr)\rho\bigl(\boldsymbol{\Psi}(\boldsymbol{t}^{(j)})\bigr)|\det(D\boldsymbol{\Psi}(\boldsymbol{t}^{(j)}))|=:Q_{N}(f).$ (9)
Here, $\Omega\subseteq\mathbb{R}^{d}$ is a suitable integration domain, possibly $\Omega=\mathbb{R}^{d}$. Often the mapping is component-wise, $\boldsymbol{\Psi}(\boldsymbol{t})=(\Psi_{1}(t_{1}),\dots,\Psi_{d}(t_{d}))$, and non-negative, in which case the Jacobian factor simplifies as
$|\det(D\boldsymbol{\Psi}(\boldsymbol{t}))|=\prod_{k=1}^{d}\Psi_{k}^{\prime}(t_{k}).$
In this paper, we consider two choices for $\Psi$: an affine map and a Möbius transformation. For the former, $\Omega$ is a suitable box; for the latter, $\Omega$ is the whole $\mathbb{R}^{d}$. We will show that both of them map $f\in H_{\rho}^{\alpha}$ to Sobolev spaces over the unit cube for which QMC methods are known to work well.
As $\{\boldsymbol{t}^{(1)},\ldots,\boldsymbol{t}^{(N)}\}$ we consider good rank-$1$ lattices and digital nets. A rank-$1$ lattice is defined by
$\boldsymbol{t}^{(j)}:=\bigg\{\frac{\boldsymbol{z}j}{N}\bigg\}\;\;\text{ for }j=1,\ldots,N,$
where $\{\cdot\}$ expresses the fractional part of each component, and the integer vector $\boldsymbol{z}\in\mathbb{N}^{d}$ is called a generating vector. A good generating vector can be found efficiently using a greedy algorithm known as the component-by-component (CBC) construction, which provably yields a vector with a satisfactorily small worst-case error in various function spaces [12].
A digital net is defined as follows. For a fixed prime $p$ and a positive integer $m$, we denote the $p$-adic expansion of a non-negative integer $i<p^{m}$ and a real number $x\in[0,1)$ by
$i=\iota_{0}+\iota_{1}p+\cdots\iota_{m-1}p^{m-1}\qquad\text{and}\qquad x=\frac{\xi_{1}}{p}+\frac{\xi_{2}}{p^{2}}+\cdots,$
respectively. We then write $\vec{i}=(\iota_{0},\iota_{1},\ldots,\iota_{m-1})^{\top}\in\mathbb{Z}_{p}^{m}$ and $\vec{x}=(\xi_{1},\xi_{2},\ldots)^{\top}\in\mathbb{Z}_{p}^{\mathbb{N}}$. Here, $\mathbb{Z}_{p}$ is the set $\{0,\dots,p-1\}$ equipped with the usual arithmetic operations modulo $p$. If $x$ admits two different $p$-adic expansions, we choose the one in which infinitely many of the digits $\xi_{i}$ differ from p1p-1.
Now, for suitably designed generating matrices $C_{1},\ldots,C_{d}\in\mathbb{Z}_{p}^{\mathbb{N}\times m}$, a digital net consisting of $N=p^{m}$ points is given by
$\vec{t}_{j,i}=C_{i}\,\overrightarrow{(j-1)}\;\;\text{ for }i=1,\ldots,d\text{ and }j=1,\ldots,N.$
Note that if every column of the generating matrices has only finitely many non-zero entries, this matrix-vector product is well-defined. The classical (first-order) digital nets constructed by Sobol’ [62], Niederreiter [50], and Niederreiter–Xing [51] are included as special cases of this construction. In [10], Dick introduced the notion of higher-order digital nets, which also form a special class within this framework. An order-$\alpha$ digital net with $\alpha\in\mathbb{N}$, $\alpha\geq 2$, is constructed using generating matrices $C_{1},\ldots,C_{d}$, which are themselves derived from the generating matrices $D_{1},\ldots,D_{\alpha d}$ for first-order digital nets in dimension $\alpha d$. We refer the reader to [28] for a review of higher-order nets.
It is worth noting that QMC methods presented in this section always have positive weights, i.e., when seeing the quadrature (9) as $Q_{N}(f)=\sum_{j=1}^{N}w_{j}f(\boldsymbol{x}_{j})$, weights $w_{j}$’s are all positive. As such, nonnegative integrands yield nonnegative approximated integral values. This nonnegativity-preserving property is important in situations where the quantity of interest must be nonnegative, for example probabilities. Positive quadrature weights are also useful when discretizing optimization problems involving expectations [46] with convex objective functions, since they preserve convexity. By contrast, other quadrature methods may involve negative weights and thus lose these desirable features.
4.1 Affinely-mapped QMC
For $\boldsymbol{b}=(b,\ldots,b)\in(0,\infty)^{d}$, define the isotropic affine transformation $\boldsymbol{\Psi}_{\mathrm{affine},b}:[0,1]^{d}\to\prod_{k=1}^{d}[-b,b]\eqqcolon[-\boldsymbol{b},\boldsymbol{b}]\subset\mathbb{R}^{d}$ by
$\boldsymbol{\Psi}_{\mathrm{affine}}:=\boldsymbol{\Psi}_{\mathrm{affine},b}(\boldsymbol{t})=2b\boldsymbol{t}-\boldsymbol{b}.$
Choosing $\boldsymbol{b},\boldsymbol{b}]$ as $\Omega$ in (8) and using $\boldsymbol{\Psi}_{\mathrm{affine}}$ in (9) yields
$Q_{N}(f)=\frac{(2b)^{d}}{N}\sum_{j=1}^{N}f\bigl(\boldsymbol{\Psi}_{\mathrm{affine}}(\boldsymbol{t}^{(j)})\bigr)\rho\bigl(\boldsymbol{\Psi}_{\mathrm{affine}}(\boldsymbol{t}^{(j)})\bigr),\quad f\in H_{\rho}^{\alpha}.$ (10)
The value $b$ controls the truncation error, namely the integral of $f\rho$ over $\mathbb{R}^{d}\setminus[-\boldsymbol{b},\boldsymbol{b}]$, whereas $Q_{N}$ controls the numerical integration error. With a suitable choice of $b$ and point set $\{\boldsymbol{t}_{1},\ldots,\boldsymbol{t}_{N}\}$, QMC methods attain the optimal convergence rate up to logarithmic factors, with the dominant part $N^{-\alpha}$ being optimal.
Proposition 4.1 (Affinely-mapped lattice).
Let $\alpha\in\mathbb{N}$, and let $Q_{N}$ be the rank-$1$ lattice rule defined in (10) with $b=\sqrt{\frac{\alpha}{2}\ln N}$. Then, with a generating vector $\boldsymbol{z}$ obtained by the CBC construction in [12, Algorithm 3.14], $Q_{N}$ satisfies
$\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}\leq C_{\alpha,d}N^{-\alpha}(\ln N)^{\frac{3\alpha d}{2}+\frac{d}{4}}.$ (11)
Proof.
First, we will show that $f\in H_{\rho}^{\alpha}$ implies $\|f\rho\|_{\text{unanchored}}+\|f\rho\|_{\text{decay}}\leq C_{d,\alpha}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$ for some constant $C_{d,\alpha}>0$, where
$\|f\rho\|_{\text{decay}}\coloneqq\sup_{\begin{subarray}{c}\boldsymbol{x}\in\mathbb{R}^{d}\\ \mathbf{r}\in\{0,\ldots,\alpha-1\}^{d}\end{subarray}}\left|\rho^{-\varepsilon}(\boldsymbol{x})\,D^{\mathbf{r}}(f\rho)(\boldsymbol{x})\right|<\infty$
for all $\varepsilon\in(0,1/2)$, and
$\|f\rho\|_{\text{unanchored}}\coloneqq\sup_{[\boldsymbol{a},\boldsymbol{b}]\subset\mathbb{R}^{d}}\|f\rho\|_{\mathscr{H}^{\alpha}([\boldsymbol{a},\boldsymbol{b}])}<\infty$
with
$\langle f,h\rangle_{\mathscr{H}^{\alpha}([\boldsymbol{a},\boldsymbol{b}])}$ $:=\sum_{\begin{subarray}{c}\mathbf{r}\in\{0,\dots,\alpha\}^{d}\\ \mathfrak{v}:=\{k:r_{k}=\alpha\}\end{subarray}}\int_{[\boldsymbol{a}_{\mathfrak{v}},\boldsymbol{b}_{\mathfrak{v}}]}\left(\int_{[\boldsymbol{a}_{-\mathfrak{v}},\boldsymbol{b}_{-\mathfrak{v}}]}D^{\mathbf{r}}f(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)$ (12)
$\qquad\qquad\qquad\qquad\times\left(\int_{[\boldsymbol{a}_{-\mathfrak{v}},\boldsymbol{b}_{-\mathfrak{v}}]}D^{\mathbf{r}}h(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)\mathrm{d}\boldsymbol{x}_{\mathfrak{v}}$
for bounded $\boldsymbol{a},\boldsymbol{b}]$. In the proof we use the following identities
$D^{\mathbf{r}}(H_{\mathbf{k}}\,\rho)$ $=(-1)^{|\mathbf{r}|}\sqrt{\frac{(\mathbf{k}+\mathbf{r})!}{\mathbf{k}!}}H_{\mathbf{k}+\mathbf{r}}\,\rho,\ \text{ and }\ D^{\mathbf{r}}\rho=(-1)^{|\mathbf{r}|}\sqrt{\mathbf{r}!}\,H_{\mathbf{r}}\,\rho,$
which can be derived from the definition of Hermite polynomial.
For $g=f\rho$ and $\mathbf{r}\in\mathbb{N}_{0}^{d}$ with $|\mathbf{r}|_{\infty}\leq\alpha$, the product rule yields
$D^{\mathbf{r}}g$ $=\sum_{\mathbf{s}\leq\mathbf{r}}\binom{\mathbf{r}}{\mathbf{s}}(D^{\mathbf{s}}\rho)\,(D^{\mathbf{r}-\mathbf{s}}f)=\sum_{\mathbf{s}\leq\mathbf{r}}(-1)^{|\mathbf{s}|}\sqrt{\mathbf{s}!}\binom{\mathbf{r}}{\mathbf{s}}\,H_{\mathbf{s}}\,\rho\,D^{\mathbf{r}-\mathbf{s}}f.$ (13)
Now,
$\sup_{[\boldsymbol{a},\boldsymbol{b}]\subset\mathbb{R}^{d}}\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{a},\boldsymbol{b}])}^{2}\leq\sum_{\begin{subarray}{c}\mathbf{r}\in\{0,\ldots,\alpha\}^{d}\\ \mathfrak{v}:=\left\{j:r_{j}=\alpha\right\}\end{subarray}}\int_{\mathbb{R}^{|\mathfrak{v}|}}\left(\int_{\mathbb{R}^{d-|\mathfrak{\mathfrak{v}}|}}|D^{\mathbf{r}}g(\boldsymbol{x})|\,\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)^{2}\mathrm{d}\boldsymbol{x}_{\mathfrak{v}}.$ (14)
Using the Cauchy–Schwarz inequality, the inner-integral can be bounded as
$\int_{\mathbb{R}^{d-|\mathfrak{\mathfrak{v}}|}}|D^{\mathbf{r}}g(\boldsymbol{x})|\,\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\leq\sum_{\mathbf{s}\leq\mathbf{r}}\sqrt{\mathbf{s}!}\binom{\mathbf{r}}{\mathbf{s}}\int_{\mathbb{R}^{d-|\mathfrak{\mathfrak{v}}|}}|H_{\mathbf{s}}(\boldsymbol{x})D^{\mathbf{r}-\mathbf{s}}f(\boldsymbol{x})|\,\rho(\boldsymbol{x})\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}$
$\ \ \kern-8.53581pt\leq\sum_{\mathbf{s}\leq\mathbf{r}}\sqrt{\mathbf{s}!}\binom{\mathbf{r}}{\mathbf{s}}|H_{\boldsymbol{s}(\mathfrak{v})}(\boldsymbol{x}_{\mathfrak{v}})|\kern 1.5pt\rho_{\mathfrak{\mathfrak{v}}}(\boldsymbol{x}_{\mathfrak{v}})\kern 1.0pt{\|H_{\boldsymbol{s}(-\mathfrak{v})}\|_{L^{2}_{\rho_{\mathfrak{-\mathfrak{v}}}}}\kern-2.84526pt}\left(\int_{\mathbb{R}^{d-|\mathfrak{\mathfrak{v}}|}}\kern-5.69054pt|D^{\mathbf{r}-\mathbf{s}}f(\boldsymbol{x})|^{2}\,\rho_{\mathfrak{-\mathfrak{v}}}(\boldsymbol{x}_{-\mathfrak{v}})\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)^{\kern-2.84526pt1/2}$
with $\|H_{\boldsymbol{s}(-\mathfrak{v})}\|_{L^{2}_{\rho_{\mathfrak{-\mathfrak{v}}}}}:=\left(\int_{\mathbb{R}^{d-|\mathfrak{\mathfrak{v}}|}}|H_{\boldsymbol{s}(-\mathfrak{v})}(\boldsymbol{x}_{-\mathfrak{v}})|^{2}\,\rho_{\mathfrak{-\mathfrak{v}}}(\boldsymbol{x}_{-\mathfrak{v}})\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)^{1/2}$. Here, we used the notations $H_{\boldsymbol{s}(\mathfrak{v})}(x_{\mathfrak{v}}):=\prod_{j\in\mathfrak{v}}H_{s_{j}}(x_{j})$, $H_{\boldsymbol{s}(-\mathfrak{v})}(x_{-\mathfrak{v}}):=\prod_{j\in\{1,\dots,d\}\setminus\mathfrak{v}}H_{s_{j}}(x_{j})$, and similarly $\rho_{\mathfrak{v}}:=\prod_{j\in\mathfrak{v}}\rho(x_{j})$ and $\rho_{-\mathfrak{v}}:=\prod_{j\in\{1,\dots,d\}\setminus\mathfrak{v}}\rho(x_{j})$. Noting $\|H_{\boldsymbol{s}(-\mathfrak{v})}\|_{L^{2}_{\rho_{\mathfrak{-\mathfrak{v}}}}}=1$, it follows
$\int_{\mathbb{R}^{|\mathfrak{\mathfrak{v}}|}}$ $\left(\int_{\mathbb{R}^{d-|\mathfrak{\mathfrak{v}}|}}|D^{\mathbf{r}}g(\boldsymbol{x})|\,\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)^{2}\mathrm{d}\boldsymbol{x}_{\mathfrak{v}}$
$\leq C_{d}\sum_{\mathbf{s}\leq\mathbf{r}}\mathbf{s}!\binom{\mathbf{r}}{\mathbf{s}}^{2}\int_{\mathbb{R}^{d}}|H_{\boldsymbol{s}(\mathfrak{v})}(\boldsymbol{x}_{\mathfrak{v}})|^{2}\rho_{\mathfrak{\mathfrak{v}}}^{2}(\boldsymbol{x}_{\mathfrak{v}})\,|D^{\mathbf{r}-\mathbf{s}}f(\boldsymbol{x})|^{2}\,\rho_{\mathfrak{-\mathfrak{v}}}(\boldsymbol{x}_{-\mathfrak{v}})\mathrm{d}\boldsymbol{x}$
$\leq C_{d}\sum_{\mathbf{s}\leq\mathbf{r}}\mathbf{s}!\binom{\mathbf{r}}{\mathbf{s}}^{2}\sup_{\boldsymbol{y}_{\mathfrak{v}}\in\mathbb{R}^{|\mathfrak{v}|}}|H_{\boldsymbol{s}(\mathfrak{v})}(\boldsymbol{y}_{\mathfrak{v}})|^{2}\rho_{\mathfrak{\mathfrak{v}}}(\boldsymbol{y}_{\mathfrak{v}})\int_{\mathbb{R}^{d}}|D^{\mathbf{r}-\mathbf{s}}f(\boldsymbol{x})|^{2}\,\rho(\boldsymbol{x})\mathrm{d}\boldsymbol{x}.$
Therefore, from (14)
$\sup_{[\boldsymbol{a},\boldsymbol{b}]\subset\mathbb{R}^{d}}\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{a},\boldsymbol{b}])}\leq C_{d,\alpha}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$
holds for some constant $C_{d,\alpha}>0$.
To show $\lVert g\rVert_{\text{decay}}\leq C\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$, we will invoke the Sobolev embedding in Proposition 2.3 to bound the supremum norm. Define $h_{\mathbf{r}}:=\rho^{-\varepsilon}D^{\mathbf{r}}g=\rho^{-\varepsilon}D^{\mathbf{r}}(f\rho)$ for $\mathbf{r}\in\{0,\dots,\alpha-1\}^{d}$, for which we will derive a bound on its Sobolev norm. For $\mathbf{k}\in\{0,1\}^{d}$
$D^{\mathbf{k}}h_{\mathbf{r}}(\boldsymbol{x})=\sum_{\mathbf{j}\leq\mathbf{k}}\binom{\mathbf{k}}{\mathbf{j}}\biggl(\prod_{j\in\mathbf{j}}{\varepsilon\,x_{j}}\biggr)\rho^{-\varepsilon}(\boldsymbol{x})(D^{\mathbf{k}-\mathbf{j}+\mathbf{r}}g(\boldsymbol{x}))$
holds. With $p_{\mathbf{j}}(\boldsymbol{x}):=\prod_{j\in\mathbf{j}}\varepsilon\,x_{j}$, taking the $L^{2}(\mathbb{R}^{d})$-norm of both sides together with triangular inequality yields
$\|D^{\mathbf{k}}h_{\mathbf{r}}\|_{L^{2}(\mathbb{R}^{d})}$ $\leq\sum_{\mathbf{j}\leq\mathbf{k}}\binom{\mathbf{k}}{\mathbf{j}}\sum_{\mathbf{s}\leq\mathbf{r}+\mathbf{k}-\mathbf{j}}\sqrt{\mathbf{s}!}\binom{\mathbf{r}+\mathbf{k}-\mathbf{j}}{\mathbf{s}}\|p_{\mathbf{j}}\rho^{-\varepsilon}\,H_{\mathbf{s}}\,\rho\,D^{\mathbf{r}+\mathbf{k}-\mathbf{j}-\mathbf{s}}f\|_{L^{2}(\mathbb{R}^{d})}$
$\leq\sum_{\mathbf{j}\leq\mathbf{k}}\binom{\mathbf{k}}{\mathbf{j}}\sum_{\mathbf{s}\leq\mathbf{r}+\mathbf{k}-\mathbf{j}}\sqrt{\mathbf{s}!}\binom{\mathbf{r}+\mathbf{k}-\mathbf{j}}{\mathbf{s}}\sup_{\boldsymbol{y}\in\mathbb{R}^{d}}|p_{\mathbf{j}}(\boldsymbol{y})H_{\mathbf{s}}(\boldsymbol{y})\rho^{1/2-\varepsilon}(\boldsymbol{y})|$
$\qquad\qquad\times\|\,\,\sqrt{\rho}\,(D^{\mathbf{r}+\mathbf{k}-\mathbf{j}-\mathbf{s}}f)\|_{L^{2}(\mathbb{R}^{d})}$
$=\sum_{\mathbf{j}\leq\mathbf{k}}\binom{\mathbf{k}}{\mathbf{j}}\sum_{\mathbf{s}\leq\mathbf{r}+\mathbf{k}-\mathbf{j}}\sqrt{\mathbf{s}!}\binom{\mathbf{r}+\mathbf{k}-\mathbf{j}}{\mathbf{s}}\sup_{\boldsymbol{y}\in\mathbb{R}^{d}}|p_{\mathbf{j}}(\boldsymbol{y})H_{\mathbf{s}}(\boldsymbol{y})\rho^{1/2-\varepsilon}(\boldsymbol{y})|$
$\qquad\qquad\times\|D^{\mathbf{r}+\mathbf{k}-\mathbf{j}-\mathbf{s}}f\|_{L_{\rho}^{2}(\mathbb{R}^{d})}<\infty,$ (15)
and we used the identity (13) for $D^{\mathbf{k}-\mathbf{j}+\mathbf{r}}g$. In (15), the multi-index satisfies $\mathbf{r}+\mathbf{k}-\mathbf{j}-\mathbf{s}\in\{0,\dots,\alpha\}^{d}$, and thus
$\sum_{\mathbf{k}\in\{0,1\}^{d}}\|D^{\mathbf{k}}(\rho^{-\varepsilon}D^{\mathbf{r}}g)\|_{L^{2}(\mathbb{R}^{d})}^{2}=\sum_{\mathbf{k}\in\{0,1\}^{d}}\|D^{\mathbf{k}}h_{\mathbf{r}}\|_{L^{2}(\mathbb{R}^{d})}^{2}$ $\leq C\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}^{2}.$
Therefore, in view of Proposition 2.3 we have $\sup_{\boldsymbol{x}\in\mathbb{R}^{d}}|\rho^{-\varepsilon}(\boldsymbol{x})D^{\mathbf{r}}g(\boldsymbol{x})|\leq C\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$, hence $\|f\rho\|_{\text{unanchored}}+\|f\rho\|_{\text{decay}}\leq C_{d,\alpha}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$ follows.
With this, from the first displayed equation in [55, Lemma 5] together with [55, Propositions 7 & 8], if $\boldsymbol{z}$ satisfies the second displayed equation in [55, Lemma 5], then this $\boldsymbol{z}$ satisfies (11); see the proof of [55, Theorem 2]. Since such $\boldsymbol{z}$ can be constructed using a CBC construction in [12, Algorithm 3.14], the proof is complete. ∎
Affinely mapped digital nets also achieve the optimal convergence rate, up to a logarithmic factor, as shown in [11, Corollary 1].
Proposition 4.2 (Affinely-mapped digital net, [11, Corollary 1]).
Let $\alpha\in\mathbb{N}$, and let $Q_{N}$ be the quadrature rule defined in (10) with $b=2\sqrt{{\alpha}\ln N}$, where points $(\boldsymbol{t}^{(j)})_{j=1,\dots,N}$ are given by a higher-order net of order $(2\alpha+1)$ with $N=p^{m}$ elements. Then, the worst case error is bounded by
$\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}\leq C_{\alpha,d}N^{-\alpha}(\ln N)^{\frac{\alpha d}{2}+\frac{3d}{4}-\frac{1}{2}}.$
In the result quoted above, higher-order nets of order $(2\alpha+1)$ are considered. On the other hand, as noted in [11], higher-order nets of order $\alpha$ also achieve the optimal polynomial rate $N^{-\alpha}$, but with a logarithmic factor of a larger exponent.
4.2 QMC with Möbius transformation
An alternative to the affine transformation is the change of variables $\Psi:(0,1)^{d}\to\mathbb{R}^{d}$ defined by
$\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\coloneqq(\phi(t_{1}),\ldots,\phi(t_{d})),\;\;\;\phi(t)\coloneqq-\cot(\pi t).$ (16)
This is a specific Möbius transformation considered in [64]. For the univariate transformation $\phi$, its derivatives satisfy the following.
Lemma 4.3.
For every $\tau\in\mathbb{N}$, $\phi^{(\tau)}\circ\phi^{-1}(x)$ is a polynomial in $x$ of degree $\tau+1$.
Proof.
Noting $\phi^{(\tau)}(t)=-\pi^{\tau}\cot^{(\tau)}(\pi t)$ and $\phi^{-1}(x)=\frac{1}{\pi}\cot^{-1}(-x)$, it suffices to show that $\cot^{(\tau)}\circ\cot^{-1}(x)$ is a polynomial of degree $\tau+1$. From [34, Equation (5)] we know that $P_{\tau}(x):=\tan^{(\tau)}\circ\tan^{-1}(x)$ is a polynomial of degree $\tau+1$ satisfying
$P_{\tau+1}=\sum_{r=0}^{\tau}\binom{\tau}{r}P_{\tau-r}(x)P_{r}(x),$
with $P_{0}(x)=x$ and $P_{1}(x)=x^{2}+1$. With $\cot(t)=\tan(\pi/2-t)$, the statement follows. ∎
Using $\boldsymbol{\Psi}_{\mathrm{cot}}$ in (9) yields the following formula:
$Q_{N}(f)=\frac{1}{N}\sum_{j=1}^{N}f\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t}^{(j)})\right)\rho\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t}^{(j)})\right)\prod_{k=1}^{d}\phi^{\prime}(t^{(j)}_{k}),\quad f\in H_{\rho}^{\alpha}.$ (17)
We derive a higher-order chain rule for component-wise composition, which may be of interest by itself. Although this is a special case of the general Faa di Bruno formula, component-wise composition makes the formula simpler and thus easier to work with for broader context in numerical analysis.
First, we derive a recurrence relation in the univariate setting. For $r\geq 1$ and $1\leq k\leq r$, let $B_{r,k}$ denote the (partial) Bell polynomial. We use the convention $B_{0,0}=1$ and $B_{r,0}=0$ for $r\geq 1$.
Define $C_{0,0}:=B_{0,0}=1$ and $C_{r,-1}:=0$ for all $r\in\mathbb{N}_{0}$. For $r\in\mathbb{N}$, $C_{r,0}:=B_{r,0}=0$, and for integers $1\leq\eta\leq r$, define
$C_{r,\eta}(t):=B_{r,\eta}\bigl(\Phi^{\prime}(t),\Phi^{\prime\prime}(t),\ldots,\Phi^{(r-\eta+1)}(t)\bigr).$
The function $C_{r,\eta}$ thus defined satisfies the following relation.
Lemma 4.4.
Let $\Phi\colon(0,1)\to\mathbb{R}$ be differentiable. For integers $0\leq\eta\leq r$, we have
$C_{r+1,\eta}(t)=\Phi^{\prime}(t)\,C_{r,\eta-1}(t)+C^{\prime}_{r,\eta}(t),\quad\text{for all }t\in(0,1).$
Proof.
The cases $r=0$ as well as $r\in\mathbb{N}$ and $\eta=0$ are easy to check. We show the statement for $1\leq\eta\leq r$. First, the partial Bell polynomials satisfy the recurrence relation
$C_{r+1,\eta}(t)$ $=B_{r+1,\eta}\bigl(\Phi^{\prime}(t),\ldots,\Phi^{(r-\eta+2)}(t)\bigr)$
$=\sum_{j=0}^{r-\eta+1}\binom{r}{j}\Phi^{(j+1)}(t)\;B_{r-j,\eta-1}\bigl(\Phi^{\prime}(t),\ldots,\Phi^{(r-j-\eta+2)}(t)\bigr).$
They also satisfy another relation
$\partial_{j}B_{r,\eta}(x_{1},\ldots,x_{r-\eta+1})$ $=\binom{r}{j}B_{r-j,\eta-1}(x_{1},\ldots,x_{r-j-(\eta-1)+1}),\quad j=1,\dots,r-\eta+1,$
and thus differentiating $C_{r,\eta}$ gives
$C^{\prime}_{r,\eta}(t)$ $=\sum_{j=1}^{r-\eta+1}\partial_{j}B_{r,\eta}\Big|_{\begin{subarray}{c}x_{m}=\Phi^{(m)}(t)\\ m=1,\dots,r-\eta+1\end{subarray}}\Phi^{(j+1)}(t)$
$=\sum_{j=1}^{r-\eta+1}\binom{r}{j}B_{r-j,\eta-1}\bigl(\Phi^{\prime}(t),\ldots,\Phi^{(r-j-\eta+2)}(t)\bigr)\;\Phi^{(j+1)}(t).$
Hence, subtracting $C_{r+1,\eta}(t)$ from $C^{\prime}_{r,\eta}(t)$ completes the proof. ∎
For $\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}$ define
$C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t}):=\prod_{j=1}^{d}C_{j;r_{j},\eta_{j}}(t_{j}).$
We will prove the higher-order chain rule using the following analogous formula for smooth functions.
Lemma 4.5.
Let $v\in C^{\infty}(\mathbb{R}^{d})$. For $\Phi_{j}\in C^{\alpha}((0,1))$, with $\alpha\in\mathbb{N}$, $1\leq j\leq d$, define $\boldsymbol{\Phi}(\boldsymbol{t})=\big(\Phi_{1}(t_{1}),\dots,\Phi_{d}(t_{d})\big)$ for $\boldsymbol{t}=(t_{1},\dots,t_{d})\in(0,1)^{d}$. For every multi-index $\mathbf{r}\in\mathbb{N}_{0}^{d}$ with $|\mathbf{r}|_{\infty}\leq\alpha$
$D^{\mathbf{r}}(v\circ\boldsymbol{\Phi})(\boldsymbol{t})=\sum_{\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}}(D^{\boldsymbol{\eta}}v\circ\boldsymbol{\Phi})(\boldsymbol{t})\;C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})$ (18)
holds for every $\boldsymbol{t}\in(0,1)^{d}$, where $\boldsymbol{\eta}\leq\mathbf{r}$ means $\eta_{j}\leq r_{j}$ for all $j=1,\dots,d$.
Proof.
We use induction on $|\mathbf{r}|=r_{1}+\dots+r_{d}$. The case $\mathbf{r}=\mathbf{0}$ is trivial. Assume that (18) holds for all $|\mathbf{r}|\leq\beta$ with $|\mathbf{r}|_{\infty}\leq\alpha$, and let $|\mathbf{r}|=\beta$ with $|\mathbf{r}|_{\infty}\leq\alpha$. If $\beta=d\alpha$, then there is nothing to show. For $\beta<d\alpha$, fix $j\in\{1,\dots,d\}$ such that $r_{j}\leq\alpha-1$ and differentiate (18) with respect to $t_{j}$:
$D^{\mathbf{r}+\mathbf{e}_{j}}(v\circ\boldsymbol{\Phi})(\boldsymbol{t})=\sum_{\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}}\partial_{{}_{j}}\bigl(v_{\boldsymbol{\eta}}\circ\boldsymbol{\Phi}\bigr)(\boldsymbol{t})C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})+\sum_{\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}}v_{\boldsymbol{\eta}}(\boldsymbol{\Phi}(\boldsymbol{t}))\,\partial_{{}_{j}}C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t}),$
where we let $v_{\boldsymbol{\eta}}:=D^{\boldsymbol{\eta}}v$ for notational ease.
We now re-index the first sum so that the summation over $\boldsymbol{\eta}$ on the right-hand side runs up to $\mathbf{r}+\mathbf{e}_{j}$. Since only the $j$-th component of $\boldsymbol{\Phi}$ depends on $t_{j}$, we have $\partial_{{}_{j}}\bigl(v_{\boldsymbol{\eta}}\circ\boldsymbol{\Phi}\bigr)(\boldsymbol{t})=v_{\boldsymbol{\eta}+\mathbf{e}_{j}}(\boldsymbol{\Phi}(\boldsymbol{t}))\,\Phi_{j}^{\prime}(t_{j})$. Moreover, $\partial_{j}C_{\mathbf{r},\boldsymbol{\eta}}=\bigl(\partial_{j}C_{j;r_{j},\eta_{j}}\bigr)\prod_{k\neq j}C_{k;r_{k},\eta_{k}}=0$ if $\eta_{j}=0$. Combining these, we obtain
$D^{\mathbf{r}+\mathbf{e}_{j}}(v\circ\boldsymbol{\Phi})(\boldsymbol{t})=\sum_{\begin{subarray}{c}\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}+\mathbf{e}_{j}\\ \eta_{j}\geq 1\end{subarray}}v_{\boldsymbol{\eta}}(\boldsymbol{\Phi}(\boldsymbol{t}))\,\Phi_{j}^{\prime}(t_{j})C_{\mathbf{r},\boldsymbol{\eta}-\mathbf{e}_{j}}(\boldsymbol{t})+\sum_{\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}}v_{\boldsymbol{\eta}}(\boldsymbol{\Phi}(\boldsymbol{t}))\,\partial_{{}_{j}}C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})$
$\!=\!\!\sum_{\begin{subarray}{c}\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}+\mathbf{e}_{j}\\ \eta_{j}=r_{j}+1\end{subarray}}\!\!v_{\boldsymbol{\eta}}(\boldsymbol{\Phi}(\boldsymbol{t}))\,\Phi_{j}^{\prime}(t_{j})C_{\mathbf{r},\boldsymbol{\eta}-\mathbf{e}_{j}}(\boldsymbol{t})+\!\sum_{\begin{subarray}{c}\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}\\ \eta_{j}\geq 1\end{subarray}}\!v_{\boldsymbol{\eta}}(\boldsymbol{\Phi}(\boldsymbol{t}))\bigl[\Phi_{j}^{\prime}(t_{j})C_{\mathbf{r},\boldsymbol{\eta}-\mathbf{e}_{j}}(\boldsymbol{t})+\partial_{{}_{j}}C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})\bigr].$
We further rewrite the right-hand side. For the first term, from $B_{r+1,r+1}(x)=x^{r+1}=x\,B_{r,r}(x)$ for $r\geq 0$ and $\eta_{j}-1=r_{j}$,
$\Phi_{j}^{\prime}(t_{j})C_{\mathbf{r},\boldsymbol{\eta}-\mathbf{e}_{j}}(\boldsymbol{t})$ $=\Phi_{j}^{\prime}(t_{j})C_{j;r_{j},r_{j}}(t_{j})\prod_{k\neq j}C_{k;r_{k},\eta_{k}}(t_{k})$
$=C_{j;r_{j}+1,r_{j}+1}(t_{j})\prod_{k\neq j}C_{k;r_{k},\eta_{k}}(t_{k})=C_{\mathbf{r}+\mathbf{e}_{j},\boldsymbol{\eta}}(\boldsymbol{t})$
holds; for the second term, we use Lemma 4.4 to see
$\Phi_{j}^{\prime}(t_{j})C_{\mathbf{r},\boldsymbol{\eta}-\mathbf{e}_{j}}(\boldsymbol{t})+\partial_{{}_{j}}C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})$ $=\Bigl(\Phi_{j}^{\prime}(t_{j})C_{j;r_{j},\eta_{j}-1}(t_{j})+\partial_{j}C_{j;r_{j},\eta_{j}}(t_{j})\Bigr)\prod_{k\neq j}C_{k;r_{k},\eta_{k}}(t_{k})$
$=\Bigl(C_{j;r_{j}+1,\eta_{j}}(t_{j})\Bigr)\prod_{k\neq j}C_{k;r_{k},\eta_{k}}(t_{k})=C_{\mathbf{r}+\mathbf{e}_{j},\boldsymbol{\eta}}(\boldsymbol{t}).$
Hence,
$D^{\mathbf{r}+\mathbf{e}_{j}}(v\circ\boldsymbol{\Phi})(\boldsymbol{t})$ $=\sum_{\begin{subarray}{c}\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}+\mathbf{e}_{j}\\ \eta_{j}=r_{j}+1\end{subarray}}v_{\boldsymbol{\eta}}(\boldsymbol{\Phi}(\boldsymbol{t}))\,C_{\mathbf{r}+\mathbf{e}_{j},\boldsymbol{\eta}}(\boldsymbol{t})+\sum_{\begin{subarray}{c}\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}\\ \eta_{j}\geq 1\end{subarray}}v_{\boldsymbol{\eta}}(\boldsymbol{\Phi}(\boldsymbol{t}))\,C_{\mathbf{r}+\mathbf{e}_{j},\boldsymbol{\eta}}(\mathbf{t})$
$=\sum_{\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}+\mathbf{e}_{j}}v_{\boldsymbol{\eta}}(\boldsymbol{\Phi}(\boldsymbol{t}))\,C_{\mathbf{r}+\mathbf{e}_{j},\boldsymbol{\eta}}(\boldsymbol{t}),$
which is (18) for $\mathbf{r}+\mathbf{e}_{j}$. This completes the proof. ∎
Proposition 4.6.
For a multi-index $\mathbf{r}\in\mathbb{N}_{0}^{d}$ with $|\mathbf{r}|_{\infty}\leq\alpha$, suppose that $v\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{d})$ admits weak derivatives $D^{\mathbf{\boldsymbol{\eta}}}v\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{d})$ for $\boldsymbol{\eta}\leq\mathbf{r}$, i.e., $\eta_{j}\leq r_{j}$ for $1\leq j\leq d$. Let $\Phi_{j}\in C^{\alpha}((0,1))$ be invertible with $\Phi_{j}^{-1}\in C^{1}((0,1))$ for $1\leq j\leq d$. Define $\boldsymbol{\Phi}(\boldsymbol{t})=\big(\Phi_{1}(t_{1}),\dots,\Phi_{d}(t_{d})\big)$ for $\boldsymbol{t}=(t_{1},\dots,t_{d})\in(0,1)^{d}$. Then,
$D^{\mathbf{r}}(v\circ\boldsymbol{\Phi})(\boldsymbol{t})=\sum_{\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}}(D^{\boldsymbol{\eta}}v\circ\boldsymbol{\Phi})(\boldsymbol{t})\;C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})$ (19)
holds for a.e. $\boldsymbol{t}\in(0,1)^{d}$.
Proof.
It suffices to show that the right-hand side of (19) is locally integrable and that
$\int_{(0,1)^{d}}(v\circ\boldsymbol{\Phi})(\boldsymbol{t})\,D^{\mathbf{r}}\varphi(\boldsymbol{t})\mathrm{d}\boldsymbol{t}=(-1)^{|\mathbf{r}|}\int_{(0,1)^{d}}\left(\sum_{\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}}(D^{\boldsymbol{\eta}}v\circ\boldsymbol{\Phi})(\boldsymbol{t})\;C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})\right)\varphi(\boldsymbol{t})\mathrm{d}\boldsymbol{t}$
holds for every compactly supported $\varphi$.
Local integrability follows from the boundedness of $(\Phi_{j}^{-1})^{\prime}$ and the derivatives $\Phi_{j}^{(\tau)}$ on compact subsets of $(0,1)^{d}$, as well as from the facts that $D^{\boldsymbol{\eta}}v$ is locally integrable on $\mathbb{R}^{d}$, and that the image of each compact subset of $(0,1)^{d}$ under $\boldsymbol{\Phi}$ is compact in $\mathbb{R}^{d}$. To show the equality above, fix $\varphi$ and take an open set $U$ such that $\mathrm{supp}(\varphi)\subset U\Subset(0,1)^{d}$. Noting that $(\Phi_{j}^{-1})^{\prime}$ and $\Phi_{j}^{(\tau)}$ are bounded on $U$, we invoke Lemma 4.5 together with a standard argument for change-of-variables for Sobolev functions to conclude that the equality holds on $U$, and thus equivalently on $(0,1)^{d}$. ∎
Note that, unlike in the common setting for Sobolev spaces, we do not assume boundedness of the derivatives of $\boldsymbol{\Phi}$, since the derivatives of the change of variables $\boldsymbol{\Phi}_{\mathrm{cot}}$ we consider are unbounded. As a result, and in contrast to standard results, the integrability of $D^{\mathbf{r}}(v\circ\boldsymbol{\Phi})$ does not follow immediately, and the membership of $(f\cdot\rho)\circ\boldsymbol{\Phi}_{\mathrm{cot}}$ in the Sobolev space is handled separately in the next proposition.
Proposition 4.7.
Define $g(\boldsymbol{t}):=f(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t}))\rho(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t}))\prod_{k=1}^{d}\phi^{\prime}(t_{k})$ for $f\in H_{\rho}^{\alpha}$ with $\alpha\in\mathbb{N}$. We have the following: for all multi-indices $\mathbf{r}\in\{0,1,\ldots,\alpha\}^{d}$
$\|D^{\mathbf{r}}g\|_{L^{2}([0,1]^{d})}\leq C\|f\|_{H_{\rho}^{\alpha}},$ (20)
where $C>0$ is a constant independent of $\mathbf{r}$ and $f$; for every $\ell=1,\dots,d$, every $\mathbf{r}\in\{0,\ldots,\alpha\}^{d}$ such that $r_{\ell}\leq\alpha-1$, and almost every $(t_{1},\dots,t_{\ell-1},t_{\ell+1},\dots,t_{d})\in(0,1)^{d-1}$,
$\lim_{t_{\ell}\to 0+}D^{\mathbf{r}}g(\boldsymbol{t})=\lim_{t_{\ell}\to 1-}D^{\mathbf{r}}g(\boldsymbol{t})=0,$ (21)
and as a consequence, the map $t_{\ell}\mapsto D^{\mathbf{r}}g(\boldsymbol{t})$ extends continuously to $[0,1]$ by zero. Furthermore, for $\mathbf{r}\in\{0,\ldots,\alpha-1\}^{d}$, (21) holds for every $(t_{1},\dots,t_{\ell-1},t_{\ell+1},\dots,t_{d})\in(0,1)^{d-1}$.
Proof.
We first prove (20) for all $\mathbf{r}\in\{0,1,\ldots,\alpha\}^{d}$. By Proposition 4.6, we have
$D^{\mathbf{s}}\bigl((f\cdot\rho)\circ\boldsymbol{\Psi}_{\mathrm{cot}}\bigr)(\boldsymbol{t})=\sum_{\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{s}}\bigl(D^{\boldsymbol{\eta}}(f\cdot\rho)\circ\boldsymbol{\Psi}_{\mathrm{cot}}\bigr)(\boldsymbol{t})\;C_{\mathbf{s},\boldsymbol{\eta}}(\boldsymbol{t}).$
for all $\boldsymbol{0}\leq\mathbf{s}\leq\mathbf{r}$. Together with the product rule, we get
$[D^{\mathbf{r}}g](\boldsymbol{t})$ $=\sum_{\mathbf{s}\leq\mathbf{r}}\binom{\mathbf{r}}{\mathbf{s}}D^{\mathbf{s}}\bigl((f\cdot\rho)\circ\boldsymbol{\Psi}_{\mathrm{cot}}\bigr)(\boldsymbol{t})\,\biggl(D^{\mathbf{r}-\mathbf{s}}\prod_{k=1}^{d}\phi^{\prime}(t_{k})\biggr)$
$=\sum_{\mathbf{s}\leq\mathbf{r}}\binom{\mathbf{r}}{\mathbf{s}}\bigg(\sum_{\boldsymbol{\eta}\leq\mathbf{s}}\bigl(D^{\boldsymbol{\eta}}(f\cdot\rho)\circ\boldsymbol{\Psi}_{\mathrm{cot}}\bigr)(\boldsymbol{t})\;C_{\mathbf{s},\boldsymbol{\eta}}(\boldsymbol{t})\bigg)\,\biggl(D^{\mathbf{r}-\mathbf{s}}\prod_{k=1}^{d}\phi^{\prime}(t_{k})\biggr),$
the right-hand side of which is a linear combination of
$w(\boldsymbol{t}):=\bigg(\bigl([D^{\boldsymbol{\lambda}_{1}}f]\circ\boldsymbol{\Psi}_{\mathrm{cot}}\bigr)(\boldsymbol{t})\bigg)\bigg(\bigl([D^{\boldsymbol{\lambda}_{2}}\rho]\circ\boldsymbol{\Psi}_{\mathrm{cot}}\bigr)(\boldsymbol{t})\bigg)\prod_{k=1}^{d}\prod_{\tau=1}^{\lambda_{3,k}}\big(D^{\tau}\phi(t_{k})\big)^{\ell_{k,\tau}}$
with some $\boldsymbol{\lambda}_{1},\boldsymbol{\lambda}_{2}\leq\mathbf{r}$, $\boldsymbol{\lambda}_{3}=(\lambda_{3,1},\ldots,\lambda_{3,d})\leq\mathbf{r}+\boldsymbol{1}$, and $\ell_{k,\tau}\in\mathbb{N}_{0}$. The $L^{2}([0,1]^{d})$ norm of each term is then bounded by
$\|w\|_{L^{2}([0,1]^{d})}^{2}=\int_{\mathbb{R}^{d}}w^{2}(\boldsymbol{\Psi}_{\mathrm{cot}}^{-1}(\boldsymbol{x}))\bigg(\prod_{k=1}^{d}(\phi^{-1})^{\prime}(x_{k})\bigg)\mathrm{d}\boldsymbol{x}$
$=\int_{\mathbb{R}^{d}}\biggl[\Bigl(D^{\boldsymbol{\lambda}_{1}}\!\!\;f(\boldsymbol{x})\Bigr)\Bigl(D^{\boldsymbol{\lambda}_{2}}\!\!\;\rho(\boldsymbol{x})\Bigr)\biggr]^{2}\bigg(\prod_{k=1}^{d}\bigg(\prod_{\tau=1}^{\lambda_{3,k}}\big([D^{\tau}\phi]\circ\phi^{-1}(x_{k})\big)^{2\ell_{k,\tau}}\bigg)(\phi^{-1})^{\prime}(x_{k})\bigg)\mathrm{d}\boldsymbol{x}$
$\leq\|D^{\boldsymbol{\lambda}_{1}}f\|_{L_{\rho}^{2}(\mathbb{R}^{d})}^{2}$
$\qquad\times\sup_{\boldsymbol{y}\in\mathbb{R}^{d}}\Bigl([D^{\boldsymbol{\lambda}_{2}}\rho](\boldsymbol{y})\Bigr)^{2}(\rho(\boldsymbol{y}))^{-1}\biggl(\prod_{k=1}^{d}\bigg(\prod_{\tau=1}^{\lambda_{3,k}}\big([D^{\tau}\phi]\circ\phi^{-1}(y_{k})\big)^{2\ell_{k,\tau}}\bigg)(\phi^{-1})^{\prime}(y_{k})\biggr)$
$=\|D^{\boldsymbol{\lambda}_{1}}f\|_{L_{\rho}^{2}(\mathbb{R}^{d})}^{2}$
$\ \ \times\prod_{k=1}^{d}\left(\sup_{y_{k}\in\mathbb{R}}\rho(y_{k})\,(\lambda_{2,k}!)\,H_{\lambda_{2,k}}^{2}(y_{k})\bigg(\prod_{\tau=1}^{\lambda_{3,k}}\big([D^{\tau}\phi]\circ\phi^{-1}(y_{k})\big)^{2\ell_{k,\tau}}\bigg)(\phi^{-1})^{\prime}(y_{k})\right)<\infty,$
where, in the last line, we used $(\phi^{-1})^{\prime}(y_{k})=1/\bigl(\pi(1+y_{k}^{2})\bigr)$ as well as the fact that $([D^{\tau}\phi]\circ\phi^{-1}(x_{k})\big)^{2\ell_{k,\tau}}$ is a polynomial of degree $2\ell_{k,\tau}(\tau+1)$; see Lemma 4.3.
We now assume $r_{\ell}\leq\alpha-1$ and prove the vanishing boundary condition for $w(\boldsymbol{t})$ in the $t_{\ell}$-direction, from which (21) will follow. First, from
$|w(\boldsymbol{t})|$ $=\bigg|\Big(([D^{\boldsymbol{\lambda}_{1}}f]\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Big)\Big(([D^{\boldsymbol{\lambda}_{2}}\rho]\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Big)\prod_{k=1}^{d}\prod_{\tau=1}^{\lambda_{3,k}}\big(D^{\tau}\phi(t_{k})\big)^{\ell_{k,\tau}}\bigg|$
$\leq\bigg|\Big(([D^{\boldsymbol{\lambda}_{1}}f]\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Big)\Big((\rho^{1/2+\varepsilon}\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Big)\bigg|$ (22)
$\;\;\times\bigg|\Big(([D^{\boldsymbol{\lambda}_{2}}\rho]\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Big)\Big((\rho^{-1/2-\varepsilon}\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Big)\prod_{k=1}^{d}\prod_{\tau=1}^{\lambda_{3,k}}\big(D^{\tau}\phi(t_{k})\big)^{\ell_{k,\tau}}\bigg|$
with $\varepsilon\in(0,1/2)$, in view of Proposition 2.3, $t_{\ell}\mapsto w(\boldsymbol{t})$ admits a continuous representative on $(0,1)$; indeed, following an argument analogous to the latter half of the proof of Proposition 4.1, the function $v:=[D^{\boldsymbol{\lambda}_{1}}f](\cdot;\boldsymbol{x}_{-\ell})\cdot\rho^{1/2+\varepsilon}(\cdot;\boldsymbol{x}_{-\ell})$ with $(\cdot,\boldsymbol{x}_{-\ell}):=(x_{1},\dots,x_{\ell-1},\cdot,x_{\ell+1},\dots,x_{d})$ is in $W_{\text{mix}}^{1,2}(\mathbb{R})$ for a.e. $(x_{1},\dots,x_{\ell-1},x_{\ell+1},\dots,x_{d})\in\mathbb{R}^{d-1}$. To see this, for simplicity consider the case $|\mathbf{r}|_{\infty}\leq\alpha-1$. Then, $\rho^{1/2+\varepsilon}D^{\boldsymbol{\lambda}_{1}}f\colon\mathbb{R}^{d}\to\mathbb{R}$ is in $W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})$. Indeed, for $\mathbf{k}\in\{0,1\}^{d}$, we have
$D^{\mathbf{k}}\bigl(\rho^{1/2+\varepsilon}D^{\boldsymbol{\lambda}_{1}}f\bigr)$ $=\sum_{\mathbf{j\leq k}}\binom{\mathbf{k}}{\mathbf{j}}D^{\mathbf{j}}(\rho^{1/2+\varepsilon})D^{\mathbf{k}-\mathbf{j}+\boldsymbol{\lambda}_{1}}f$
$=\sum_{\mathbf{j\leq k}}\binom{\mathbf{k}}{\mathbf{j}}\Bigl(\prod_{j\in\mathbf{j}}-\bigg(\frac{1}{2}+\varepsilon\bigg)x_{j}\Bigr)\rho^{1/2+\varepsilon}D^{\mathbf{k}-\mathbf{j}+\boldsymbol{\lambda}_{1}}f,$
so $\bigl\|D^{\mathbf{k}}\bigl(\rho^{1/2+\varepsilon}D^{\boldsymbol{\lambda}_{1}}f\bigr)\bigr\|_{L^{2}(\mathbb{R}^{d})}\leq\sum_{\mathbf{j\leq k}}\binom{\mathbf{k}}{\mathbf{j}}c_{\mathbf{j},\varepsilon}\|\rho^{1/2}D^{\mathbf{k}-\mathbf{j}+\boldsymbol{\lambda}_{1}}f\|_{L^{2}(\mathbb{R}^{d})}<\infty$. Showing $v\in W_{\mathrm{mix}}^{1,2}(\mathbb{R})$ when we only have $r_{\ell}\leq\alpha-1$ is analogous. As a result, noting that $(\Psi_{\mathrm{cot}}^{-1},\dots,\Psi_{\mathrm{cot}}^{-1})\colon\mathbb{R}^{d-1}\to(0,1)^{d-1}$ maps null sets to null sets, we see that $t_{\ell}\mapsto D^{\mathbf{r}}g(\boldsymbol{t})$ admits a continuous representative $(0,1)^{d-1}$-a.e. For $|\mathbf{r}|_{\infty}\leq\alpha-1$, the function $\rho^{1/2+\varepsilon}\cdot D^{\boldsymbol{\lambda}_{1}}f$ is in $W_{\text{mix}}^{1,2}(\mathbb{R}^{d})$ and thus the analogous claim holds everywhere in $(0,1)^{d-1}$.
For (21), we show that the first factor in (22) involving $f$ is bounded on $\mathbb{R}^{d}$, and that the second factor approaches zero as $\boldsymbol{t}$ goes to the boundary. For the first factor, again we use the argument in the proof of Proposition 4.1 to deduce $v\in W_{\text{mix}}^{1,2}(\mathbb{R})$, and thus from Proposition 2.3 $\sup_{x_{\ell}\in\mathbb{R}}|v(x_{\ell})|<\infty$. This holds everywhere in $\mathbb{R}^{d-1}$ if $|\mathbf{r}|_{\infty}\leq\alpha-1$, and a.e. in $\mathbb{R}^{d-1}$ otherwise. For the second factor, we have
$\lim_{t_{j}\to 0+,1-}\bigg|\Bigl(([D^{\boldsymbol{\lambda}_{2}}\rho]\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Bigr)\Bigl((\rho^{-1/2-\varepsilon}\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Bigr)\prod_{k=1}^{d}\prod_{\tau=1}^{\lambda_{3,k}}\big(D^{\tau}\phi(t_{k})\big)^{\ell_{k,\tau}}\bigg|$
$=\lim_{|x_{j}|\to\infty}\bigg|\Bigl([D^{\boldsymbol{\lambda}_{2}}\rho](\boldsymbol{x})\Bigr)\Bigl(\rho^{-1/2-\varepsilon}(\boldsymbol{x})\Bigr)\prod_{k=1}^{d}\bigg(\prod_{\tau=1}^{\lambda_{3,k}}\big([D^{\tau}\phi]\circ\phi^{-1}(x_{k})\big)^{\ell_{k,\tau}}\bigg)\bigg|$
$=\lim_{|x_{j}|\to\infty}\bigg|\Bigl(H_{\boldsymbol{\lambda}_{2}}(\boldsymbol{x})\;\rho^{1/2-\varepsilon}(\boldsymbol{x})\Bigr)\prod_{k=1}^{d}\bigg(\prod_{\tau=1}^{\lambda_{3,k}}\big([D^{\tau}\phi]\circ\phi^{-1}(x_{k})\big)^{\ell_{k,\tau}}\bigg)\bigg|=0,$
where in the last line, we again used the fact that $([D^{\tau}\phi]\circ\phi^{-1}(x_{k})\big)^{\ell_{k,\tau}}$ is a polynomial of degree $(\tau+1)\ell_{k,\tau}$; see Lemma 4.3. Thus we have proved (21). ∎
Using Proposition 4.7, we will show that if $f\in H_{\rho}^{\alpha}$, then after a Möbius transformation it belongs to Sobolev-type spaces on the unit cube, where QMC rules yield small errors. We use two such spaces. The first is the Korobov space, given by
${H}_{\text{Kor}}^{\alpha}$ $:=\bigg\{f\in L^{2}([0,1]^{d})\bigg|\|f\|^{2}_{\text{Kor},\alpha}:=\sum_{\mathbf{k}\in\mathbb{Z}^{d}}|\widehat{f}(\mathbf{k})|^{2}r_{\alpha}^{2}(\mathbf{k})<\infty,$
$\qquad r_{\alpha}(\mathbf{k}):=\prod_{j=1}^{d}\max(1,|k_{j}|^{\alpha}),\;\;\widehat{f}(\mathbf{k}):=\int_{[0,1]^{d}}f(\boldsymbol{t})\mathrm{e}^{-2\pi\mathrm{i}\boldsymbol{t}\cdot\mathbf{k}}\mathrm{d}\boldsymbol{t}\bigg\}.$
The second is the unanchored Sobolev space, defined by (12) with $\boldsymbol{a}=\boldsymbol{0}$ and $\boldsymbol{b}=\boldsymbol{1}$.
For the Korobov space, the rank-$1$ lattice rule constructed by [12, Algorithm 3.14] achieves the convergence rate $N^{-\alpha}(\ln N)^{d\alpha}$. For the unanchored Sobolev space, higher-order digital nets achieve the optimal rate $N^{-\alpha}(\ln N)^{(d-1)/2}$ [29]. This rate is optimal, including the logarithmic factor, among all linear and nonlinear quadrature classes that rely solely on function evaluations. QMC rules for $\mathbb{R}^{d}$ with Möbius transformation inherit these results: those based on rank-$1$ lattice rules achieve the optimal rate up to a logarithmic factor, whereas those based on higher-order digital nets inherit full optimality, including the logarithmic factor, as we show below.
Remark 1 (Evaluation at the boundary).
Lattice points and digital nets on the unit cube include boundary points such as $\boldsymbol{0}$, where the Möbius transformation is undefined. Proposition 4.7 tells us that this is not a problem, since the integrand $g(\boldsymbol{t})=f(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t}))\rho(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t}))\prod_{k=1}^{d}\phi^{\prime}(t_{k})$ for $f\in H_{\rho}^{\alpha}$ admits an extension by zero to the boundary. Thus, in the implementation of QMC methods using such points, boundary values may be set to zero.
Proposition 4.8 (Lattice with Möbius transformation).
Let $Q_{N}$ be the rank-$1$ lattice rule with cotangent transform as in (17). Then, with a generating vector $\boldsymbol{z}$ obtained by the CBC construction in [12, Algorithm 3.14], $Q_{N}$ satisfies
$\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}\leq C_{\alpha,d}N^{-\alpha}(\ln N)^{d\alpha}.$ (23)
Proof.
For $f\in H_{\rho}^{\alpha}$, set $g(\boldsymbol{t})=f\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\right)\rho\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\right)\prod_{k=1}^{d}\phi^{\prime}(t_{k})$ for $\boldsymbol{t}\in(0,1)^{d}$. Fix $\ell\in\{1,\dots,d\}$ and choose $\mathbf{r}\in\{0,\ldots,\alpha\}^{d}$, $r_{\ell}\leq\alpha-1$. Noting that $t_{\ell}\mapsto D^{\mathbf{r}}g(\boldsymbol{t})\mathrm{e}^{-2\pi\mathrm{i}\boldsymbol{t}\cdot\mathbf{k}}$ is absolutely continuous on the interval $(0,1)$ for a.e. $\boldsymbol{t}_{-\ell}\in(0,1)^{d}$, we use integration by parts in $(0,1)$ [44, Corollary 3.23], as well as uniformly continuity up to $[0,1]$ and the integrability of functions under consideration on $(0,1)$, to obtain
$\int_{0}^{1}D^{\mathbf{r}+\mathbf{e}_{\ell}}g(\boldsymbol{t})\mathrm{e}^{-2\pi\mathrm{i}\boldsymbol{t}\cdot\mathbf{k}}\mathrm{d}t_{\ell}$ $=2\pi\mathrm{i}k_{\ell}\int_{0}^{1}D^{\mathbf{r}}g(\boldsymbol{t})\mathrm{e}^{-2\pi\mathrm{i}\boldsymbol{t}\cdot\mathbf{k}}\mathrm{d}t_{\ell},$
where we used Proposition 4.7 to handle the boundary terms. Here, $\mathbf{e}_{\ell}$ denotes the $\ell$-th standard basis in $\mathbb{R}^{d}$. Using Proposition 4.7 once more, we see
$\|g\|_{\mathrm{Kor},\alpha}^{2}$ $\leq\sum_{\mathbf{k}\in\mathbb{Z}^{d}}\Bigl(\prod_{j=1}^{d}\sum_{r=0}^{\alpha}(2\pi k_{j})^{2r}\Bigr)\big|\widehat{g}(\mathbf{k})\big|^{2}=\sum_{\mathbf{k}\in\mathbb{Z}^{d}}\sum_{|\mathbf{r}|_{\infty}\leq\alpha}\Bigl(\prod_{j=1}^{d}(2\pi k_{j})^{2r_{j}}\Bigr)\big|\widehat{g}(\mathbf{k})\big|^{2}$
$=\sum_{|\mathbf{r}|_{\infty}\leq\alpha}\|D^{\mathbf{r}}g\|_{L^{2}((0,1)^{d})}^{2}\leq C\|f\|_{H_{\rho}^{\alpha}}^{2}<\infty.$
Finally, since $Q_{N}$ satisfies $|I(f)-Q_{N}(f)|\leq C\|g\|_{\mathrm{Kor},\alpha}N^{-\alpha}(\ln N)^{d\alpha}$, the statement follows. ∎
Proposition 4.9 (Digital net with Möbius transformation).
Let $Q_{N}$ be the QMC with digital net with cotangent transform as in (17), where points $(\boldsymbol{t}^{(j)})_{j=1,\dots,N}$ are given by a higher-order net of order $(2\alpha+1)$ with $N=p^{m}$ elements. Then, $Q_{N}$ satisfies
$\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}\leq C_{\alpha,d}N^{-\alpha}(\ln N)^{(d-1)/2}.$ (24)
Proof.
From [29, Theorem 1], the QMC under study $Q_{N}$ satisfies $|I(f)-Q_{N}(f)|\leq C_{d,\alpha}\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{0},\boldsymbol{1}])}N^{-\alpha}(\ln N)^{(d-1)/2}$, where $g(\boldsymbol{t})=f\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\right)\rho\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\right)\prod_{k=1}^{d}\phi^{\prime}(t_{k})$, and $\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{0},\boldsymbol{1}])}$ is defined in (12). From Proposition 4.7, this norm can be bounded in terms of $\|f\|_{H_{\rho}^{\alpha}}$ as
$\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{0},\boldsymbol{1}])}^{2}\leq\sum_{\mathbf{r}\in\{0,\dots,\alpha\}^{d}}\|D^{\mathbf{r}}g\|_{L^{2}((0,1)^{d})}^{2}\leq C_{d,\alpha}\|f\|_{H_{\rho}^{\alpha}}.$
This completes the proof. ∎
We summarise the results thus far in the following theorem.
Theorem 4.10.
The quasi-Monte Carlo rule $Q_{N}$ (9) achieves the error bound
$\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}$ $\leq C_{d,\alpha}\frac{(\ln N)^{\gamma}}{N^{\alpha}},$
where the exponent $\gamma$ depends on the QMC construction:
[…]
The polynomial convergence rate $\alpha$ is optimal in $H_{\rho}^{\alpha}$: no algorithm can achieve a better polynomial rate. Furthermore, the exponent (d1)/2(d-1)/2 in the logarithmic factor is also optimal.
5 Concluding remarks
In this paper, we showed that quasi-Monte Carlo methods with a change of variables achieve optimal worst-case convergence rate in Gaussian Sobolev spaces, whereas sparse-grid quadrature based on the one-dimensional Gauss–Hermite rule is suboptimal in this sense.
These findings motivate further studies in several directions. One is to analyse sparse-grid algorithms for functions over $\mathbb{R}^{d}$ built from univariate rules other than the Gauss–Hermite rule. A notable alternative is the sparse-grid method using weighted Leja sequences [22, 48], which was proposed by Narayan and Jakeman [48]. To obtain results analogous to what we showed in this paper, it is worth noting that the fooling function we constructed in Theorems 3.1 and  3.2 depend solely on the distribution of quadrature nodes, and not on the quadrature weights. Hence, understanding the distribution of Leja points for the weight of interest would allow the construction of a corresponding fooling function.
We did not discuss dimension independence of the constant. The implied constant in our error estimate may depend on the dimension $d$, which may be large in practice. Addressing this is left for future work. A standard approach is to employ the so-called weighted spaces, as introduced by Hickernell [33], under which QMC methods can attain dimension-independent convergence rates with constants that are likewise independent of dimension, as shown by Sloan and Woźniakowski [59]. These ideas have been widely applied in the analysis of parametric partial differential equations [30, 32, 39, 40, 58]. In fact, the results in [59] show that such weights are even necessary in the setting considered there, suggesting that additional structure is required to make the constants in our bounds dimension-independent.
This paper focused on Sobolev functions. Another important regime in high dimension is that of analytic functions [8, 35, 17], where the question of optimal rates in this setting remains largely open. In establishing the lower bound for the sparse-grid Gauss–Hermite rule in this paper, a crucial component was the fooling function constructed for univariate Sobolev functions [41]. A similar technique for analytic functions is available [27, 63], which may prove useful for further analysis in this regime.

### v2

@@
 In contrast, several quasi-Monte Carlo methods with a change of variables are shown to achieve the optimal rate, some up to, and one including, the logarithmic factor.
+keywords:
+high-dimensional integration, quasi-Monte Carlo, digital net, rank-1 lattice rule, Möbius transformation, Smolyak algorithm, Gauss–Hermite quadrature, optimal algorithm
+{MSCcodes}
+65C05, 65D30, 65D32, 65D40
 1 Introduction
 This paper is concerned with numerical integration of multivariate functions over the whole space in high dimensions. Our focus is on integration with respect to the standard Gaussian measure:
-$I(f):=\int_{\mathbb{R}^{d}}f(\boldsymbol{x})\frac{\mathrm{e}^{-|\boldsymbol{x}|^{2}/2}}{(2\pi)^{d/2}}\mathrm{d}\boldsymbol{x}\approx Q_{N}(f),$
-where $|\boldsymbol{x}|=\sqrt{x_{1}^{2}+\dotsb+x_{d}^{2}}$ for $\boldsymbol{x}\in\mathbb{R}^{d}$ is the Euclidean norm, and $Q_{N}$ is a suitable numerical integration rule using $N$ evaluations of $f$.
-Such problems arise across a broad range of applications, including data assimilation [36], robotics [3], life insurance [23], physics [45].
-They also serve as simpler problems to which more complex integration tasks can be reduced [20].
-In high dimensions $d\gg 1$, two popular classes of numerical integration rules $Q_{N}$ are sparse-grid quadrature [6] and quasi-Monte Carlo (QMC) methods [13].
+$I(f):=\int_{\mathbb{R}^{d}}f(\boldsymbol{x})\frac{\mathrm{e}^{-|\boldsymbol{x}|_{{2}}^{2}/2}}{(2\pi)^{d/2}}\mathrm{d}\boldsymbol{x}\approx Q_{N}(f),$
+where $|\boldsymbol{x}|_{2}=\sqrt{x_{1}^{2}+\dotsb+x_{d}^{2}}$ for $\boldsymbol{x}\in\mathbb{R}^{d}$ is the Euclidean norm, and $Q_{N}$ is a suitable numerical integration rule using $N$ evaluations of $f$.
+Such problems arise across a broad range of applications, including data assimilation [undefak], robotics [undefb], life insurance [undefw], physics [undefau].
+They also serve as simpler problems to which more complex integration tasks can be reduced [undeft].
+In high dimensions $d\gg 1$, two popular classes of numerical integration rules $Q_{N}$ are sparse-grid quadrature [undefe] and quasi-Monte Carlo (QMC) methods [undefm].
 Sparse-grid quadrature is a numerical integration technique based on univariate quadrature rules.
 Unlike fully tensorized quadrature, it significantly reduces the number of evaluation points, making it attractive for high-dimensional problems.
-A commonly used univariate quadrature for approximating $I(f)$ is the Gauss–Hermite rule [38, 56, 19, 4, 66], which is the focus of this paper.
+A commonly used univariate quadrature for approximating $I(f)$ is the Gauss–Hermite rule [undefam, undefaaf, undefs, undefc, undefaap], which is the focus of this paper.
 Quasi-Monte Carlo, on the other hand, is an equal-weight numerical integration.
@@
 Examples of point sets include low-discrepancy point sets and sequences, lattice rules, and digital nets; these classes are not mutually exclusive.
-These methods are also widely used for integration with respect to the Gaussian measure and have found applications across a diverse range of scientific fields, such as uncertainty quantification [9, 42], computational physics [37], finance [24, 43], and machine learning [1, 47].
+These methods are also widely used for integration with respect to the Gaussian measure and have found applications across a diverse range of scientific fields, such as uncertainty quantification [undefi, undefaq], computational physics [undefal], finance [undefx, undefar], and machine learning [undef, undefaw].
 In this paper, we show that sparse-grid Gauss–Hermite quadrature is suboptimal, whereas several quasi-Monte Carlo methods are optimal in terms of convergence rate.
@@
 The rate $O(N^{-\alpha})$ is in fact optimal, as there exists a matching lower bound established by Dick et al.
-[11].
+[undefk].
 In this sense, QMC methods considered herein are rate-optimal.
@@
 Such a gap in convergence rates between the sparse-grid Gauss–Hermite rule and QMC methods has also been observed numerically by Dick et al.
-[11] and in subsequent work by Nuyens and one of the present authors [55].
+[undefk] and in subsequent work by Nuyens and one of the present authors [undefaae].
 Our theoretical results provide a rigorous explanation of these empirical findings.
-Analysis of the sparse-grid Gauss–Hermite rule and QMC methods on the whole space has been carried out before, with many papers set in the context of partial differential equations with random coefficients [7, 15, 17, 21, 30, 31, 32, 40, 49, 58].
+Analysis of the sparse-grid Gauss–Hermite rule and QMC methods on the whole space has been carried out before, with many papers set in the context of partial differential equations with random coefficients [undefg, undefo, undefq, undefu, undefad, undefae, undefaf, undefao, undefay, undefaah].
 This paper provides theoretical results for a standard function space, without specific structure of the underlying model problem such as the form of the random field or partial differential equations whose differential operator is given by such a field.
 In the same vein, aforementioned work by Dick et al.
-[11] studied quasi-Monte Carlo methods in a Sobolev space, the space we also consider herein.
+[undefk] studied quasi-Monte Carlo methods in a Sobolev space, the space we also consider herein.
 They showed that the QMC method considered there attains the optimal convergence rate by proving matching upper and lower bounds, although the logarithmic factor was not optimal.
-Nuyens and one of the present authors considered a rank-$1$ lattice QMC over $\mathbb{R}^{d}$ [55] for functions with suitable decay.
+Nuyens and one of the present authors considered a rank-$1$ lattice QMC over $\mathbb{R}^{d}$ [undefaae] for functions with suitable decay.
 We revisit their results and show that the resulting QMC method achieves the optimal rate in the Sobolev space considered herein, although the logarithmic factor is again not optimal.
-In [64], Hyvönen, Karvonen, and one of the present authors considered a change of variables that maps functions on the real line to the unit interval.
+In [undefaan], Hyvönen, Karvonen, and one of the present authors considered a change of variables that maps functions on the real line to the unit interval.
 Although such an approach typically incurs undesirable unboundedness near the boundary, they showed that, with a carefully chosen Möbius transformation, the resulting mapping sends univariate functions from the Gaussian Sobolev space on the real line to the periodic Sobolev space of the same smoothness on the unit interval.
@@
 These rates turn out to be optimal in Gaussian Sobolev spaces: rank-$1$ lattice rules achieve the optimal rate up to a logarithmic factor, whereas higher-order digital nets achieve the optimal rate including the logarithmic factor.
-Dũng and Nguyen [16] proposed two numerical integration methods that are optimal – including the logarithmic factor – in $L^{p}$-Sobolev spaces for $1<p<\infty$.
+Dũng and Nguyen [undefp] proposed two numerical integration methods that are optimal – including the logarithmic factor – in $L^{p}$-Sobolev spaces for $1<p<\infty$.
 One method requires implementing a partition of unity with suitable properties.
@@
 Sections 3 and 4 present our main results on the sparse-grid Gauss–Hermite quadrature and QMC methods, respectively.
-These results build on the authors’ previous works [29, 41, 64].
+These results build on the authors’ previous works [undefac, undefap, undefaan].
 Finally, Section 5 concludes the paper.
@@
 $\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}:=\|f\|_{H_{\rho}^{\alpha}}:=\left(\sum_{|\mathbf{r}|_{\infty}\leq\alpha}\|D^{\mathbf{r}}f\|_{L_{\rho}^{2}(\mathbb{R}^{d})}^{2}\right)^{\frac{1}{2}},$
-where $\rho(\boldsymbol{x})=\frac{\mathrm{e}^{-|\boldsymbol{x}|^{2}/2}}{(2\pi)^{d/2}}$ and $D^{\mathbf{r}}f$ is the weak derivative of $f$.
+where $\rho(\boldsymbol{x})=\frac{\mathrm{e}^{-|\boldsymbol{x}|_{{2}}^{2}/2}}{(2\pi)^{d/2}}$ and $D^{\mathbf{r}}f$ is the weak derivative of $f$.
 Here, $\|\cdot\|_{L_{\rho}^{2}(\mathbb{R}^{d})}$ is the norm induced by the weighted $L^{2}$-inner product $\langle f,g\rangle_{L_{\rho}^{2}}:=\int_{\mathbb{R}^{d}}f(\boldsymbol{x})g(\boldsymbol{x})\rho(\boldsymbol{x})\mathrm{d}\boldsymbol{x}$ for $f,g\in L_{\rho}^{2}(\mathbb{R}^{d})$.
-It is known that this space and the so-called Hermite space coincide as vector spaces, with equivalent norms.
+Throughout this paper, we denote by $L_{\mathrm{loc}}^{1}(\mathbb{R}^{d})$ the space of locally integrable functions on $\mathbb{R}^{d}$.
+It is known that this space and the so-called Hermite space introduced in [undefaj] and generalized in [undefk] coincide as vector spaces, with equivalent norms.
 More specifically, Dick et al.
-[11] studied quasi-Monte Carlo methods in the Hermite space and showed that every element of this space also belongs to $H_{\rho}^{\alpha}(\mathbb{R}^{d})$, with the embedding being continuous; the reverse inclusion is discussed in [41, 16].
-This space is also known to coincide with the so-called modulation space, as discussed in [18].
-Further properties of Hermite spaces are discussed in [25, 26].
+[undefk] studied quasi-Monte Carlo methods in the Hermite space and showed that every element of this space also belongs to $H_{\rho}^{\alpha}(\mathbb{R}^{d})$, with the embedding being continuous; the reverse inclusion is discussed in [undefap, undefp].
+This space is also known to coincide with the so-called modulation space, as discussed in [undefr].
+Further properties of Hermite spaces are discussed in [undefy, undefas, undefz].
 In the space $H_{\rho}^{\alpha}(\mathbb{R}^{d})$, a lower bound on the worst-case integration error is known for general algorithms.
-Theorem 2.1 ([11]).
+Theorem 2.1 ([undefk]).
 For $N\geq 2$, let $A_{N}\colon H_{\rho}^{\alpha}(\mathbb{R}^{d})\to\mathbb{R}$ be a mapping (linear or nonlinear) that uses only $N$ values as information about the argument, i.e., it is of the form $A_{N}(f)=\mathcal{I}_{N}(f(\boldsymbol{x}_{1}),\dots,f(\boldsymbol{x}_{N}))$ for a mapping $\mathcal{I}_{N}\colon\mathbb{R}^{N}\to\mathbb{R}$.
@@
 We will show that several QMC methods attain a rate of the form $\frac{(\ln N)^{\gamma}}{N^{\alpha}}$, where the exponent $\gamma$ depends on the method.
-The rate $\gamma=\frac{d-1}{2}$ turns out to be attainable, hence optimal in the sense above.
+The exponent $\gamma=\frac{d-1}{2}$ turns out to be attainable, hence optimal in the sense above.
 In contrast, the sparse-grid quadrature based on the Gauss–Hermite rule is of the rate $N^{-\alpha/2}$, up to a logarithmic factor, which is subotimal in the sense of worst-case error convergence rate.
-Note that [11, Theorem 1] is formulated for linear algorithms.
-As noted in [11, Section 3], however, in view of for example [2] (see also [53, Section 4.2]), their lower bound also holds for nonlinear algorithms.
-Moreover, the implied constant in the result above may differ from that in [11, Theorem 1], since the present bound is formulated for the Sobolev space, whereas [11, Theorem 1] considers a norm-equivalent space, namely the Hermite space.
+Note that [undefk, Theorem 1] is formulated for linear algorithms.
+As noted in [undefk, Section 3], however, in view of for example [undefa] (see also [undefaac, Section 4.2]), their lower bound also holds for nonlinear algorithms.
+Moreover, the implied constant in the result above may differ from that in [undefk, Theorem 1], since the present bound is formulated for the Sobolev space, whereas [undefk, Theorem 1] considers a norm-equivalent space, namely the Hermite space.
+Further characterizations of Hermite spaces are investigated in [undefas].
 Let $H_{k}$ for $k\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$ denote the normalized probabilistic Hermite polynomial of degree $k$, defined by
@@
 For example, $f\lesssim_{d}g$ means there exists a constant $c_{d}\in(0,\infty)$, which may depend on $d$, such that $f(x)\leq c_{d}\,g(x)$ for all $x\in S$; likewise, $f\asymp_{d}g$ means $f\lesssim_{d}g$ and $g\lesssim_{d}f$.
+We also use the following notations: $\mathbf{a}!:=\prod_{j=1}^{d}a_{j}!$ and $\binom{\mathbf{a}}{\mathbf{b}}:=\prod_{j=1}^{d}\binom{a_{j}}{b_{j}}$ for $\mathbf{a},\mathbf{b}\in\mathbb{N}^{d}_{0}$.
+Furthermore, for $\mathbf{k}\in\mathbb{N}_{0}^{d}$ we let $|\mathbf{k}|:=k_{1}+\dotsb+k_{d}$.
 2.2 Tensorised operators
@@
 The $\mathbb{R}$-vector space spanned by $d$-variate polynomials $\mathcal{P}:=\mathrm{span}\Big\{\prod_{j=1}^{d}H_{k_{j}}\mid k_{j}\in\mathbb{N}_{0},\,j=1,\dots,d\Big\}$ is dense in $H_{\rho}^{\alpha}(\mathbb{R}^{d})$.
-Proof.
-Fix $f\in H_{\rho}^{\alpha}$. Following the argument in [41, Lemma 2.1] we see
-$\langle D^{\mathbf{r}}f,H_{\mathbf{k}}\rangle_{L_{\rho}^{2}}=\biggl(\prod_{j=1}^{d}\prod_{m=1}^{\mathrm{r}_{j}}\bigl(k_{j}+m\bigr)\biggr)^{1/2}\langle f,H_{\mathbf{k}+\mathbf{r}}\rangle_{L_{\rho}^{2}}\quad\text{for }\text{$\mathbf{k}$}\in\mathbb{N}_{0}^{d},\ \mathbf{r}\in\{0,\dots,\alpha\}^{d}.$
+Proof 2.3.
+Fix $f\in H_{\rho}^{\alpha}$. Following the argument in [undefap, Lemma 2.1] we see
+$\langle D^{\mathbf{r}}f,H_{\mathbf{k}}\rangle_{L_{\rho}^{2}}=\biggl(\prod_{j=1}^{d}\prod_{m=1}^{\mathrm{r}_{j}}\bigl(k_{j}+m\bigr)\biggr)^{1/2}\langle f,H_{\mathbf{k}+\mathbf{r}}\rangle_{L_{\rho}^{2}}\quad\text{for }\mathbf{k}\in\mathbb{N}_{0}^{d},\ \mathbf{r}\in\{0,\dots,\alpha\}^{d}.$
 With this, using the Parseval identity for derivatives we see that the polynomial $f_{N}=\sum_{|\mathbf{k}|\leq N}\langle f,H_{\mathbf{k}}\rangle_{L_{\rho}^{2}}H_{\mathbf{k}}$ can be made arbitrarily close to $f$ in $H_{\rho}^{\alpha}$:
@@
 Define $T_{1}\otimes\dotsb\otimes T_{d}:\mathcal{P}\to H_{\rho}^{\alpha}(\mathbb{R}^{d})$ by defining
-$\left[T_{1}\otimes\dotsb\otimes T_{d}\Bigg(\prod_{j=1}^{d}H_{k_{j}}\Bigg)\right](\boldsymbol{x})=\prod_{j=1}^{d}[T_{j}(H_{k_{1}})](x_{j})\qquad\text{for }\prod_{j=1}^{d}H_{k_{j}}\in\mathcal{P}$
+$\left[T_{1}\otimes\dotsb\otimes T_{d}\Bigg(\prod_{j=1}^{d}H_{k_{j}}\Bigg)\right](\boldsymbol{x})=\prod_{j=1}^{d}[T_{j}(H_{k_{{j}}})](x_{j})\qquad\text{for }\prod_{j=1}^{d}H_{k_{j}}\in\mathcal{P}$
 and extending it linearly to $\mathcal{P}$.
-Then, using the argument for tensor product Hilbert spaces [57, Section VIII.10] we see $\sup_{p\in\mathcal{P},\|p\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}=1}\|T_{1}\otimes\dotsb\otimes T_{d}(p)\|=\prod_{j=1}^{d}\|T_{j}\|_{H_{\rho}^{\alpha}(\mathbb{R})\to H_{\rho}^{\alpha}(\mathbb{R})}$ on $\mathcal{P}$, and from Proposition 2.2, the operator $T_{1}\otimes\dotsb\otimes T_{d}$ extends continuously to $H_{\rho}^{\alpha}(\mathbb{R}^{d})$ with the same operator norm.
-Note that any continuous linear functional on $H_{\rho}^{\alpha}(\mathbb{R}^{d})$ is a bounded linear operator in our setting.
-Indeed, since the total measure is normalised to $1$, any constant belongs to $H_{\rho}^{\alpha}(\mathbb{R}^{d})$, and if $T_{j}$ is a functional then we have $T_{j}(f)\in H_{\rho}^{\alpha}(\mathbb{R})$ with $\|T_{j}(f)\|_{H_{\rho}^{\alpha}(\mathbb{R})}=|T_{j}(f)|$ for $f\in H_{\rho}^{\alpha}(\mathbb{R})$.
+Then, using the argument for tensor product Hilbert spaces [undefaag, Section VIII.10] we see $\sup_{p\in\mathcal{P},\|p\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}=1}\|T_{1}\otimes\dotsb\otimes T_{d}(p)\|_{{H_{\rho}^{\alpha}(\mathbb{R}^{d})}}=\prod_{j=1}^{d}\|T_{j}\|_{H_{\rho}^{\alpha}(\mathbb{R})\to H_{\rho}^{\alpha}(\mathbb{R})}$ on $\mathcal{P}$, and from Proposition 2.2, the operator $T_{1}\otimes\dotsb\otimes T_{d}$ extends continuously to $H_{\rho}^{\alpha}(\mathbb{R}^{d})$ with the same operator norm.
+Note that this construction applies also when the $T_{j}$ are bounded linear functionals.
+In such a case, since $T_{j}(f)$ is a constant function for $f\in H_{\rho}^{\alpha}(\mathbb{R})$ and the total measure is normalised to $1$, we have $\|T_{j}(f)\|_{H_{\rho}^{\alpha}(\mathbb{R})}=|T_{j}(f)|$.
+From this, we see that $T_{j}$ is also a bounded operator from $H_{\rho}^{\alpha}(\mathbb{R})$ to itself with $\|T_{j}\|_{H_{\rho}^{\alpha}(\mathbb{R})\to H_{\rho}^{\alpha}(\mathbb{R})}=\|T_{j}\|_{H_{\rho}^{\alpha}(\mathbb{R})\to\mathbb{R}}$.
+Similarly, for the functional $T_{1}\otimes\dotsb\otimes T_{d}$ on $\mathcal{P}$ constructed as above, $T_{1}\otimes\dotsb\otimes T_{d}(p)$ is a constant function so we have $|T_{1}\otimes\dotsb\otimes T_{d}(p)|=\|T_{1}\otimes\dotsb\otimes T_{d}(p)\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$.
+Hence, $T_{1}\otimes\dotsb\otimes T_{d}$ satisfies $\sup_{p\in\mathcal{P},\|p\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}=1}|T_{1}\otimes\dotsb\otimes T_{d}(p)|=\prod_{j=1}^{d}\|T_{j}\|_{H_{\rho}^{\alpha}(\mathbb{R})\to\mathbb{R}}$ and admits a continuous extension to $H_{\rho}^{\alpha}(\mathbb{R}^{d})$ with the same functional norm.
 Operators with the symbol $\otimes$ in this paper are constructed explicitly in this manner.
@@
 Every element of $W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})$ admits a bounded continuous representative $v$ satisfying
-$\sup_{\boldsymbol{x}\in\mathbb{R}^{d}}|v(\boldsymbol{x})|$ $\leq\|v\|_{W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})},$ (1)
+(1) $\sup_{\boldsymbol{x}\in\mathbb{R}^{d}}|v(\boldsymbol{x})|$ $\leq\|v\|_{W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})},$
 and
-$|v(\boldsymbol{x})-v(\boldsymbol{y})|$ $\leq d^{3/4}\|v\|_{W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})}|\boldsymbol{x}-\boldsymbol{y}|_{2}^{1/2}\qquad\text{for all }\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}.$ (2)
-Proof.
+(2) $|v(\boldsymbol{x})-v(\boldsymbol{y})|$ $\leq d^{3/4}\|v\|_{W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})}|\boldsymbol{x}-\boldsymbol{y}|_{2}^{1/2}\qquad\text{for all }\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}.$
+Proof 2.5.
 We will first prove the statement for infinitely differentiable functions with compact support, $C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$, and then extend it by a density argument.
-For $v\in C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$, following [5, Proof of Theorem 8.8] we get
+For $v\in C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$, following [undefd, Proof of Theorem 8.8] we get
 $|v(x_{1},\dots,x_{d})|^{2}$ $\leq 2\sqrt{\int_{\mathbb{R}}|v(s,x_{2},\dots,x_{d})|^{2}\mathrm{d}s}\sqrt{\int_{\mathbb{R}}|D^{(1,0,\dots,0)}v(s,x_{2},\dots,x_{d})|^{2}\mathrm{d}s}$
@@
 $\sum_{j=1}^{d}|x_{j}-y_{j}|^{1/2}\leq\bigg(\sum_{j=1}^{d}|x_{j}-y_{j}|^{2}\bigg)^{1/4}\big(\sum_{j=1}^{d}\;1\big)^{3/4}=d^{3/4}|\boldsymbol{x}-\boldsymbol{y}|_{2}^{1/2},$
-which shows the second inequality for $C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$. Using the density of $C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$ completes the proof. ∎
+which shows the second inequality for $C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$. Using the density of $C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$ completes the proof.
 3 Gauss–Hermite sparse-grid quadrature
@@
 Given an index set $\Lambda\subset\mathbb{N}^{d}$, the sparse-grid quadrature $S_{\Lambda}\colon H_{\rho}^{\alpha}(\mathbb{R}^{d})\to\mathbb{R}$ is defined by
-$S_{\Lambda}:=\sum_{(\ell_{1},\dots,\ell_{d})\in\Lambda}\Delta_{\ell_{1}}\otimes\dotsb\otimes\Delta_{\ell_{d}}.$ (3)
+(3) $S_{\Lambda}:=\sum_{(\ell_{1},\dots,\ell_{d})\in\Lambda}\Delta_{\ell_{1}}\otimes\dotsb\otimes\Delta_{\ell_{d}}.$
 The purpose of this section is to analyse the error of $S_{\Lambda}$ when $Q_{\ell}^{\mathrm{uni}}$ is the Gauss–Hermite quadrature.
@@
 where the constant $c_{\alpha}>0$ is independent of $N$ and $d$.
-Proof.
-Our strategy follows [41, Section 3.1].
+Proof 3.2.
+Our strategy follows [undefap, Section 3.1].
 The heart of the matter is to construct a suitable fooling function, which we denote by $h_{\Lambda}$ below.
-The function is based on the earlier construction from [41, Section 3.1].
 Without loss of generality, we assume that $\Lambda$ is downward-closed in the first coordinate.
 Then, for some $\overline{\ell}_{1}\in\mathbb{N}$ we have $(\ell_{1},1\dots,1)\in\Lambda$ for all $\ell_{1}=1,\dots,\overline{\ell}_{1}$.
-For example, if $\Lambda$ is the classical isotropic index set $\Lambda=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid\,|\boldsymbol{\ell}|_{1}\leq L\}$ then $\overline{\ell}_{1}=L-d+1$.
+For example, if $\Lambda$ is the classical isotropic index set $\Lambda=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid\,|{\boldsymbol{\ell}|}\leq L\}$ then $\overline{\ell}_{1}=L-d+1$.
 Let $n:=n_{\overline{\ell}_{1}}$ denote the number of quadrature points used in the univariate Gauss–Hermite rule $Q_{\overline{\ell}_{1}}^{\mathrm{uni}}$, and let $\xi_{1},\dots,\xi_{n}$ be the roots of the univariate Hermite polynomial $H_{n}$ of degree $n$.
-Following [41], define the function $p_{n}(t)\colon\mathbb{R}\to\mathbb{R}$ by
+Following [undefap], define the function $p_{n}(t)\colon\mathbb{R}\to\mathbb{R}$ by
 […]
 With this, define
-$h_{\Lambda}(\boldsymbol{x}):=p_{n}(x_{1})\quad\text{for }\boldsymbol{x}=(x_{1},\dots,x_{d})\in\mathbb{R}^{d}.$ (4)
+(4) $h_{\Lambda}(\boldsymbol{x}):=p_{n}(x_{1})\quad\text{for }\boldsymbol{x}=(x_{1},\dots,x_{d})\in\mathbb{R}^{d}.$
 First, we will show
-$\frac{|I(h_{\Lambda})-S_{\Lambda}(h_{\Lambda})|}{\|h_{\Lambda}\|_{H_{\rho}^{\alpha}}}=\frac{\int_{\mathbb{R}}p_{n}(x)\rho(x)\mathrm{d}x}{\left(\sum_{r=0}^{\alpha}\|D^{r}p_{n}\|_{L_{\rho}^{2}}^{2}\right)},$ (5)
+(5) $\frac{|I(h_{\Lambda})-S_{\Lambda}(h_{\Lambda})|}{\|h_{\Lambda}\|_{H_{\rho}^{\alpha}}}=\frac{\int_{\mathbb{R}}p_{n}(x)\rho(x)\mathrm{d}x}{\left(\sum_{r=0}^{\alpha}\|D^{r}p_{n}\|_{L_{\rho}^{2}}^{2}\right)^{{1/2}}},$
 so that the problem reduces to a one-dimensional integration problem.
@@
 As for the norm, since $h_{\Lambda}(\boldsymbol{x})=p_{n}(x_{1})$ we obtain
-$\|h_{\Lambda}\|_{H_{\rho}^{\alpha}}$ $=\left(\sum_{|\mathbf{r}|_{\infty}\leq\alpha}\|D^{(r_{1},0,\dots,0)}h_{\Lambda}\|_{L_{\rho}^{2}}^{2}\right)^{\frac{1}{2}}=\left(\sum_{r=0}^{\alpha}\|D^{r}p_{n}\|_{L_{\rho}^{2}}^{2}\right)^{\frac{1}{2}}$
+$\|h_{\Lambda}\|_{H_{\rho}^{\alpha}}$ $=\left(\sum_{|\mathbf{r}|_{\infty}\leq\alpha}\|D^{(r_{1},0,\dots,0)}h_{\Lambda}\|_{L_{\rho}^{2}}^{2}\right)^{\frac{1}{2}}=\left(\sum_{r=0}^{\alpha}\|D^{r}p_{n}\|_{L_{\rho}^{2}}^{2}\right)^{\frac{1}{2}}{.}$
 Therefore, (5) holds.
-The right-hand side of (5) corresponds to the expression in the last displayed equation in [41, Proof of Lemma 3.1], which is further analysed in [41, Theorem 3.2]. There, it is shown
-$\frac{\int_{\mathbb{R}}p_{n}(x)\rho(x)\mathrm{d}x}{\left(\sum_{r=0}^{\alpha}\|D^{r}p_{n}\|_{L_{\rho}^{2}}^{2}\right)}\geq c_{\alpha}n^{-\alpha/2},$
+The right-hand side of (5) corresponds to the expression in the last displayed equation in [undefap, Proof of Lemma 3.1], which is further analysed in [undefap, Theorem 3.2]. There, it is shown
+$\frac{\int_{\mathbb{R}}p_{n}(x)\rho(x)\mathrm{d}x}{\left(\sum_{r=0}^{\alpha}\|D^{r}p_{n}\|_{L_{\rho}^{2}}^{2}\right)^{{1/2}}}\geq c_{\alpha}n^{-\alpha/2},$
 for some constant $c_{\alpha}>0$ independent of $N$. From
 $n\leq\#\left(\sum_{\begin{subarray}{c}(\ell_{1},\dots,\ell_{d})\in\Lambda\\ \ell_{2}=\dotsb=\ell_{d}=1\end{subarray}}\Delta_{\ell_{1}}\otimes\dotsb\otimes\Delta_{\ell_{d}}\right)\leq\#\left(\sum_{(\ell_{1},\dots,\ell_{d})\in\Lambda}\Delta_{\ell_{1}}\otimes\dotsb\otimes\Delta_{\ell_{d}}\right)=N,$
-the statement follows. ∎
+where #(Q)\#(Q) denotes the total number of quadrature nodes used in a quadrature rule $Q$, the statement follows.
 For the particular choice of index set
-$\Lambda_{L}=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid|\boldsymbol{\ell}|_{1}\leq L\},$
+$\Lambda_{L}=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid|{\boldsymbol{\ell}|}\leq L\},\quad{\text{where }\ |\boldsymbol{\ell}|=\ell_{1}+\dots+\ell_{d},}$
 we establish a lower bound of a slower decay, now including a logarithmic factor that grows with the dimension; see Theorem 3.3.
@@
 Note that $S_{\Lambda_{L}}=S_{\Lambda_{L}^{0}}$, where
-$\Lambda_{L}^{0}=\{\boldsymbol{\ell}\in\mathbb{N}_{0}^{d}\mid|\boldsymbol{\ell}|_{1}\leq L\}$
+$\Lambda_{L}^{0}=\{\boldsymbol{\ell}\in\mathbb{N}_{0}^{d}\mid|{\boldsymbol{\ell}|}\leq L\}$
 because $\Delta_{0}=0$.
@@
 Theorem 3.3.
-Let $S_{\Lambda_{L}}$ be the sparse-grid quadrature based on the Gauss–Hermite rule for $\Lambda_{L}=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid|\boldsymbol{\ell}|_{1}\leq L\}$.
+Let $S_{\Lambda_{L}}$ be the sparse-grid quadrature based on the Gauss–Hermite rule for $\Lambda_{L}=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid|{\boldsymbol{\ell}|}\leq L\}$.
 Suppose that the Gauss–Hermite rule $Q_{\ell}^{\mathrm{uni}}$ for univariate functions employs $n_{\ell}$ quadrature points for every $\ell\geq 1$, and that for some $M_{0}>0$ and $M>1$ we choose $n_{\ell}$ as
@@
 where the constant $c_{\alpha,d}>0$ is independent of $N$.
-Proof.
+Proof 3.4.
 Define $h_{\Lambda_{L}}(\boldsymbol{x}):=p_{n_{L-d+1}}(x_{1})$ for $\boldsymbol{x}=(x_{1},\dots,x_{d})\in\mathbb{R}^{d}$. From the proof of Theorem 3.1 we know
@@
 so it remains to relate NN and $n_{L-d+1}$.
-Given the choice $\Lambda_{L}=\bigl\{\boldsymbol{\ell}=(\ell_{1},\dots,\ell_{d})\in\mathbb{N}^{d}:|\boldsymbol{\ell}|_{1}\leq L\bigr\}$, and in view of the well-known representation formula of the Smolyak algorithm (see e.g., [54, Lemma 15.1]) the total number of point evaluation $N$ can be written explicitly as
-$N=\sum_{\boldsymbol{\ell}\in\Lambda_{L}}\prod_{k=1}^{d}n_{\ell_{k}}.$ (6)
+Given the choice $\Lambda_{L}=\bigl\{\boldsymbol{\ell}=(\ell_{1},\dots,\ell_{d})\in\mathbb{N}^{d}:|{\boldsymbol{\ell}|}\leq L\bigr\}$, and in view of the well-known representation formula of the Smolyak algorithm (see e.g., [undefaad, Lemma 15.1]) the total number of point evaluation $N$ can be written explicitly as
+(6) $N=\sum_{\boldsymbol{\ell}\in\Lambda_{L}}\prod_{k=1}^{d}n_{\ell_{k}}.$
 Hence $M^{\ell-1}\leq n_{\ell}\leq M_{0}(M^{\ell}-1)$ implies
-$M^{-d}\sum_{\boldsymbol{\ell}\in\Lambda_{L}}M^{|\boldsymbol{\ell}|_{1}}\leq N\leq M_{0}^{d}\sum_{\boldsymbol{\ell}\in\Lambda_{L}}M^{|\boldsymbol{\ell}|_{1}},$
-and thus together with $\sum_{\boldsymbol{\ell}\in\Lambda_{L}}M^{|\boldsymbol{\ell}|_{1}}=\sum_{k=d}^{L}M^{k}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{N}^{d}\\ |\boldsymbol{\ell}|_{1}=k\end{subarray}}1$ and $\{\boldsymbol{\ell}\in\mathbb{N}^{d}:|\boldsymbol{\ell}|_{1}=k\}=\frac{(k-1)!}{(k-d)!(d-1)!}=\Theta_{d}(k^{d-1})$ we obtain
-$\underline{C}_{d}M^{L}L^{d-1}\leq\underline{C}^{\prime}_{d}M^{L}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{N}^{d}\\ |\boldsymbol{\ell}|_{1}=L\end{subarray}}1\leq N\leq\overline{C}_{d}L^{d-1}\sum_{k=0}^{L}M^{k}\leq\frac{\overline{C}_{d}M}{M-1}L^{d-1}M^{L}.$ (7)
+$M^{-d}\sum_{\boldsymbol{\ell}\in\Lambda_{L}}M^{|{\boldsymbol{\ell}|}}\leq N\leq M_{0}^{d}\sum_{\boldsymbol{\ell}\in\Lambda_{L}}M^{|{\boldsymbol{\ell}|}},$
+and thus together with $\sum_{\boldsymbol{\ell}\in\Lambda_{L}}M^{|{\boldsymbol{\ell}|}}=\sum_{k=d}^{L}M^{k}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{N}^{d}\\ |{\boldsymbol{\ell}|}=k\end{subarray}}1$ and $\{\boldsymbol{\ell}\in\mathbb{N}^{d}:|{\boldsymbol{\ell}|}=k\}=\frac{(k-1)!}{(k-d)!(d-1)!}=\Theta_{d}(k^{d-1})$ we obtain
+(7) $\underline{C}_{d}M^{L}L^{d-1}\leq\underline{C}^{\prime}_{d}M^{L}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{N}^{d}\\ |{\boldsymbol{\ell}|}=L\end{subarray}}1\leq N\leq\overline{C}_{d}L^{d-1}\sum_{k=0}^{L}M^{k}\leq\frac{\overline{C}_{d}M}{M-1}L^{d-1}M^{L}.$
 Here, the implied constants depend on $d$. Hence, $N\asymp_{d}n_{L}(\ln n_{L})^{d-1}\asymp_{d}n_{L-d+1}(\ln n_{L-d+1})^{d-1}$.
@@
 $n_{L-d+1}^{-\alpha/2}\asymp_{d}N^{-\alpha/2}(\ln n_{L-d+1})^{\frac{a}{2}(d-1)}\gtrsim_{d}N^{-\alpha/2}(\ln N)^{\frac{a}{2}(d-1)},$
-which completes the proof. ∎
+which completes the proof.
 These bounds turn out to be sharp in the sense that for the standard isotropic Smolyak’s algorithm we have a matching upper bound, as we will see in Section 3.3.
@@
 $\sup_{f\neq 0}\frac{\|f-\mathscr{I}_{\Lambda_{L}}(f)\|_{L_{\rho}^{1}}}{\|f\|_{H_{\rho}^{\alpha}}}\geq c_{\alpha,d}N^{-\alpha/2}(\ln N)^{\frac{a}{2}(d-1)}.$
-Proof.
+Proof 3.6.
 Without loss of generality, we assume that $\Lambda$ is downward-closed in the first coordinate.
@@
 $\frac{\bigl|\int_{\mathbb{R}}p_{n}(x)\rho(x)\mathrm{d}x-Q_{\overline{\ell}_{1}}^{\mathrm{uni}}(p_{n})\bigr|}{\left(\sum_{r=0}^{\alpha}\|D^{r}p_{n}\|_{L_{\rho}^{2}(\mathbb{R})}^{2}\right)^{\frac{1}{2}}}$ $\leq\frac{\|h_{\Lambda}-\mathscr{I}_{\Lambda}(h_{\Lambda})\|_{L_{\rho}^{1}(\mathbb{R}^{d})}}{\|h_{\Lambda}\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}}.$
-Following the arguments in the proofs of Theorems 3.1 and 3.2, the statement follows. ∎
+Following the arguments in the proofs of Theorems 3.1 and 3.3, the statement follows.
 3.3 Upper bound
 In this section, we use the index set
-$\Lambda_{L}:=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid\,|\boldsymbol{\ell}|_{1}\leq L\}.$
+$\Lambda_{L}:=\{\boldsymbol{\ell}\in\mathbb{N}^{d}\mid\,|{\boldsymbol{\ell}|}\leq L\}.$
 As mentioned earlier, this choice corresponds to the classical sparse-grid quadrature, also known as the isotropic sparse grid or the classical Smolyak sparse grid.
-To derive an upper bound, we will invoke the general theory of Smolyak algorithm [65]; see also Smolyak’s foundational work [61].
+To derive an upper bound, we will invoke the general theory of Smolyak algorithm [undefaao]; see also Smolyak’s foundational work [undefaak].
 The theory therein is for general approximation problem of tensor product of continuous linear functionals defined on the tensor product of a reproducing kernel Hilbert space.
@@
 Instead, we will directly verify that an analogous result holds in $H_{\rho}^{\alpha}$.
-For notational convenience, define $e(S_{\Lambda}(f)):=\sup_{f\neq 0}\frac{|I(f)-S_{\Lambda}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}.$ The following result is essentially by Wasilkowski and Woźniakowski [65].
+For notational convenience, define $e(S_{\Lambda}(f)):=\sup_{f\neq 0}\frac{|I(f)-S_{\Lambda}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}.$ The following result is essentially by Wasilkowski and Woźniakowski [undefaao].
 Proposition 3.7.
@@
 $e(S_{\Lambda_{L}}(f))\leq C_{2}\,M^{-\alpha L/2}\binom{L}{d-1}$
-with $C_{2}:=C_{1}\Big(\max\Big\{M^{\alpha/2},c_{\alpha}M^{\alpha/2}+c_{\alpha}M^{\alpha}\Big\}\Big)^{d-1}$.
-Proof.
-We closely follow the argument in [65, Proof of Lemma 2].
-The same reasoning applies if the functional norm on $H_{\rho}^{\alpha}(\mathbb{R}^{d})$ is bounded by the product of functional norms on $H_{\rho}^{\alpha}(\mathbb{R})$.
-Since such a bound is established in Section 2, the result follows.
-∎
+with $C_{2}:=C_{1}\Bigl(\max\bigl\{M^{\alpha/2}\,,\,{C_{1}(1+M^{\alpha/2})}\bigr\}\Bigr)^{d-1}$.
+Proof 3.8.
+We closely follow the argument in [undefaao, Proof of Lemma 2].
+The same reasoning applies if bounded linear functionals on $H_{\rho}^{\alpha}(\mathbb{R}^{s})$ with $\rho(\boldsymbol{x})=\frac{\mathrm{e}^{-|\boldsymbol{x}|_{2}^{2}/2}}{(2\pi)^{s/2}}$ of the form $T_{1}\otimes\dotsb\otimes T_{s}$, where each $T_{j}\colon H_{\rho}^{\alpha}(\mathbb{R})\to\mathbb{R}$ is a bounded linear funcitonal on $H_{\rho}^{\alpha}(\mathbb{R})$ with $\rho(x)=\frac{\mathrm{e}^{-|x|^{2}/2}}{(2\pi)^{1/2}}$, satisfies $\|T_{1}\otimes\dotsb\otimes T_{s}\|_{H_{\rho}^{\alpha}(\mathbb{R}^{s})\to\mathbb{R}}=\prod_{j=1}^{s}\|T_{j}\|_{H_{\rho}^{\alpha}(\mathbb{R})\to\mathbb{R}}$ for $s=1,\dots,d$.
+Since such an equality holds as per Section 2.2, the result follows.
 Then, the final piece for deriving the numerical integration error is an error analysis in the space that gives $H_{\rho}^{\alpha}$ upon taking the tensor product.
-Such result is obtained in our previous paper [41].
+Such result is obtained in our previous paper [undefap].
 Proposition 3.9.
@@
 where the constant $c_{\alpha,d}$ is independent of $f$.
-Proof.
-From [41, Proposition 3.4], we know that the worst-case error in the Sobolev space for univariate functions is given by
+Proof 3.10.
+From [undefap, Proposition 3.4], we know that the worst-case error in the Sobolev space for univariate functions is given by
 $\sup_{g\neq 0}\frac{\big|\int_{\mathbb{R}}g(x)\rho(x)\mathrm{d}x-Q_{n_{\ell}}^{\mathrm{uni}}(g)\big|}{\|g\|_{H_{\rho}^{\alpha}}}\leq c_{\alpha}n_{\ell}^{-\alpha/2}\leq C_{1}\,M^{-\alpha\ell/2}$
-with $C_{1}:=c_{\alpha}M^{\alpha/2}$. This bound enables us to invoke Proposition 3.4, yielding
+with a constant $c_{\alpha}$ depending on $\alpha$, and $C_{1}:=c_{\alpha}M^{\alpha/2}$. This bound enables us to invoke Proposition 3.7, yielding
 $e(S_{\Lambda_{L}}(f))\leq C_{2}\,M^{-\alpha L/2}\binom{L}{d-1}$
@@
 The lower bound $M^{L-d}\leq n_{L-d+1}\leq N$ can be obtained by considering the contribution for the quadrature corresponding to the index $(\ell_{1},\dots,\ell_{d})$ with $\ell_{1}=L-d+1$ and $\ell_{2}=\dotsb=\ell_{d}=1$.
-On the other hand, following the calculations in [65, Section 5] we get $N\leq M_{0}^{d}\frac{M}{M-1}M^{L}\binom{L-1}{d-1}$.
+On the other hand, following the calculations in [undefaao, Section 5] we get $N\leq M_{0}^{d}\frac{M}{M-1}M^{L}\binom{L-1}{d-1}$.
 From $\binom{L-1}{d-1}=\big(1-\frac{d-1}{L}\big)\binom{L}{d-1}\leq\binom{L}{d-1}$
@@
 $\leq\frac{C_{2}}{((d-1)!)^{1+\alpha/2}}\left(\frac{M_{0}^{d}M}{M-1}\right)^{\alpha/2}N^{-\alpha/2}L^{(d-1)(1+\alpha/2)}.$
-[…]
+To bound the factor of $L$ in terms of $\ln N$, note that $2\leq L-d+1$ implies $M\leq n_{2}\leq N$; combined with $M^{L-d}\leq N$, this gives $M^{L}\leq M^{d}N\leq N^{1+d}$ and thus $L\leq{(}\ln N{)}(1+d)/\ln M$.
+Therefore, we obtain
 $e(S_{\Lambda_{L}}(f))\leq\frac{C_{2}((1+d)/\ln M)^{(d-1)(1+\alpha/2)}}{((d-1)!)^{1+\alpha/2}}\left(\frac{M_{0}^{d}M}{M-1}\right)^{\alpha/2}\,N^{-\alpha/2}(\ln N)^{(d-1)(1+\alpha/2)}.$
-Hence, the statement follows with $c_{\alpha,d}:=\frac{C_{2}((1+d)/\ln M)^{(d-1)(1+\alpha/2)}}{((d-1)!)^{1+\alpha/2}}\left(\frac{M_{0}^{d}M}{M-1}\right)^{\alpha/2}$. ∎
+Hence, the statement follows with $c_{\alpha,d}:=\frac{C_{2}((1+d)/\ln M)^{(d-1)(1+\alpha/2)}}{((d-1)!)^{1+\alpha/2}}\left(\frac{M_{0}^{d}M}{M-1}\right)^{\alpha/2}$.
 4 Quasi-Monte Carlo methods
@@
 with nodes $\boldsymbol{t}^{(j)}\in[0,1]^{d}$, $j=1,\dots,N$.
-Popular classes for the choice of $\boldsymbol{t}_{j}$ include lattice points [60], higher-order digital nets [14], and low discrepancy points/sequences [52, Chapter 3].
-These classes are not mutually exclusive: lattice points and digital nets may also have low discrepancy; see, for instance, [12, Chapter 5] and [14, Chapter 5].
+Popular classes for the choice of $\boldsymbol{t}_{j}$ include lattice points [undefaab, undefaaj, undefl], (higher-order) digital nets [undefaab, undefn], and low discrepancy points/sequences (such as Halton and Kronecker sequences) [undefaab, Chapter 3].
+These classes are not mutually exclusive: lattice points and digital nets may also have low discrepancy; see, for instance, [undefl, Chapter 5] and [undefn, Chapter 5].
 An important point to mention is that if one views $\{\boldsymbol{t}_{j}\}$ as a low-discrepancy point set and uses the Koksma–Hlawka inequality as a framework to bound the QMC error, then the best achievable rate is linear convergence.
@@
 QMC points are typically defined on the unit cube $0,1]^{d}$. To integrate functions over $\mathbb{R}^{d}$, we therefore apply domain truncation and a change of variables:
-$\int_{\mathbb{R}^{d}}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}$ $=\int_{\Omega}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}+\int_{\mathbb{R}^{d}\setminus\Omega}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}$ (8)
+(8) $\int_{\mathbb{R}^{d}}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}$ $=\int_{\Omega}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}+\int_{\mathbb{R}^{d}\setminus\Omega}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}$
 and
-$\int_{\Omega}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}$ $=\int_{[0,1]^{d}}f(\boldsymbol{\Psi}(\boldsymbol{t}))\,\rho(\boldsymbol{\Psi}(\boldsymbol{t}))\,|\det(D\boldsymbol{\Psi}(\boldsymbol{t}))|\,\mathrm{d}\boldsymbol{t}$
-$\approx\frac{1}{N}\sum_{j=1}^{N}f\bigl(\boldsymbol{\Psi}(\boldsymbol{t}^{(j)})\bigr)\rho\bigl(\boldsymbol{\Psi}(\boldsymbol{t}^{(j)})\bigr)|\det(D\boldsymbol{\Psi}(\boldsymbol{t}^{(j)}))|=:Q_{N}(f).$ (9)
-Here, $\Omega\subseteq\mathbb{R}^{d}$ is a suitable integration domain, possibly $\Omega=\mathbb{R}^{d}$.
+$\int_{\Omega}f(\boldsymbol{x})\,\rho(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}$ $=\int_{{(}0,1{)}^{d}}f(\boldsymbol{\Psi}(\boldsymbol{t}))\,\rho(\boldsymbol{\Psi}(\boldsymbol{t}))\,|\det(D\boldsymbol{\Psi}(\boldsymbol{t}))|\,\mathrm{d}\boldsymbol{t}$
+(9) $\approx\frac{1}{N}\sum_{j=1}^{N}f\bigl(\boldsymbol{\Psi}(\boldsymbol{t}^{(j)})\bigr)\rho\bigl(\boldsymbol{\Psi}(\boldsymbol{t}^{(j)})\bigr)|\det(D\boldsymbol{\Psi}(\boldsymbol{t}^{(j)}))|=:Q_{N}(f).$
+Here, $\Omega\subseteq\mathbb{R}^{d}$ is a suitable integration domain, possibly $\Omega=\mathbb{R}^{d}$, and $\boldsymbol{\Psi}:(0,1)^{d}\to\Omega$ is a suitable mapping.
 Often the mapping is component-wise, $\boldsymbol{\Psi}(\boldsymbol{t})\!\!\;=\!\!\;(\Psi_{1}(t_{1}),\dots,\Psi_{d}(t_{d}))$, and non-negative, in which case the Jacobian factor simplifies as
 $|\det(D\boldsymbol{\Psi}(\boldsymbol{t}))|=\prod_{k=1}^{d}\Psi_{k}^{\prime}(t_{k}).$
-In this paper, we consider two choices for $\Psi$: an affine map and a Möbius transformation.
+In this paper, we consider two choices for $\Psi$: an affine map and a Möbius transformation, namely the cotangent transform.
 For the former, $\Omega$ is a suitable box; for the latter, $\Omega$ is the whole $\mathbb{R}^{d}$.
@@
 where $\{\cdot\}$ expresses the fractional part of each component, and the integer vector $\boldsymbol{z}\in\mathbb{N}^{d}$ is called a generating vector.
-A good generating vector can be found efficiently using a greedy algorithm known as the component-by-component (CBC) construction, which provably yields a vector with a satisfactorily small worst-case error in various function spaces [12].
-A digital net is defined as follows. For a fixed prime $p$ and a positive integer $m$, we denote the $p$-adic expansion of a non-negative integer $i<p^{m}$ and a real number $x\in[0,1)$ by
-$i=\iota_{0}+\iota_{1}p+\cdots\iota_{m-1}p^{m-1}\qquad\text{and}\qquad x=\frac{\xi_{1}}{p}+\frac{\xi_{2}}{p^{2}}+\cdots,$
+A good generating vector can be found efficiently using a greedy algorithm known as the component-by-component (CBC) construction, which provably yields a vector with a satisfactorily small worst-case error in various function spaces [undefl].
+A digital net is defined as follows. For a fixed prime $p$ and a positive integer $m$, we denote the $p$-adic expansion of a non-negative integer ${j}<p^{m}$ and a real number $x\in[0,1)$ by
+${j=\iota_{0}+\iota_{1}p+\cdots+\iota_{m-1}p^{m-1}}\qquad\text{and}\qquad x=\frac{\xi_{1}}{p}+\frac{\xi_{2}}{p^{2}}+\cdots,$
 respectively.
-We then write $\vec{i}=(\iota_{0},\iota_{1},\ldots,\iota_{m-1})^{\top}\in\mathbb{Z}_{p}^{m}$ and $\vec{x}=(\xi_{1},\xi_{2},\ldots)^{\top}\in\mathbb{Z}_{p}^{\mathbb{N}}$.
-Here, $\mathbb{Z}_{p}$ is the set $\{0,\dots,p-1\}$ equipped with the usual arithmetic operations modulo $p$.
-If $x$ admits two different $p$-adic expansions, we choose the one in which infinitely many of the digits $\xi_{i}$ differ from p1p-1.
-Now, for suitably designed generating matrices $C_{1},\ldots,C_{d}\in\mathbb{Z}_{p}^{\mathbb{N}\times m}$, a digital net consisting of $N=p^{m}$ points is given by
-$\vec{t}_{j,i}=C_{i}\,\overrightarrow{(j-1)}\;\;\text{ for }i=1,\ldots,d\text{ and }j=1,\ldots,N.$
+Here, the digits $\iota_{k}$ and $\xi_{k}$ are in $\mathbb{Z}_{p}$, the set $\{0,1,\dots,p-1\}$ equipped with the usual arithmetic operations modulo $p$.
+We then write $\vec{j}=(\iota_{0},\iota_{1},\ldots,\iota_{m-1})^{\top}\in\mathbb{Z}_{p}^{m}$ and $\vec{x}=(\xi_{1},\xi_{2},\ldots)^{\top}\in\mathbb{Z}_{p}^{\mathbb{N}}$.
+If $x$ admits two different $p$-adic expansions, we choose the one in which infinitely many of the digits $\xi_{k}$ differ from p1p-1.
+Now, for suitably designed generating matrices $C_{1},\ldots,C_{d}\in\mathbb{Z}_{p}^{\mathbb{N}\times m}$, consider the corresponding digit vectors $\vec{t}_{j,i}\in\mathbb{Z}^{\mathbb{N}}_{p}$ obtained by the matrix-vector product modulo $p$:
+$\vec{t}_{j,i}=C_{i}\,\vec{j}\;\;\text{ for }i=1,\ldots,d\text{ and }j=0,1,\ldots,N-1.$
+Then, the digital net consisting of $N=p^{m}$ points corresponding to $C_{1},\dots,C_{d}$ is defined as the set $\{\boldsymbol{t}^{(0)},\ldots,\boldsymbol{t}^{(N-1)}\}$, where each component $t_{j,i}$ of the point $\boldsymbol{t}^{(j)}=(t_{j,1},\ldots,t_{j,d})\in[0,1]^{d}$ is determined by the corresponding digit vector $\vec{t}_{j,i}$.
 Note that if every column of the generating matrices has only finitely many non-zero entries, this matrix-vector product is well-defined.
-The classical (first-order) digital nets constructed by Sobol’ [62], Niederreiter [50], and Niederreiter–Xing [51] are included as special cases of this construction.
-In [10], Dick introduced the notion of higher-order digital nets, which also form a special class within this framework.
+The classical (first-order) digital nets constructed by Sobol’ [undefaal], Niederreiter [undefaz], and Niederreiter–Xing [undefaaa] are included as special cases of this construction.
+In [undefj], Dick introduced the notion of higher-order digital nets, which also form a special class within this framework.
 An order-$\alpha$ digital net with $\alpha\in\mathbb{N}$, $\alpha\geq 2$, is constructed using generating matrices $C_{1},\ldots,C_{d}$, which are themselves derived from the generating matrices $D_{1},\ldots,D_{\alpha d}$ for first-order digital nets in dimension $\alpha d$.
-We refer the reader to [28] for a review of higher-order nets.
+We refer the reader to [undefab] for a review of higher-order nets.
 It is worth noting that QMC methods presented in this section always have positive weights, i.e., when seeing the quadrature (9) as $Q_{N}(f)=\sum_{j=1}^{N}w_{j}f(\boldsymbol{x}_{j})$, weights $w_{j}$’s are all positive.
@@
 This nonnegativity-preserving property is important in situations where the quantity of interest must be nonnegative, for example probabilities.
-Positive quadrature weights are also useful when discretizing optimization problems involving expectations [46] with convex objective functions, since they preserve convexity.
+Positive quadrature weights are also useful when discretizing optimization problems involving expectations [undefav] with convex objective functions, since they preserve convexity.
 By contrast, other quadrature methods may involve negative weights and thus lose these desirable features.
+Before going into details, we first summarize our main results of this section in the following theorem.
+Theorem 4.1.
+The quasi-Monte Carlo rule $Q_{N}$ (9) achieves the error bound
+$\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}$ $\leq C_{d,\alpha}\frac{(\ln N)^{\gamma}}{N^{\alpha}},$
+where the exponent $\gamma$ depends on the QMC construction:
+[…]
+The polynomial convergence rate $\alpha$ is optimal in $H_{\rho}^{\alpha}$: no algorithm can achieve a better polynomial rate.
+Furthermore, the exponent (d1)/2(d-1)/2 in the logarithmic factor is also optimal.
+The bound for higher-order digital nets with an affine map is due to Dick et al.
+[undefk].
+In the following subsections, we present detailed algorithms and establish the remaining upper bounds.
+The optimality statements then follow from the lower bound in Theorem 2.1.
 4.1 Affinely-mapped QMC
@@
 Choosing $\boldsymbol{b},\boldsymbol{b}]$ as $\Omega$ in (8) and using $\boldsymbol{\Psi}_{\mathrm{affine}}$ in (9) yields
-$Q_{N}(f)=\frac{(2b)^{d}}{N}\sum_{j=1}^{N}f\bigl(\boldsymbol{\Psi}_{\mathrm{affine}}(\boldsymbol{t}^{(j)})\bigr)\rho\bigl(\boldsymbol{\Psi}_{\mathrm{affine}}(\boldsymbol{t}^{(j)})\bigr),\quad f\in H_{\rho}^{\alpha}.$ (10)
+(10) $Q_{N}(f)=\frac{(2b)^{d}}{N}\sum_{j=1}^{N}f\bigl(\boldsymbol{\Psi}_{\mathrm{affine}}(\boldsymbol{t}^{(j)})\bigr)\rho\bigl(\boldsymbol{\Psi}_{\mathrm{affine}}(\boldsymbol{t}^{(j)})\bigr),\quad f\in H_{\rho}^{\alpha}.$
 The value $b$ controls the truncation error, namely the integral of $f\rho$ over $\mathbb{R}^{d}\setminus[-\boldsymbol{b},\boldsymbol{b}]$, whereas $Q_{N}$ controls the numerical integration error.
@@
 Proposition 4.2 (Affinely-mapped lattice).
-Let $\alpha\in\mathbb{N}$, and let $Q_{N}$ be the rank-$1$ lattice rule defined in (10) with $b=\sqrt{\frac{\alpha}{2}\ln N}$.
-Then, with a generating vector $\boldsymbol{z}$ obtained by the CBC construction in [12, Algorithm 3.14], $Q_{N}$ satisfies
-$\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}\leq C_{\alpha,d}N^{-\alpha}(\ln N)^{\frac{3\alpha d}{2}+\frac{d}{4}}.$ (11)
-Proof.
-First, we will show that $f\in H_{\rho}^{\alpha}$ implies $\|f\rho\|_{\text{unanchored}}+\|f\rho\|_{\text{decay}}\leq C_{d,\alpha}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$ for some constant $C_{d,\alpha}>0$, where
-$\|f\rho\|_{\text{decay}}\coloneqq\sup_{\begin{subarray}{c}\boldsymbol{x}\in\mathbb{R}^{d}\\ \mathbf{r}\in\{0,\ldots,\alpha-1\}^{d}\end{subarray}}\left|\rho^{-\varepsilon}(\boldsymbol{x})\,D^{\mathbf{r}}(f\rho)(\boldsymbol{x})\right|<\infty$
-for all $\varepsilon\in(0,1/2)$, and
+Let $\alpha\in\mathbb{N}$, and let $Q_{N}$ be the rank-$1$ lattice rule defined in (10) with $b={(2+\eta)}\sqrt{\alpha\ln N}$, where $\eta>0$ is arbitrary.
+Then, with a generating vector $\boldsymbol{z}$ obtained by the CBC construction in [undefl, Algorithm 3.14], $Q_{N}$ satisfies
+(11) $\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}\leq C_{\alpha,d,{\eta}}\;\!N^{-\alpha}(\ln N)^{\frac{3\alpha d}{2}+\frac{d}{4}}.$
+Proof 4.3.
+First, we will show that $f\in H_{\rho}^{\alpha}$ implies $\|f\rho\|_{\text{unanchored}}+\|f\rho\|_{\text{decay}{,\varepsilon}}\leq C_{d,\alpha,{\varepsilon}}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$ for every $\varepsilon\in(0,1/2)$ with some constant $C_{d,\alpha{,\varepsilon}}>0$, where
+$\|f\rho\|_{\text{decay}{,\varepsilon}}\coloneqq\sup_{\begin{subarray}{c}\boldsymbol{x}\in\mathbb{R}^{d}\\ \mathbf{r}\in\{0,\ldots,\alpha-1\}^{d}\end{subarray}}\left|\rho^{-\varepsilon}(\boldsymbol{x})\,D^{\mathbf{r}}(f\rho)(\boldsymbol{x})\right|<\infty,$
+and
 $\|f\rho\|_{\text{unanchored}}\coloneqq\sup_{[\boldsymbol{a},\boldsymbol{b}]\subset\mathbb{R}^{d}}\|f\rho\|_{\mathscr{H}^{\alpha}([\boldsymbol{a},\boldsymbol{b}])}<\infty$
 with
-$\langle f,h\rangle_{\mathscr{H}^{\alpha}([\boldsymbol{a},\boldsymbol{b}])}$ $:=\sum_{\begin{subarray}{c}\mathbf{r}\in\{0,\dots,\alpha\}^{d}\\ \mathfrak{v}:=\{k:r_{k}=\alpha\}\end{subarray}}\int_{[\boldsymbol{a}_{\mathfrak{v}},\boldsymbol{b}_{\mathfrak{v}}]}\left(\int_{[\boldsymbol{a}_{-\mathfrak{v}},\boldsymbol{b}_{-\mathfrak{v}}]}D^{\mathbf{r}}f(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)$ (12)
+(12) $\langle f,h\rangle_{\mathscr{H}^{\alpha}([\boldsymbol{a},\boldsymbol{b}])}$ $:=\sum_{\begin{subarray}{c}\mathbf{r}\in\{0,\dots,\alpha\}^{d}\\ \mathfrak{v}:=\{k:r_{k}=\alpha\}\end{subarray}}\int_{[\boldsymbol{a}_{\mathfrak{v}},\boldsymbol{b}_{\mathfrak{v}}]}\left(\int_{[\boldsymbol{a}_{-\mathfrak{v}},\boldsymbol{b}_{-\mathfrak{v}}]}D^{\mathbf{r}}f(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)$
 $\qquad\qquad\qquad\qquad\times\left(\int_{[\boldsymbol{a}_{-\mathfrak{v}},\boldsymbol{b}_{-\mathfrak{v}}]}D^{\mathbf{r}}h(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)\mathrm{d}\boldsymbol{x}_{\mathfrak{v}}$
-for bounded $\boldsymbol{a},\boldsymbol{b}]$. In the proof we use the following identities
+for bounded $\boldsymbol{a},\boldsymbol{b}]$.
+The condition $\|f\rho\|_{\text{decay},\varepsilon}<\infty$ corresponds to the exponential decay condition [undefaae, Equation (22)], whereas $\|f\rho\|_{\text{unanchored}}<\infty$ corresponds to the requirement that [undefaae, Equation (24)] be finite.
+In the proof we use the following identities
 $D^{\mathbf{r}}(H_{\mathbf{k}}\,\rho)$ $=(-1)^{|\mathbf{r}|}\sqrt{\frac{(\mathbf{k}+\mathbf{r})!}{\mathbf{k}!}}H_{\mathbf{k}+\mathbf{r}}\,\rho,\ \text{ and }\ D^{\mathbf{r}}\rho=(-1)^{|\mathbf{r}|}\sqrt{\mathbf{r}!}\,H_{\mathbf{r}}\,\rho,$
@@
 For $g=f\rho$ and $\mathbf{r}\in\mathbb{N}_{0}^{d}$ with $|\mathbf{r}|_{\infty}\leq\alpha$, the product rule yields
-$D^{\mathbf{r}}g$ $=\sum_{\mathbf{s}\leq\mathbf{r}}\binom{\mathbf{r}}{\mathbf{s}}(D^{\mathbf{s}}\rho)\,(D^{\mathbf{r}-\mathbf{s}}f)=\sum_{\mathbf{s}\leq\mathbf{r}}(-1)^{|\mathbf{s}|}\sqrt{\mathbf{s}!}\binom{\mathbf{r}}{\mathbf{s}}\,H_{\mathbf{s}}\,\rho\,D^{\mathbf{r}-\mathbf{s}}f.$ (13)
+(13) $D^{\mathbf{r}}g$ $=\sum_{\mathbf{s}\leq\mathbf{r}}\binom{\mathbf{r}}{\mathbf{s}}(D^{\mathbf{s}}\rho)\,(D^{\mathbf{r}-\mathbf{s}}f)=\sum_{\mathbf{s}\leq\mathbf{r}}(-1)^{|\mathbf{s}|}\sqrt{\mathbf{s}!}\binom{\mathbf{r}}{\mathbf{s}}\,H_{\mathbf{s}}\,\rho\,D^{\mathbf{r}-\mathbf{s}}f.$
 Now,
-$\sup_{[\boldsymbol{a},\boldsymbol{b}]\subset\mathbb{R}^{d}}\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{a},\boldsymbol{b}])}^{2}\leq\sum_{\begin{subarray}{c}\mathbf{r}\in\{0,\ldots,\alpha\}^{d}\\ \mathfrak{v}:=\left\{j:r_{j}=\alpha\right\}\end{subarray}}\int_{\mathbb{R}^{|\mathfrak{v}|}}\left(\int_{\mathbb{R}^{d-|\mathfrak{\mathfrak{v}}|}}|D^{\mathbf{r}}g(\boldsymbol{x})|\,\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)^{2}\mathrm{d}\boldsymbol{x}_{\mathfrak{v}}.$ (14)
+(14) $\sup_{[\boldsymbol{a},\boldsymbol{b}]\subset\mathbb{R}^{d}}\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{a},\boldsymbol{b}])}^{2}\leq\sum_{\begin{subarray}{c}\mathbf{r}\in\{0,\ldots,\alpha\}^{d}\\ \mathfrak{v}:=\left\{j:r_{j}=\alpha\right\}\end{subarray}}\int_{\mathbb{R}^{|\mathfrak{v}|}}\left(\int_{\mathbb{R}^{d-|\mathfrak{\mathfrak{v}}|}}|D^{\mathbf{r}}g(\boldsymbol{x})|\,\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)^{2}\mathrm{d}\boldsymbol{x}_{\mathfrak{v}}.$
 Using the Cauchy–Schwarz inequality, the inner-integral can be bounded as
@@
 with $\|H_{\boldsymbol{s}(-\mathfrak{v})}\|_{L^{2}_{\rho_{\mathfrak{-\mathfrak{v}}}}}:=\left(\int_{\mathbb{R}^{d-|\mathfrak{\mathfrak{v}}|}}|H_{\boldsymbol{s}(-\mathfrak{v})}(\boldsymbol{x}_{-\mathfrak{v}})|^{2}\,\rho_{\mathfrak{-\mathfrak{v}}}(\boldsymbol{x}_{-\mathfrak{v}})\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)^{1/2}$.
-Here, we used the notations $H_{\boldsymbol{s}(\mathfrak{v})}(x_{\mathfrak{v}}):=\prod_{j\in\mathfrak{v}}H_{s_{j}}(x_{j})$, $H_{\boldsymbol{s}(-\mathfrak{v})}(x_{-\mathfrak{v}}):=\prod_{j\in\{1,\dots,d\}\setminus\mathfrak{v}}H_{s_{j}}(x_{j})$, and similarly $\rho_{\mathfrak{v}}:=\prod_{j\in\mathfrak{v}}\rho(x_{j})$ and $\rho_{-\mathfrak{v}}:=\prod_{j\in\{1,\dots,d\}\setminus\mathfrak{v}}\rho(x_{j})$.
+Here, we used the notations $H_{\boldsymbol{s}(\mathfrak{v})}({\boldsymbol{x}}_{\mathfrak{v}}):=\prod_{j\in\mathfrak{v}}H_{s_{j}}(x_{j})$, $H_{\boldsymbol{s}(-\mathfrak{v})}({\boldsymbol{x}}_{-\mathfrak{v}}):=\prod_{j\in\{1,\dots,d\}\setminus\mathfrak{v}}H_{s_{j}}(x_{j})$, and similarly $\rho_{\mathfrak{v}}{(\boldsymbol{x_{\mathfrak{v}}})}:=\prod_{j\in\mathfrak{v}}\rho(x_{j})$ and $\rho_{-\mathfrak{v}}{(\boldsymbol{x_{-\mathfrak{v}}})}:=\prod_{j\in\{1,\dots,d\}\setminus\mathfrak{v}}\rho(x_{j})$.
 Noting $\|H_{\boldsymbol{s}(-\mathfrak{v})}\|_{L^{2}_{\rho_{\mathfrak{-\mathfrak{v}}}}}=1$, it follows
 $\int_{\mathbb{R}^{|\mathfrak{\mathfrak{v}}|}}$ $\left(\int_{\mathbb{R}^{d-|\mathfrak{\mathfrak{v}}|}}|D^{\mathbf{r}}g(\boldsymbol{x})|\,\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)^{2}\mathrm{d}\boldsymbol{x}_{\mathfrak{v}}$
-$\leq C_{d}\sum_{\mathbf{s}\leq\mathbf{r}}\mathbf{s}!\binom{\mathbf{r}}{\mathbf{s}}^{2}\int_{\mathbb{R}^{d}}|H_{\boldsymbol{s}(\mathfrak{v})}(\boldsymbol{x}_{\mathfrak{v}})|^{2}\rho_{\mathfrak{\mathfrak{v}}}^{2}(\boldsymbol{x}_{\mathfrak{v}})\,|D^{\mathbf{r}-\mathbf{s}}f(\boldsymbol{x})|^{2}\,\rho_{\mathfrak{-\mathfrak{v}}}(\boldsymbol{x}_{-\mathfrak{v}})\mathrm{d}\boldsymbol{x}$
-$\leq C_{d}\sum_{\mathbf{s}\leq\mathbf{r}}\mathbf{s}!\binom{\mathbf{r}}{\mathbf{s}}^{2}\sup_{\boldsymbol{y}_{\mathfrak{v}}\in\mathbb{R}^{|\mathfrak{v}|}}|H_{\boldsymbol{s}(\mathfrak{v})}(\boldsymbol{y}_{\mathfrak{v}})|^{2}\rho_{\mathfrak{\mathfrak{v}}}(\boldsymbol{y}_{\mathfrak{v}})\int_{\mathbb{R}^{d}}|D^{\mathbf{r}-\mathbf{s}}f(\boldsymbol{x})|^{2}\,\rho(\boldsymbol{x})\mathrm{d}\boldsymbol{x}.$
+$\leq C_{d,{\alpha}}\sum_{\mathbf{s}\leq\mathbf{r}}\mathbf{s}!\binom{\mathbf{r}}{\mathbf{s}}^{2}\int_{\mathbb{R}^{d}}|H_{\boldsymbol{s}(\mathfrak{v})}(\boldsymbol{x}_{\mathfrak{v}})|^{2}\rho_{\mathfrak{\mathfrak{v}}}^{2}(\boldsymbol{x}_{\mathfrak{v}})\,|D^{\mathbf{r}-\mathbf{s}}f(\boldsymbol{x})|^{2}\,\rho_{\mathfrak{-\mathfrak{v}}}(\boldsymbol{x}_{-\mathfrak{v}})\mathrm{d}\boldsymbol{x}$
+$\leq C_{d,{\alpha}}\sum_{\mathbf{s}\leq\mathbf{r}}\mathbf{s}!\binom{\mathbf{r}}{\mathbf{s}}^{2}\sup_{\boldsymbol{y}_{\mathfrak{v}}\in\mathbb{R}^{|\mathfrak{v}|}}|H_{\boldsymbol{s}(\mathfrak{v})}(\boldsymbol{y}_{\mathfrak{v}})|^{2}\rho_{\mathfrak{\mathfrak{v}}}(\boldsymbol{y}_{\mathfrak{v}})\int_{\mathbb{R}^{d}}|D^{\mathbf{r}-\mathbf{s}}f(\boldsymbol{x})|^{2}\,\rho(\boldsymbol{x})\mathrm{d}\boldsymbol{x}.$
 Therefore, from (14)
-$\sup_{[\boldsymbol{a},\boldsymbol{b}]\subset\mathbb{R}^{d}}\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{a},\boldsymbol{b}])}\leq C_{d,\alpha}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$
+${\|f\rho\|_{\text{unanchored}}=}\sup_{[\boldsymbol{a},\boldsymbol{b}]\subset\mathbb{R}^{d}}\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{a},\boldsymbol{b}])}\leq C_{d,\alpha}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$
 holds for some constant $C_{d,\alpha}>0$.
-To show $\lVert g\rVert_{\text{decay}}\leq C\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$, we will invoke the Sobolev embedding in Proposition 2.3 to bound the supremum norm.
+To show $\lVert g\rVert_{\text{decay}{,\varepsilon}}\leq C\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$, we will invoke the Sobolev embedding in Proposition 2.4 to bound the supremum norm.
 Define $h_{\mathbf{r}}:=\rho^{-\varepsilon}D^{\mathbf{r}}g=\rho^{-\varepsilon}D^{\mathbf{r}}(f\rho)$ for $\mathbf{r}\in\{0,\dots,\alpha-1\}^{d}$, for which we will derive a bound on its Sobolev norm.
@@
 $=\sum_{\mathbf{j}\leq\mathbf{k}}\binom{\mathbf{k}}{\mathbf{j}}\sum_{\mathbf{s}\leq\mathbf{r}+\mathbf{k}-\mathbf{j}}\sqrt{\mathbf{s}!}\binom{\mathbf{r}+\mathbf{k}-\mathbf{j}}{\mathbf{s}}\sup_{\boldsymbol{y}\in\mathbb{R}^{d}}|p_{\mathbf{j}}(\boldsymbol{y})H_{\mathbf{s}}(\boldsymbol{y})\rho^{1/2-\varepsilon}(\boldsymbol{y})|$
-$\qquad\qquad\times\|D^{\mathbf{r}+\mathbf{k}-\mathbf{j}-\mathbf{s}}f\|_{L_{\rho}^{2}(\mathbb{R}^{d})}<\infty,$ (15)
+(15) $\qquad\qquad\times\|D^{\mathbf{r}+\mathbf{k}-\mathbf{j}-\mathbf{s}}f\|_{L_{\rho}^{2}(\mathbb{R}^{d})}<\infty,$
 and we used the identity (13) for $D^{\mathbf{k}-\mathbf{j}+\mathbf{r}}g$. In (15), the multi-index satisfies $\mathbf{r}+\mathbf{k}-\mathbf{j}-\mathbf{s}\in\{0,\dots,\alpha\}^{d}$, and thus
-$\sum_{\mathbf{k}\in\{0,1\}^{d}}\|D^{\mathbf{k}}(\rho^{-\varepsilon}D^{\mathbf{r}}g)\|_{L^{2}(\mathbb{R}^{d})}^{2}=\sum_{\mathbf{k}\in\{0,1\}^{d}}\|D^{\mathbf{k}}h_{\mathbf{r}}\|_{L^{2}(\mathbb{R}^{d})}^{2}$ $\leq C\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}^{2}.$
-Therefore, in view of Proposition 2.3 we have $\sup_{\boldsymbol{x}\in\mathbb{R}^{d}}|\rho^{-\varepsilon}(\boldsymbol{x})D^{\mathbf{r}}g(\boldsymbol{x})|\leq C\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$, hence $\|f\rho\|_{\text{unanchored}}+\|f\rho\|_{\text{decay}}\leq C_{d,\alpha}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$ follows.
-With this, from the first displayed equation in [55, Lemma 5] together with [55, Propositions 7 & 8], if $\boldsymbol{z}$ satisfies the second displayed equation in [55, Lemma 5], then this $\boldsymbol{z}$ satisfies (11); see the proof of [55, Theorem 2].
-Since such $\boldsymbol{z}$ can be constructed using a CBC construction in [12, Algorithm 3.14], the proof is complete.
-∎
-Affinely mapped digital nets also achieve the optimal convergence rate, up to a logarithmic factor, as shown in [11, Corollary 1].
-Proposition 4.2 (Affinely-mapped digital net, [11, Corollary 1]).
+$\sum_{\mathbf{k}\in\{0,1\}^{d}}\|D^{\mathbf{k}}(\rho^{-\varepsilon}D^{\mathbf{r}}g)\|_{L^{2}(\mathbb{R}^{d})}^{2}=\sum_{\mathbf{k}\in\{0,1\}^{d}}\|D^{\mathbf{k}}h_{\mathbf{r}}\|_{L^{2}(\mathbb{R}^{d})}^{2}$ $\leq C_{{\varepsilon}}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}^{2}.$
+Therefore, in view of Proposition 2.4 we have $\sup_{\boldsymbol{x}\in\mathbb{R}^{d}}|\rho^{-\varepsilon}(\boldsymbol{x})D^{\mathbf{r}}g(\boldsymbol{x})|\leq C_{{\varepsilon}}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$, hence $\|f\rho\|_{\text{unanchored}}+\|f\rho\|_{\text{decay}{,\varepsilon}}\leq C_{d,\alpha,{\varepsilon}}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$ follows.
+With this, from the first displayed equation in [undefaae, Lemma 5] together with [undefaae, Propositions 7 & 8], if $\boldsymbol{z}$ satisfies the second displayed equation in [undefaae, Lemma 5], then this $\boldsymbol{z}$ satisfies (11); see the proof of [undefaae, Theorem 2].
+Since such $\boldsymbol{z}$ can be constructed using a CBC construction in [undefl, Algorithm 3.14], we have
+$|I(f)-Q_{N}(f)|\leq C_{\alpha,d,\beta}\|f\|_{H^{\alpha}_{\rho}(\mathbb{R}^{d})}\bigg(\frac{(\ln N)^{\alpha d}}{N^{\alpha}}b^{d(\alpha+1/2)}+b^{d\alpha}\mathrm{e}^{-\beta b^{2}}+b^{d-2}\mathrm{e}^{-\beta b^{2}}\bigg),$
+where $\beta:=\varepsilon/2\in(0,1/4)$ corresponds to $\beta$ in [undefaae], and the constant $C_{\alpha,d,\beta}$ is independent of $N$ and $b$.
+For $\eta>0$ arbitrary, we now choose $b=(2+\eta)\sqrt{\alpha\ln(N)}$.
+Then, since the bound above holds for any $1/(2+\eta)^{2}\leq\beta<1/4$, for such $\beta$ this choice of $b$ ensures $\mathrm{e}^{-\beta b^{2}}=N^{-\beta(2+\eta)^{2}\alpha}\leq N^{-\alpha}$.
+This completes the proof.
+Affinely mapped digital nets also achieve the optimal convergence rate, up to a logarithmic factor, as shown in [undefk, Corollary 1].
+Proposition 4.4 (Affinely-mapped digital net, [undefk, Corollary 1]).
 Let $\alpha\in\mathbb{N}$, and let $Q_{N}$ be the quadrature rule defined in (10) with $b=2\sqrt{{\alpha}\ln N}$, where points $(\boldsymbol{t}^{(j)})_{j=1,\dots,N}$ are given by a higher-order net of order $(2\alpha+1)$ with $N=p^{m}$ elements.
@@
 In the result quoted above, higher-order nets of order $(2\alpha+1)$ are considered.
-On the other hand, as noted in [11], higher-order nets of order $\alpha$ also achieve the optimal polynomial rate $N^{-\alpha}$, but with a logarithmic factor of a larger exponent.
+On the other hand, as noted in [undefk], higher-order nets of order $\alpha$ also achieve the optimal polynomial rate $N^{-\alpha}$, but with a logarithmic factor of a larger exponent.
 4.2 QMC with Möbius transformation
 An alternative to the affine transformation is the change of variables $\Psi:(0,1)^{d}\to\mathbb{R}^{d}$ defined by
-$\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\coloneqq(\phi(t_{1}),\ldots,\phi(t_{d})),\;\;\;\phi(t)\coloneqq-\cot(\pi t).$ (16)
-This is a specific Möbius transformation considered in [64]. For the univariate transformation $\phi$, its derivatives satisfy the following.
+(16) $\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\coloneqq(\phi(t_{1}),\ldots,\phi(t_{d})),\;\;\;\phi(t)\coloneqq-\cot(\pi t).$
+This is a specific Möbius transformation considered in [undefaan]. Using $\boldsymbol{\Psi}_{\mathrm{cot}}$ in (9) yields the following formula:
+(17) $Q_{N}(f)=\frac{1}{N}\sum_{j=1}^{N}f\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t}^{(j)})\right)\rho\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t}^{(j)})\right)\prod_{k=1}^{d}\phi^{\prime}(t^{(j)}_{k}),\quad f\in H_{\rho}^{\alpha}.$
+It turns out that the transformed integrand, defined on the unit cube by $g:=f(\boldsymbol{\Psi}_{\mathrm{cot}}(\cdot))\rho(\boldsymbol{\Psi}_{\mathrm{cot}}(\cdot))\prod_{k=1}^{d}\phi^{\prime}$, belongs to suitable Sobolev spaces for which QMC methods work well.
+To show this, we first derive results on the derivatives of $g$.
+The derivatives of the transformation $\phi(t)=-\cot(\pi t)$ satisfy the following.
 Lemma 4.5.
 For every $\tau\in\mathbb{N}$, $\phi^{(\tau)}\circ\phi^{-1}(x)$ is a polynomial in $x$ of degree $\tau+1$.
-Proof.
+Proof 4.6.
 Noting $\phi^{(\tau)}(t)=-\pi^{\tau}\cot^{(\tau)}(\pi t)$ and $\phi^{-1}(x)=\frac{1}{\pi}\cot^{-1}(-x)$, it suffices to show that $\cot^{(\tau)}\circ\cot^{-1}(x)$ is a polynomial of degree $\tau+1$.
-From [34, Equation (5)] we know that $P_{\tau}(x):=\tan^{(\tau)}\circ\tan^{-1}(x)$ is a polynomial of degree $\tau+1$ satisfying
+From [undefah, Equation (5)] we know that $P_{\tau}(x):=\tan^{(\tau)}\circ\tan^{-1}(x)$ is a polynomial of degree $\tau+1$ satisfying
 $P_{\tau+1}=\sum_{r=0}^{\tau}\binom{\tau}{r}P_{\tau-r}(x)P_{r}(x),$
-with $P_{0}(x)=x$ and $P_{1}(x)=x^{2}+1$. With $\cot(t)=\tan(\pi/2-t)$, the statement follows. ∎
-Using $\boldsymbol{\Psi}_{\mathrm{cot}}$ in (9) yields the following formula:
-$Q_{N}(f)=\frac{1}{N}\sum_{j=1}^{N}f\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t}^{(j)})\right)\rho\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t}^{(j)})\right)\prod_{k=1}^{d}\phi^{\prime}(t^{(j)}_{k}),\quad f\in H_{\rho}^{\alpha}.$ (17)
-We derive a higher-order chain rule for component-wise composition, which may be of interest by itself.
+with $P_{0}(x)=x$ and $P_{1}(x)=x^{2}+1$. With $\cot(t)=\tan(\pi/2-t)$, the statement follows.
+Next, we derive a higher-order chain rule for component-wise composition, which may be of interest by itself.
 Although this is a special case of the general Faa di Bruno formula, component-wise composition makes the formula simpler and thus easier to work with for broader context in numerical analysis.
-First, we derive a recurrence relation in the univariate setting.
-For $r\geq 1$ and $1\leq k\leq r$, let $B_{r,k}$ denote the (partial) Bell polynomial.
+First, we derive a recurrence relation in the univariate setting. For $r\geq 1$ and $1\leq{\eta}\leq r$, let $B_{r,{\eta}}$ denote the (partial) Bell polynomial, which may be defined by
+(18) $B_{r,\eta}(x_{1},\dots,x_{r})=\sum_{\boldsymbol{0}\leq\mathbf{k}\leq\boldsymbol{\eta}}r!\bigg(\prod_{\ell=1}^{r}\frac{x_{\ell}^{k_{\ell}}}{k_{\ell}!(\ell!)^{k_{\ell}}}\bigg)\mathbbm{1}_{\{|\mathbf{k}|=\eta,\,\sum_{\ell=1}^{r}\ell k_{\ell}=r\}}(\mathbf{k}),$
+where $\mathbbm{1}_{A}(\mathbf{k})$ denotes the indicator function of a set $A$, and $\boldsymbol{\eta}:=(\eta,\ldots,\eta)$; see for example [undeff, Definition 11.2].
 We use the convention $B_{0,0}=1$ and $B_{r,0}=0$ for $r\geq 1$.
-Define $C_{0,0}:=B_{0,0}=1$ and $C_{r,-1}:=0$ for all $r\in\mathbb{N}_{0}$. For $r\in\mathbb{N}$, $C_{r,0}:=B_{r,0}=0$, and for integers $1\leq\eta\leq r$, define
+We note that although $B_{r,\eta}=B_{r,\eta}(x_{1},\dots,x_{r})$ is defined as a polynomial in $x_{1},\dots,x_{r}$, the constraints $\sum_{i=1}^{r}k_{i}=\eta$ and $\sum_{i=1}^{r}ik_{i}=r$ in its definition imply that $B_{r,\eta}$ in fact depends only on $x_{1},\dots,x_{r-\eta+1}$.
+We will use the following identity, which shows that $\partial_{j}B_{r,\eta}$ is, up to a binomial coefficient, again a partial Bell polynomial.
+Although the result appears to be well known, a proof seems difficult to find in the literature, so we provide one.
+Lemma 4.7.
+Let $r\in\mathbb{N}$ and $1\leq\eta\leq r$. Then, for $j=1,\dots,r-\eta+1$, we have
+$\partial_{j}B_{r,\eta}(x_{1},\ldots,x_{r-\eta+1})=\binom{r}{j}\,B_{r-j,\eta-1}\!\left(x_{1},\ldots,x_{r-j-(\eta-1)+1}\right).$
+Proof 4.8.
+If we differentiate $B_{r,\eta}$ defined in (18) with respect to $x_{j}$, the terms with $k_{j}=0$ vanish: $\partial_{j}B_{r,\eta}=\sum_{\mathbf{e}_{j}\leq\mathbf{k}\leq\boldsymbol{\eta}}r!\frac{k_{j}\,x_{j}^{k_{j}-1}}{k_{j}!(j!)^{k_{j}}}\big(\prod_{\begin{subarray}{c}\ell=1\\ \ell\neq j\end{subarray}}^{r}\frac{x_{\ell}^{k_{\ell}}}{k_{\ell}!(\ell!)^{k_{\ell}}}\big)\mathbbm{1}_{\{|\mathbf{k}|=\eta,\,\sum_{\ell=1}^{r}\ell k_{\ell}=r\}}(\mathbf{k})$, where $\mathbf{e}_{j}$ is the unit vector whose $j$-th component is $1$ and whose other components are 0.
+We then extract the binomial coefficient $\binom{r}{j}=\frac{r!}{j!(r-j)!}$ as follows:
+$\partial_{j}B_{r,\eta}$ $=\sum_{\mathbf{e}_{j}\leq\mathbf{k}\leq\boldsymbol{\eta}}\binom{r}{j}\frac{(r-j)!x_{j}^{k_{j}-1}}{(k_{j}-1)!(j!)^{k_{j}-1}}\bigg(\prod_{\begin{subarray}{c}\ell=1\\ \ell\neq j,\end{subarray}}^{r}\frac{x_{\ell}^{k_{\ell}}}{k_{\ell}!(\ell!)^{k_{\ell}}}\bigg)\mathbbm{1}_{\{|\mathbf{k}|=\eta,\,\sum_{\ell=1}^{r}\ell k_{\ell}=r\}}(\mathbf{k}).$
+Re-indexing $k_{j}$ to $k_{j}+1$ yields
+$\partial_{j}B_{r,\eta}$ $=\binom{r}{j}\sum_{\mathbf{0}\leq\mathbf{k}\leq\boldsymbol{\eta}-\mathbf{e}_{j}}(r-j)!\frac{x_{j}^{k_{j}}}{k_{j}!(j!)^{k_{j}}}\bigg(\prod_{\begin{subarray}{c}\ell=1\\ \ell\neq j\end{subarray}}^{r}\frac{x_{\ell}^{k_{\ell}}}{k_{\ell}!(\ell!)^{k_{\ell}}}\bigg)\mathbbm{1}_{\{|\mathbf{k}|=\eta-1,\,\sum_{\ell=1}^{r}\ell k_{\ell}=r-j\}}(\mathbf{k})$
+$=\binom{r}{j}\sum_{\mathbf{0}\leq\mathbf{k}\leq\boldsymbol{\eta}-\mathbf{1}}(r-j)!\frac{x_{j}^{k_{j}}}{k_{j}!(j!)^{k_{j}}}\bigg(\prod_{\begin{subarray}{c}\ell=1\\ \ell\neq j\end{subarray}}^{r}\frac{x_{\ell}^{k_{\ell}}}{k_{\ell}!(\ell!)^{k_{\ell}}}\bigg)\mathbbm{1}_{\{|\mathbf{k}|=\eta-1,\,\sum_{\ell=1}^{r}\ell k_{\ell}=r-j\}}(\mathbf{k})$
+$=\binom{r}{j}B_{r-j,\eta-1},$
+where in the second line we used the fact that given the constraint $|\mathbf{k}|=\eta-1$, the range of summation may be replaced by $\mathbf{0}\leq\mathbf{k}\leq\boldsymbol{\eta}-\mathbf{1}$.
+This completes the proof.
+To derive a higher-order chain rule, we consider the functions $C_{r,\eta}(t)$, defined by evaluating partial Bell polynomials at the derivatives of $\Phi$.
+Define $C_{0,0}:=B_{0,0}=1$ and $C_{r,-1}:=0$ for all $r\in\mathbb{N}_{0}$.
+For $r\in\mathbb{N}$, $C_{r,0}:=B_{r,0}=0$, and for integers $1\leq\eta\leq r$, define
 $C_{r,\eta}(t):=B_{r,\eta}\bigl(\Phi^{\prime}(t),\Phi^{\prime\prime}(t),\ldots,\Phi^{(r-\eta+1)}(t)\bigr).$
@@
 $C_{r+1,\eta}(t)=\Phi^{\prime}(t)\,C_{r,\eta-1}(t)+C^{\prime}_{r,\eta}(t),\quad\text{for all }t\in(0,1).$
-Proof.
-The cases $r=0$ as well as $r\in\mathbb{N}$ and $\eta=0$ are easy to check. We show the statement for $1\leq\eta\leq r$. First, the partial Bell polynomials satisfy the recurrence relation
+Proof 4.10.
+The cases $r=0$ as well as $r\in\mathbb{N}$ and $\eta=0$ are easy to check.
+We show the statement for $1\leq\eta\leq r$.
+First, the partial Bell polynomials satisfy the recurrence relation [undeff, Equation (11.11)],
 $C_{r+1,\eta}(t)$ $=B_{r+1,\eta}\bigl(\Phi^{\prime}(t),\ldots,\Phi^{(r-\eta+2)}(t)\bigr)$
-$=\sum_{j=0}^{r-\eta+1}\binom{r}{j}\Phi^{(j+1)}(t)\;B_{r-j,\eta-1}\bigl(\Phi^{\prime}(t),\ldots,\Phi^{(r-j-\eta+2)}(t)\bigr).$
-They also satisfy another relation
-$\partial_{j}B_{r,\eta}(x_{1},\ldots,x_{r-\eta+1})$ $=\binom{r}{j}B_{r-j,\eta-1}(x_{1},\ldots,x_{r-j-(\eta-1)+1}),\quad j=1,\dots,r-\eta+1,$
-and thus differentiating $C_{r,\eta}$ gives
+$=\sum_{{q}=0}^{r-\eta+1}\binom{r}{{q}}\Phi^{({q}+1)}(t)\;B_{r-{q},\eta-1}\bigl(\Phi^{\prime}(t),\ldots,\Phi^{(r-{q}-\eta+2)}(t)\bigr).$
+and thus, together with Lemma 4.7, differentiating $C_{r,\eta}$ gives
 $C^{\prime}_{r,\eta}(t)$ $=\sum_{j=1}^{r-\eta+1}\partial_{j}B_{r,\eta}\Big|_{\begin{subarray}{c}x_{m}=\Phi^{(m)}(t)\\ m=1,\dots,r-\eta+1\end{subarray}}\Phi^{(j+1)}(t)$
 $=\sum_{j=1}^{r-\eta+1}\binom{r}{j}B_{r-j,\eta-1}\bigl(\Phi^{\prime}(t),\ldots,\Phi^{(r-j-\eta+2)}(t)\bigr)\;\Phi^{(j+1)}(t).$
-Hence, subtracting $C_{r+1,\eta}(t)$ from $C^{\prime}_{r,\eta}(t)$ completes the proof. ∎
+Hence, subtracting $C_{r+1,\eta}(t)$ from $C^{\prime}_{r,\eta}(t)$ completes the proof.
 For $\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}$ define
-$C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t}):=\prod_{j=1}^{d}C_{j;r_{j},\eta_{j}}(t_{j}).$
+$C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t}):=\prod_{j=1}^{d}C_{r_{j},\eta_{j}}(t_{j}).$
 We will prove the higher-order chain rule using the following analogous formula for smooth functions.
@@
 For every multi-index $\mathbf{r}\in\mathbb{N}_{0}^{d}$ with $|\mathbf{r}|_{\infty}\leq\alpha$
-$D^{\mathbf{r}}(v\circ\boldsymbol{\Phi})(\boldsymbol{t})=\sum_{\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}}(D^{\boldsymbol{\eta}}v\circ\boldsymbol{\Phi})(\boldsymbol{t})\;C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})$ (18)
+(19) $D^{\mathbf{r}}(v\circ\boldsymbol{\Phi})(\boldsymbol{t})=\sum_{\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}}(D^{\boldsymbol{\eta}}v\circ\boldsymbol{\Phi})(\boldsymbol{t})\;C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})$
 holds for every $\boldsymbol{t}\in(0,1)^{d}$, where $\boldsymbol{\eta}\leq\mathbf{r}$ means $\eta_{j}\leq r_{j}$ for all $j=1,\dots,d$.
-Proof.
+Proof 4.12.
 We use induction on $|\mathbf{r}|=r_{1}+\dots+r_{d}$.
@@
 Since only the $j$-th component of $\boldsymbol{\Phi}$ depends on $t_{j}$, we have $\partial_{{}_{j}}\bigl(v_{\boldsymbol{\eta}}\circ\boldsymbol{\Phi}\bigr)(\boldsymbol{t})=v_{\boldsymbol{\eta}+\mathbf{e}_{j}}(\boldsymbol{\Phi}(\boldsymbol{t}))\,\Phi_{j}^{\prime}(t_{j})$.
-Moreover, $\partial_{j}C_{\mathbf{r},\boldsymbol{\eta}}=\bigl(\partial_{j}C_{j;r_{j},\eta_{j}}\bigr)\prod_{k\neq j}C_{k;r_{k},\eta_{k}}=0$ if $\eta_{j}=0$.
+Moreover, $\partial_{j}C_{\mathbf{r},\boldsymbol{\eta}}=\bigl(\partial_{j}C_{r_{j},\eta_{j}}\bigr)\prod_{k\neq j}C_{r_{k},\eta_{k}}=0$ if $\eta_{j}=0$.
 Combining these, we obtain
@@
 We further rewrite the right-hand side. For the first term, from $B_{r+1,r+1}(x)=x^{r+1}=x\,B_{r,r}(x)$ for $r\geq 0$ and $\eta_{j}-1=r_{j}$,
-$\Phi_{j}^{\prime}(t_{j})C_{\mathbf{r},\boldsymbol{\eta}-\mathbf{e}_{j}}(\boldsymbol{t})$ $=\Phi_{j}^{\prime}(t_{j})C_{j;r_{j},r_{j}}(t_{j})\prod_{k\neq j}C_{k;r_{k},\eta_{k}}(t_{k})$
-$=C_{j;r_{j}+1,r_{j}+1}(t_{j})\prod_{k\neq j}C_{k;r_{k},\eta_{k}}(t_{k})=C_{\mathbf{r}+\mathbf{e}_{j},\boldsymbol{\eta}}(\boldsymbol{t})$
+$\Phi_{j}^{\prime}(t_{j})C_{\mathbf{r},\boldsymbol{\eta}-\mathbf{e}_{j}}(\boldsymbol{t})$ $=\Phi_{j}^{\prime}(t_{j})C_{r_{j},r_{j}}(t_{j})\prod_{k\neq j}C_{r_{k},\eta_{k}}(t_{k})$
+$=C_{r_{j}+1,r_{j}+1}(t_{j})\prod_{k\neq j}C_{r_{k},\eta_{k}}(t_{k})=C_{\mathbf{r}+\mathbf{e}_{j},\boldsymbol{\eta}}(\boldsymbol{t})$
 holds; for the second term, we use Lemma 4.9 to see
-$\Phi_{j}^{\prime}(t_{j})C_{\mathbf{r},\boldsymbol{\eta}-\mathbf{e}_{j}}(\boldsymbol{t})+\partial_{{}_{j}}C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})$ $=\Bigl(\Phi_{j}^{\prime}(t_{j})C_{j;r_{j},\eta_{j}-1}(t_{j})+\partial_{j}C_{j;r_{j},\eta_{j}}(t_{j})\Bigr)\prod_{k\neq j}C_{k;r_{k},\eta_{k}}(t_{k})$
-$=\Bigl(C_{j;r_{j}+1,\eta_{j}}(t_{j})\Bigr)\prod_{k\neq j}C_{k;r_{k},\eta_{k}}(t_{k})=C_{\mathbf{r}+\mathbf{e}_{j},\boldsymbol{\eta}}(\boldsymbol{t}).$
+$\Phi_{j}^{\prime}(t_{j})C_{\mathbf{r},\boldsymbol{\eta}-\mathbf{e}_{j}}(\boldsymbol{t})+\partial_{{}_{j}}C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})$ $=\Bigl(\Phi_{j}^{\prime}(t_{j})C_{r_{j},\eta_{j}-1}(t_{j})+\partial_{j}C_{r_{j},\eta_{j}}(t_{j})\Bigr)\prod_{k\neq j}C_{r_{k},\eta_{k}}(t_{k})$
+$=\Bigl(C_{r_{j}+1,\eta_{j}}(t_{j})\Bigr)\prod_{k\neq j}C_{r_{k},\eta_{k}}(t_{k})=C_{\mathbf{r}+\mathbf{e}_{j},\boldsymbol{\eta}}(\boldsymbol{t}).$
 Hence,
@@
 $=\sum_{\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}+\mathbf{e}_{j}}v_{\boldsymbol{\eta}}(\boldsymbol{\Phi}(\boldsymbol{t}))\,C_{\mathbf{r}+\mathbf{e}_{j},\boldsymbol{\eta}}(\boldsymbol{t}),$
-which is (18) for $\mathbf{r}+\mathbf{e}_{j}$. This completes the proof. ∎
+which is (19) for $\mathbf{r}+\mathbf{e}_{j}$. This completes the proof.
 Proposition 4.13.
@@
 Then,
-$D^{\mathbf{r}}(v\circ\boldsymbol{\Phi})(\boldsymbol{t})=\sum_{\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}}(D^{\boldsymbol{\eta}}v\circ\boldsymbol{\Phi})(\boldsymbol{t})\;C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})$ (19)
+(20) $D^{\mathbf{r}}(v\circ\boldsymbol{\Phi})(\boldsymbol{t})=\sum_{\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}}(D^{\boldsymbol{\eta}}v\circ\boldsymbol{\Phi})(\boldsymbol{t})\;C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})$
 holds for a.e. $\boldsymbol{t}\in(0,1)^{d}$.
-Proof.
+Proof 4.14.
 It suffices to show that the right-hand side of (20) is locally integrable and that
@@
 Local integrability follows from the boundedness of $(\Phi_{j}^{-1})^{\prime}$ and the derivatives $\Phi_{j}^{(\tau)}$ on compact subsets of $(0,1)^{d}$, as well as from the facts that $D^{\boldsymbol{\eta}}v$ is locally integrable on $\mathbb{R}^{d}$, and that the image of each compact subset of $(0,1)^{d}$ under $\boldsymbol{\Phi}$ is compact in $\mathbb{R}^{d}$.
-To show the equality above, fix $\varphi$ and take an open set $U$ such that $\mathrm{supp}(\varphi)\subset U\Subset(0,1)^{d}$.
+To show the equality above, fix $\varphi$ and take an open set $U$ such that $\mathrm{supp}(\varphi)\subset U\Subset(0,1)^{d}$, i.e., the closure $\overline{U}$ of $U\subset\mathbb{R}^{d}$ is compact and $\overline{U}\subset(0,1)^{d}$.
 Noting that $(\Phi_{j}^{-1})^{\prime}$ and $\Phi_{j}^{(\tau)}$ are bounded on $U$, we invoke Lemma 4.11 together with a standard argument for change-of-variables for Sobolev functions to conclude that the equality holds on $U$, and thus equivalently on $(0,1)^{d}$.
-∎
+Having established these identities for derivatives, we are now in a position to derive bounds for the transformed integrand and its derivatives.
 Note that, unlike in the common setting for Sobolev spaces, we do not assume boundedness of the derivatives of $\boldsymbol{\Phi}$, since the derivatives of the change of variables $\boldsymbol{\Phi}_{\mathrm{cot}}$ we consider are unbounded.
@@
 We have the following: for all multi-indices $\mathbf{r}\in\{0,1,\ldots,\alpha\}^{d}$
-$\|D^{\mathbf{r}}g\|_{L^{2}([0,1]^{d})}\leq C\|f\|_{H_{\rho}^{\alpha}},$ (20)
-where $C>0$ is a constant independent of $\mathbf{r}$ and $f$; for every $\ell=1,\dots,d$, every $\mathbf{r}\in\{0,\ldots,\alpha\}^{d}$ such that $r_{\ell}\leq\alpha-1$, and almost every $(t_{1},\dots,t_{\ell-1},t_{\ell+1},\dots,t_{d})\in(0,1)^{d-1}$,
-$\lim_{t_{\ell}\to 0+}D^{\mathbf{r}}g(\boldsymbol{t})=\lim_{t_{\ell}\to 1-}D^{\mathbf{r}}g(\boldsymbol{t})=0,$ (21)
+(21) $\|D^{\mathbf{r}}g\|_{L^{2}([0,1]^{d})}\leq C{{}_{d.\alpha}}\|f\|_{H_{\rho}^{\alpha}},$
+where $C{{}_{d.\alpha}}>0$ is a constant independent of $\mathbf{r}$ and $f$ but depends on $d$ and $\alpha$; for every $\ell=1,\dots,d$, every $\mathbf{r}\in\{0,\ldots,\alpha\}^{d}$ such that $r_{\ell}\leq\alpha-1$, and almost every $(t_{1},\dots,t_{\ell-1},t_{\ell+1},\dots,t_{d})\in(0,1)^{d-1}$,
+(22) $\lim_{t_{\ell}\to 0+}D^{\mathbf{r}}g(\boldsymbol{t})=\lim_{t_{\ell}\to 1-}D^{\mathbf{r}}g(\boldsymbol{t})=0,$
 and as a consequence, the map $t_{\ell}\mapsto D^{\mathbf{r}}g(\boldsymbol{t})$ extends continuously to $[0,1]$ by zero.
 Furthermore, for $\mathbf{r}\in\{0,\ldots,\alpha-1\}^{d}$, (22) holds for every $(t_{1},\dots,t_{\ell-1},t_{\ell+1},\dots,t_{d})\in(0,1)^{d-1}$.
-Proof.
+Proof 4.16.
 We first prove (21) for all $\mathbf{r}\in\{0,1,\ldots,\alpha\}^{d}$. By Proposition 4.13, we have
@@
 $|w(\boldsymbol{t})|$ $=\bigg|\Big(([D^{\boldsymbol{\lambda}_{1}}f]\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Big)\Big(([D^{\boldsymbol{\lambda}_{2}}\rho]\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Big)\prod_{k=1}^{d}\prod_{\tau=1}^{\lambda_{3,k}}\big(D^{\tau}\phi(t_{k})\big)^{\ell_{k,\tau}}\bigg|$
-$\leq\bigg|\Big(([D^{\boldsymbol{\lambda}_{1}}f]\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Big)\Big((\rho^{1/2+\varepsilon}\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Big)\bigg|$ (22)
+(23) $\leq\bigg|\Big(([D^{\boldsymbol{\lambda}_{1}}f]\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Big)\Big((\rho^{1/2+\varepsilon}\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Big)\bigg|$
 $\;\;\times\bigg|\Big(([D^{\boldsymbol{\lambda}_{2}}\rho]\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Big)\Big((\rho^{-1/2-\varepsilon}\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Big)\prod_{k=1}^{d}\prod_{\tau=1}^{\lambda_{3,k}}\big(D^{\tau}\phi(t_{k})\big)^{\ell_{k,\tau}}\bigg|$
@@
 so $\bigl\|D^{\mathbf{k}}\bigl(\rho^{1/2+\varepsilon}D^{\boldsymbol{\lambda}_{1}}f\bigr)\bigr\|_{L^{2}(\mathbb{R}^{d})}\leq\sum_{\mathbf{j\leq k}}\binom{\mathbf{k}}{\mathbf{j}}c_{\mathbf{j},\varepsilon}\|\rho^{1/2}D^{\mathbf{k}-\mathbf{j}+\boldsymbol{\lambda}_{1}}f\|_{L^{2}(\mathbb{R}^{d})}<\infty$.
-Showing $v\in W_{\mathrm{mix}}^{1,2}(\mathbb{R})$ when we only have $r_{\ell}\leq\alpha-1$ is analogous.
+Showing $v\in W_{\mathrm{mix}}^{1,2}(\mathbb{R})$ a.e.
+in $\mathbb{R}^{d-1}$ when we only have $r_{\ell}\leq\alpha-1$ is analogous.
 As a result, noting that $(\Psi_{\mathrm{cot}}^{-1},\dots,\Psi_{\mathrm{cot}}^{-1})\colon\mathbb{R}^{d-1}\to(0,1)^{d-1}$ maps null sets to null sets, we see that $t_{\ell}\mapsto D^{\mathbf{r}}g(\boldsymbol{t})$ admits a continuous representative $(0,1)^{d-1}$-a.e.
@@
 Thus we have proved (22).
-∎
 Using Proposition 4.15, we will show that if $f\in H_{\rho}^{\alpha}$, then after a Möbius transformation it belongs to Sobolev-type spaces on the unit cube, where QMC rules yield small errors.
@@
 The second is the unanchored Sobolev space, defined by (12) with $\boldsymbol{a}=\boldsymbol{0}$ and $\boldsymbol{b}=\boldsymbol{1}$.
-For the Korobov space, the rank-$1$ lattice rule constructed by [12, Algorithm 3.14] achieves the convergence rate $N^{-\alpha}(\ln N)^{d\alpha}$.
-For the unanchored Sobolev space, higher-order digital nets achieve the optimal rate $N^{-\alpha}(\ln N)^{(d-1)/2}$ [29].
+For the Korobov space, the rank-$1$ lattice rule constructed by [undefl, Algorithm 3.14] achieves the convergence rate $N^{-\alpha}(\ln N)^{d\alpha}$.
+For the unanchored Sobolev space, higher-order digital nets achieve the optimal rate $N^{-\alpha}(\ln N)^{(d-1)/2}$ [undefac].
 This rate is optimal, including the logarithmic factor, among all linear and nonlinear quadrature classes that rely solely on function evaluations.
 QMC rules for $\mathbb{R}^{d}$ with Möbius transformation inherit these results: those based on rank-$1$ lattice rules achieve the optimal rate up to a logarithmic factor, whereas those based on higher-order digital nets inherit full optimality, including the logarithmic factor, as we show below.
-Remark 1 (Evaluation at the boundary).
+Remark 4.17 (Evaluation at the boundary).
 Lattice points and digital nets on the unit cube include boundary points such as $\boldsymbol{0}$, where the Möbius transformation is undefined.
@@
 Let $Q_{N}$ be the rank-$1$ lattice rule with cotangent transform as in (17).
-Then, with a generating vector $\boldsymbol{z}$ obtained by the CBC construction in [12, Algorithm 3.14], $Q_{N}$ satisfies
-$\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}\leq C_{\alpha,d}N^{-\alpha}(\ln N)^{d\alpha}.$ (23)
-Proof.
+Then, with a generating vector $\boldsymbol{z}$ obtained by the CBC construction in [undefl, Algorithm 3.14], $Q_{N}$ satisfies
+(24) $\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}\leq C_{\alpha,d}N^{-\alpha}(\ln N)^{d\alpha}.$
+Proof 4.19.
 For $f\in H_{\rho}^{\alpha}$, set $g(\boldsymbol{t})=f\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\right)\rho\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\right)\prod_{k=1}^{d}\phi^{\prime}(t_{k})$ for $\boldsymbol{t}\in(0,1)^{d}$.
@@
 Noting that $t_{\ell}\mapsto D^{\mathbf{r}}g(\boldsymbol{t})\mathrm{e}^{-2\pi\mathrm{i}\boldsymbol{t}\cdot\mathbf{k}}$ is absolutely continuous on the interval $(0,1)$ for a.e.
-$\boldsymbol{t}_{-\ell}\in(0,1)^{d}$, we use integration by parts in $(0,1)$ [44, Corollary 3.23], as well as uniformly continuity up to $[0,1]$ and the integrability of functions under consideration on $(0,1)$, to obtain
+$\boldsymbol{t}_{-\ell}\in(0,1)^{d}$, we use integration by parts in $(0,1)$ [undefat, Corollary 3.23], as well as uniformly continuity up to $[0,1]$ and the integrability of functions under consideration on $(0,1)$, to obtain
 $\int_{0}^{1}D^{\mathbf{r}+\mathbf{e}_{\ell}}g(\boldsymbol{t})\mathrm{e}^{-2\pi\mathrm{i}\boldsymbol{t}\cdot\mathbf{k}}\mathrm{d}t_{\ell}$ $=2\pi\mathrm{i}k_{\ell}\int_{0}^{1}D^{\mathbf{r}}g(\boldsymbol{t})\mathrm{e}^{-2\pi\mathrm{i}\boldsymbol{t}\cdot\mathbf{k}}\mathrm{d}t_{\ell},$
@@
 $=\sum_{|\mathbf{r}|_{\infty}\leq\alpha}\|D^{\mathbf{r}}g\|_{L^{2}((0,1)^{d})}^{2}\leq C\|f\|_{H_{\rho}^{\alpha}}^{2}<\infty.$
-Finally, since $Q_{N}$ satisfies $|I(f)-Q_{N}(f)|\leq C\|g\|_{\mathrm{Kor},\alpha}N^{-\alpha}(\ln N)^{d\alpha}$, the statement follows. ∎
+Finally, since $Q_{N}$ satisfies $|I(f)-Q_{N}(f)|\leq C\|g\|_{\mathrm{Kor},\alpha}N^{-\alpha}(\ln N)^{d\alpha}$, the statement follows.
 Proposition 4.20 (Digital net with Möbius transformation).
@@
 Then, $Q_{N}$ satisfies
-$\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}\leq C_{\alpha,d}N^{-\alpha}(\ln N)^{(d-1)/2}.$ (24)
-Proof.
-From [29, Theorem 1], the QMC under study $Q_{N}$ satisfies $|I(f)-Q_{N}(f)|\leq C_{d,\alpha}\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{0},\boldsymbol{1}])}N^{-\alpha}(\ln N)^{(d-1)/2}$, where $g(\boldsymbol{t})=f\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\right)\rho\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\right)\prod_{k=1}^{d}\phi^{\prime}(t_{k})$, and $\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{0},\boldsymbol{1}])}$ is defined in (12).
+(25) $\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}\leq C_{\alpha,d}N^{-\alpha}(\ln N)^{(d-1)/2}.$
+Proof 4.21.
+From [undefac, Theorem 1], the QMC under study $Q_{N}$ satisfies $|I(f)-Q_{N}(f)|\leq C_{d,\alpha}\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{0},\boldsymbol{1}])}N^{-\alpha}(\ln N)^{(d-1)/2}$, where $g(\boldsymbol{t})=f\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\right)\rho\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\right)\prod_{k=1}^{d}\phi^{\prime}(t_{k})$, and $\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{0},\boldsymbol{1}])}$ is defined in (12).
 From Proposition 4.15, this norm can be bounded in terms of $\|f\|_{H_{\rho}^{\alpha}}$ as
-$\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{0},\boldsymbol{1}])}^{2}\leq\sum_{\mathbf{r}\in\{0,\dots,\alpha\}^{d}}\|D^{\mathbf{r}}g\|_{L^{2}((0,1)^{d})}^{2}\leq C_{d,\alpha}\|f\|_{H_{\rho}^{\alpha}}.$
-This completes the proof. ∎
-We summarise the results thus far in the following theorem.
-Theorem 4.10.
-The quasi-Monte Carlo rule $Q_{N}$ (9) achieves the error bound
-$\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}$ $\leq C_{d,\alpha}\frac{(\ln N)^{\gamma}}{N^{\alpha}},$
-where the exponent $\gamma$ depends on the QMC construction:
-[…]
-The polynomial convergence rate $\alpha$ is optimal in $H_{\rho}^{\alpha}$: no algorithm can achieve a better polynomial rate.
-Furthermore, the exponent (d1)/2(d-1)/2 in the logarithmic factor is also optimal.
+$\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{0},\boldsymbol{1}])}^{2}\leq\sum_{\mathbf{r}\in\{0,\dots,\alpha\}^{d}}\|D^{\mathbf{r}}g\|_{L^{2}((0,1)^{d})}^{2}\leq C_{d,\alpha}\|f\|_{H_{\rho}^{\alpha}}^{{2}}.$
+This completes the proof.
 5 Concluding remarks
@@
 One is to analyse sparse-grid algorithms for functions over $\mathbb{R}^{d}$ built from univariate rules other than the Gauss–Hermite rule.
-A notable alternative is the sparse-grid method using weighted Leja sequences [22, 48], which was proposed by Narayan and Jakeman [48].
-To obtain results analogous to what we showed in this paper, it is worth noting that the fooling function we constructed in Theorems 3.1 and  3.2 depend solely on the distribution of quadrature nodes, and not on the quadrature weights.
+A notable alternative is the sparse-grid method using weighted Leja sequences [undefv, undefax], which was proposed by Narayan and Jakeman [undefax].
+To obtain results analogous to what we showed in this paper, it is worth noting that the fooling function we constructed in Theorems 3.1 and 3.3 depend solely on the distribution of quadrature nodes, and not on the quadrature weights.
 Hence, understanding the distribution of Leja points for the weight of interest would allow the construction of a corresponding fooling function.
@@
 Addressing this is left for future work.
-A standard approach is to employ the so-called weighted spaces, as introduced by Hickernell [33], under which QMC methods can attain dimension-independent convergence rates with constants that are likewise independent of dimension, as shown by Sloan and Woźniakowski [59].
-These ideas have been widely applied in the analysis of parametric partial differential equations [30, 32, 39, 40, 58].
-In fact, the results in [59] show that such weights are even necessary in the setting considered there, suggesting that additional structure is required to make the constants in our bounds dimension-independent.
+A standard approach is to employ the so-called weighted spaces, as introduced by Hickernell [undefag], under which QMC methods can attain dimension-independent convergence rates with constants that are likewise independent of dimension, as shown by Sloan and Woźniakowski [undefaai].
+These ideas have been widely applied in the analysis of parametric partial differential equations [undefad, undefaf, undefan, undefao, undefaah].
+In fact, the results in [undefaai] show that such weights are even necessary in the setting considered there, suggesting that additional structure is required to make the constants in our bounds dimension-independent.
 This paper focused on Sobolev functions.
-Another important regime in high dimension is that of analytic functions [8, 35, 17], where the question of optimal rates in this setting remains largely open.
-In establishing the lower bound for the sparse-grid Gauss–Hermite rule in this paper, a crucial component was the fooling function constructed for univariate Sobolev functions [41].
-A similar technique for analytic functions is available [27, 63], which may prove useful for further analysis in this regime.
+Another important regime in high dimension is that of analytic functions [undefh, undefai, undefq], where the question of optimal rates in this setting remains largely open.
+In establishing the lower bound for the sparse-grid Gauss–Hermite rule in this paper, a crucial component was the fooling function constructed for univariate Sobolev functions [undefap].
+A similar technique for analytic functions is available [undefaa, undefaam], which may prove useful for further analysis in this regime.